### v2

# Resilience for Regular Path Queries: Towards a Complexity Classification

Abstract.
The resilience problem for a query and an input set or bag database is to compute the minimum number of facts to remove from the database to make the query false. In this paper, we study how to compute the resilience of Regular Path Queries (RPQs) over graph databases. Our goal is to characterize the regular languages $L$ for which it is tractable to compute the resilience of the existentially-quantified RPQ built from $L$.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter. We last highlight some remaining obstacles towards a full dichotomy.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
{toappendix}
1. Introduction
[…]
Figure 1. Overview of our results on the complexity of resilience (all results apply both to bag and set semantics), with some examples of languages. All languages are reduced. For legibility, we omit Section 5.6.
This paper studies the resilience problem which asks, for a Boolean query $Q$ and a given database $D$ (in set or bag semantics) what is a minimum-cardinality set of tuples to remove from $D$ so that $Q$ is no longer satisfied. (In particular: if $Q$ does not hold on $D$ then resilience is zero; if $Q$ holds on $D$ and all subinstances of $D$ then resilience is infinite.) The notion of resilience is motivated by the need to identify how “robust” a query is, especially when the facts of the input database may be incorrect. Specifically, resilience quantifies how many tuples of $D$ must be removed before $Q$ no longer holds, so that higher resilience values mean that $Q$ is less sensitive to tuples of $D$ being removed.
The resilience problem has been introduced for Boolean conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; bodirsky2024complexity; makhija2024unified). The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified)). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
This leads to a natural question: Can one give a characterization of the queries for which resilience can be tractably computed? Further, can one show a dichotomy result to establish that resilience is always either in PTIME or NP-complete, depending on the query? This question has remained elusive, already in the case of conjunctive queries (CQs).
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, the only other work which studies this problem is the recent paper of Bodirsky et al. (bodirsky2024complexity), in which the authors use CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for Boolean RPQs (and also for the broader class of so-called two-way RPQs aka 2RPQs). However, their lower bounds do not apply to databases in set semantics. In addition, their dichotomy criterion, while decidable, does not give a intuitive language-theoretic characterization of the tractable cases. We discuss this in more details at the end of Section 2.
In contrast with CQs, RPQs are in certain ways harder to analyze, and in other ways easier. On the one hand, RPQs feature the union operator and the recursion operator; and they inherently correspond to queries with self-joins, because the same relation symbols (corresponding to alphabet letters) may occur multiple times in query matches. On the other hand, RPQs do not allow all features of CQs: they ask for a simple pattern shape (namely, a walk, which importantly is always directed), and they are posed over graph databases (in which all facts have arity two).
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem, allowing multiple sources and sinks without loss of generality, is easily seen to be equivalent to the resilience problem in bag semantics for the RPQ $ax^{*}b$ on the alphabet $\{a,b,x\}$, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
Contributions and outline.
After giving preliminary definitions and stating the problem in Section 2, we start in Section 3 by giving our main tractability result (Theorem 3.3): resilience is tractable for RPQs defined from so-called local languages (mpri; sakarovitch2009elements). Those languages are intuitively those that are defined by specifying the possible initial labels of a path, the possible final labels of a path, and the consecutive pairs of labels that are allowed. We show that, for RPQs defined with a local language $L$, we can compute resilience by reducing to the MinCut problem via a natural product construction. The construction applies to queries both in set and bag semantics, and it is also tractable in combined complexity, running in time $\tilde{O}(|A|\times|\Sigma|\times|D|)$ where $A$ is an automaton representing $L$, where $\Sigma$ is the signature, and where $D$ is the input database.
We then turn to intractability results in Sections 4, 5 and 6. In all these results, we assume that the language has been reduced to eliminate redundant words, e.g., resilience for $abbc|bb$ and $bb$ is the same because the corresponding (existentially-quantified) Boolean RPQs are equivalent.
We define in Section 4 the notion of a hardness gadget for a language $L$, namely, a specific database to be used as a building block in a reduction. These gadgets are the analogue, in our setting, of the Independent Join Paths (IJPs) used in earlier work (makhija2024unified) to show hardness for CQs. When an RPQ $Q$ has a hardness gadget $\Gamma$, we can reduce the minimum vertex cover problem on an undirected graph $G$ to a resilience computation (in set semantics) for $Q$ on a database built using $\Gamma$, by replacing the edges of $G$ by a copy of $\Gamma$. We illustrate in Section 4 how our gadgets can be used to show hardness for two specific RPQs: $axb|cxd$ (Proposition 4.8), and $aa$ (Proposition 4.1).
We then generalize these two hardness results in turn to apply to a broader class of languages. First, in Section 5, we introduce the notion of a four-legged language, as a language featuring a specific kind of counterexample to being local. More precisely, the local languages of Section 3 have the following known equivalent characterization (sakarovitch2009elements): for any letter $x$ and words $\alpha,\beta,\gamma,\delta$, if a language $L$ contains the words $\alpha x\beta$ and $\gamma x\delta$ but does not contain the “cross-product” word $\alpha x\delta$, then $L$ is not local. A four-legged language is a non-local language that features a counterexample of this form in which all the words $\alpha,\beta,\gamma,\delta$ are required to be non-empty. This covers, for instance, $axb|cxd$, but not $ab|bc$. We show (Section 5.2) that resilience is intractable for any four-legged language. This result implies hardness for many languages: it covers all non-star-free languages (Section 5.5), a well-known class of infinite regular languages; and, along with the tractability of local languages, it implies a dichotomy for languages having a neutral letter (Section 5.6).
Second, in Section 6, we show that resilience is hard for any finite reduced language containing a word with a repeated letter (Section 6.1). This generalizes the case of $aa$, and shows that self-joins always imply hardness in the following sense: finite RPQs (those equivalent to UCQs), once reduced, have intractable resilience as soon as one of their constituent CQs features a self-join. This contrasts with the setting of CQs, where characterizing the tractable queries with self-joins is still open.
Note that our two hardness results (Section 5.2 and Section 6.1) are incomparable: languages with a repeated letter are non-local, but they are not all four-legged (e.g., $aa$); conversely, some four-legged languages have no words with a repeated letter (e.g., $axb|cxd$). Assuming finiteness in Section 6.1 is also crucial, as $ax^{*}b$ features words with repeated letters and still enjoys tractable resilience. The proof of Section 6.1 is our main technical achievement: the proof relies on Section 5.2 and uses different gadgets depending on which additional words the language contains.
Section 7 explores how our results do not classify all RPQs. We focus on finite RPQs: the remaining open cases are then the RPQs which have no words with repeated letters, are not local, but are not four-legged. We introduce the class of chain queries (e.g., $axb|byc$) to cover some of these RPQs, and show tractability (and reducibility to MinCut) assuming a bipartiteness condition. We also study the RPQ $abc|be$ and variants thereof, and show that it enjoys tractable resilience via techniques from submodular optimization (Section 7.4), which seems new in this context. This query (i.e., $R(x,y),S(y,z),T(z,u)\vee S(x,y),U(y,z)$ as a UCQ), and variants thereof, are currently the only known tractable RPQs with no known reduction to MinCut. We conclude by showing that some other unclassified RPQs are hard, and by discussing some remaining open cases.
We finally conclude in Section 8. For lack of space, most proofs are deferred to the appendix.
2. Preliminaries and Problem Statement
Formal languages.
Let $\Sigma$ be an alphabet, consisting of letters. A word over $\Sigma$ is a finite sequence of elements of $\Sigma$. We write $\epsilon$ for the empty word. We write $\Sigma^{*}$ for the set of words over $\Sigma$, and write $\Sigma^{+}\coloneq\Sigma^{*}\setminus\{\epsilon\}$. We use lowercase Latin letters such as $a,b,x,y$ to denote letters in $\Sigma$, and lowercase Greek letters like $\alpha,\beta,\gamma,\tau$ to denote words in $\Sigma^{*}$. The concatenation of two words $\alpha,\beta$ is written $\alpha\beta$. We write $|\alpha|$ for the length of a word. A word $\alpha$ is an infix of another word $\beta$ if $\beta$ is of the form $\delta\alpha\gamma$, and it is a strict infix if moreover $\delta\gamma\neq\epsilon$. If $\beta=\alpha\gamma$ then $\alpha$ is a prefix of $\beta$, and a strict prefix if $\gamma\neq\epsilon$. Likewise if $\beta=\delta\alpha$ then $\alpha$ is a suffix of $\beta$, and a strict suffix if $\delta\neq\epsilon$.
A language $L$ is a (possibly infinite) subset of $\Sigma^{*}$. For $L,L^{\prime}$ two languages, we write their concatenation as $LL^{\prime}\coloneq\{\alpha\beta\mid\alpha\in L,\beta\in L^{\prime}\}$. We mainly focus in this work on regular languages, also known as rational languages, which can be represented by finite state automata. We use the following automaton formalisms, which recognize precisely the regular languages (sakarovitch2009elements): deterministic finite automata (DFA), nondeterministic finite automata (NFA), and NFAs with $\epsilon$-transitions ($\epsilon$NFA, also called automata with spontaneous transitions).
An $\epsilon$NFA $A=(S,I,F,\Delta)$ is a finite set $S$ of states, a set $I\subseteq S$ of initial states, a set $F\subseteq S$ of final states, and a transition relation $\Delta\subseteq S\times(\Sigma\cup\{\epsilon\})\times S$. Let $\alpha=a_{1}\cdots a_{m}$ be a word. A configuration for $A$ on $\alpha$ is an ordered pair $(s,i)$ for $s\in S$ and $0\leq i\leq m$. A run of $A$ on $\alpha$ is a sequence of configurations $(s_{0},i_{0}),\ldots,(s_{\ell},i_{\ell})$ with $\ell\geq m$ such that $s_{0}\in I$ is initial, $i_{0}=0$, $i_{\ell}=m$, and for each $0\leq j<\ell$ we have either $i_{j+1}=i_{j}$ and $(s_{j},\epsilon,s_{j+1})\in\Delta$ (an $\epsilon$-transition) or $i_{j+1}=i_{j}+1$ and $(s_{j},a_{i_{j}+1},s_{j+1})\in\Delta$ (a letter transition). The run is accepting if $s_{\ell}\in F$. The language of $A$, denoted $\mathcal{L}(A)$, is the set of words having an accepting run: we say that $A$ recognizes $\mathcal{L}(A)$. An example $\epsilon$NFA $A_{3}$ is depicted in Figure 2(c), and an example accepting run of $A_{3}$ on the word $ad$ is $(s_{1},0),(s_{2},1),(s_{4},1),(s_{5},2)$. The size $|A|$ of the $\epsilon$NFA $A=(S,I,F,\Delta)$ is $|S|+|\Delta|$, its total number of states and transitions.
An NFA is an $\epsilon$NFA which does not feature $\epsilon$-transitions, and a DFA is an NFA having exactly one initial state and having at most one transition $(s,a,s^{\prime})$ in $\Delta$ for each $s\in S$ and $a\in\Sigma$.
We also use regular expressions to denote regular languages: we omit their formal definition, see (sakarovitch2009elements). We write |||| for the union operator and * for the Kleene star operator.
Databases, queries, resilience.
[…]
Definition 2.1 (Resilience).
Let $D$ be a database and $Q$ a Boolean query. The resilience of $Q$ on $D$, denoted $\mathtt{RES}_{\mathrm{set}}(Q,D)$, is defined [^1] By convention, if $Q$ is satisfied by all subsets of $D$, then the minimum is defined to be +$\infty$+ ∞. as $\mathtt{RES}_{\mathrm{set}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}|D^{\prime}|$.
In other words, the resilience is the minimum number of facts to remove from $D$ so that $Q$ is not satisfied. A subset of facts $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$ is called a cut (or a contingency set) of $D$ for $Q$. Notice that if $D\not\models Q$, then $\emptyset$ is a cut and we have [^2] We deviate from the original definition in (DBLP:journals/pvldb/FreireGIM15) which left resilience undefined for the case of $D\not\models Q$. $\mathtt{RES}_{\mathrm{set}}(Q,D)=0$.
Regular Path Queries.
From a language $L$ over $\Sigma$ we now define a query over graph databases that tests the existence of walks labeled by words of $L$. Given a database $D$, a walk in $D$ is a (possibly empty) sequence of consecutive edges $v_{1}\xrightarrow{a_{1}}v_{2}$, $v_{2}\xrightarrow{a_{2}}v_{3}$, $\ldots$, $v_{k}\xrightarrow{a_{k}}v_{k+1}$ in $D$. Note that the $a_{i}$ and $v_{i}$ are not necessarily pairwise distinct. We say that the walk is labeled by the word $\alpha=a_{1}\cdots a_{k}$, and call it an $\alpha$-walk. We also say that the walk contains a word $\beta\in\Sigma^{*}$ if $\beta$ is an infix of $\alpha$. An $L$-walk in $D$ is then an $\alpha$-walk for some $\alpha\in L$. We define the Boolean query $Q_{L}$ as: $Q_{L}(D)=1$ iff there is some $L$-walk in $D$. In particular, if $\epsilon\in L$ then $Q_{L}(D)=1$ on every database $D$ (including $D=\emptyset$). A regular path query (RPQ) is then a Boolean query of the form $Q_{L}$ for $L$ a regular language.
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the reduced language of $L$ is $\mathtt{reduce}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{reduce}(L)}$ are the exact same queries. The reduced language of $L$ is also called the infix-free sublanguage of $L$ in (barcelo2019boundedness): if $L$ is regular then so is $\mathtt{reduce}(L)$. We say that $L$ is reduced if $L=\mathtt{reduce}(L)$.
Bag semantics.
We also consider the resilience problems on bag graph databases, by opposition with set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. In the input, the multiplicities are expressed as integers written in binary. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set for $Q$ and $D$ is then a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
Networks and cuts.
Most of our resilience algorithms work by reducing to the problem of finding a minimum cut of a flow network (Cormen:2009dz). Recall that a flow network consists of a directed graph $\mathcal{N}=(V,t_{\text{source}},t_{\text{target}},E,c)$ with $V$ being the vertices, $t_{\text{source}},t_{\text{target}}\in V$ being special vertices called the source and the target respectively, $E\subseteq V\times V$, and $c:E\to\mathbb{R}_{+}\cup\{+\infty\}$ giving the capacity of each edge. For $E^{\prime}\subseteq E$, we define the cost of $E^{\prime}$ as $c(E^{\prime})\coloneq\sum_{e\in E^{\prime}}c(e)$, and denote $\mathcal{N}\setminus E^{\prime}$ as the network in which all edges from $E^{\prime}$ are removed. A cut is a set of edges $E^{\prime}\subseteq E$ such that there is no directed path from $t_{\text{source}}$ to $t_{\text{target}}$ in $\mathcal{N}\setminus E^{\prime}$. The min-cut problem, written MinCut, takes as input a flow network and returns the minimum cost of a cut. From the max-flow min-cut theorem (ford1956maximal) and Menger’s theorem (menger:1927), we know that the min-cut problem can be solved in PTIME. More precisely, practical algorithms can run in $O(|V|^{2}\sqrt{|E|})$ (ahuja1997computational), and recent theoretical algorithms run in $\tilde{O}(|\mathcal{N}|)$ (henzinger2024deterministic), where $|\mathcal{N}|\coloneq|V|+|E|$ and where $\tilde{O}$ hides polylogarithmic factors.
Problem statement.
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-hard (makhija2024unified, Table 1) (even when the multiplicities of the input bag database are written in unary).
Our definitions correspond to the data complexity perspective, where we have one computational problem per language $L$ and where the complexity is measured as a function of $|D|$ only. We mostly focus on data complexity, but most of our upper bounds also apply in combined complexity, i.e., when both the database $D$ and an automaton $A$ representing $L$ are given as input.
It is easy to see that, for any regular language $L$, the problem $\mathtt{RES}_{\mathrm{bag}}(L)$, hence $\mathtt{RES}_{\mathrm{set}}(L)$, is in NP (see Section 2). {toappendix} We explain here why $\mathtt{RES}_{\mathrm{bag}}(L)$ for $L$ a regular language is always in NP. Note that this implies the same for $\mathtt{RES}_{\mathrm{set}}(L)$. Given a bag database $D$, it suffices to guess a subdatabase $D^{\prime}\subseteq D$, and verify in PTIME that $\sum_{f\in D^{\prime}}\mathtt{mult}(f)$ is small enough and that $D^{\prime}$ is a contingency set, i.e., $D\setminus D^{\prime}\not\models Q$. The latter point, i.e., the inexistence of an $L$-walk in $D\setminus D^{\prime}$, can easily be checked in PTIME by a product construction followed by a reachability test (see, e.g., (DBLP:journals/siamcomp/MendelzonW95, Lemma 3.1)). However, the NP upper bound is not always tight: there are languages $L$ (e.g., $L=a|b$) such that $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. The goal of our paper is to characterize the languages $L$ for which $\mathtt{RES}_{\mathrm{set}}(L)$ and $\mathtt{RES}_{\mathrm{bag}}(L)$ are in PTIME, and give efficient algorithms for them.
Connection to VCSPs.
From the recent work of Bodirsky et al. (bodirsky2024complexity) using earlier results on Valued Constraint Satisfaction Problems (VCSP), the following dichotomy on $\mathtt{RES}_{\mathrm{bag}}$ is already known:
Theorem 2.2 ((bodirsky2024complexity, Remark 8.12)).
Let $L$ be a regular language. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ is either in PTIME or it is NP-complete.
We point out that this result can be extended to so-called two-way RPQs, and to the case where some relations are declared exogenous (i.e., they cannot be part of a contingency set). Hence, as discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages, but it also has some shortcomings. (1.) First, the VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity), using (kolmogorov2019testing), can be computationally checked but takes time exponential in a representation of the dual of $L$. (2.) Second, in tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that difficulties (1.) and (2.) also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds: this is the case for instance in (hirai2021minimum). (3.) Third, the VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and database are given as input. (4.) Fourth, the VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$. Our results will improve on some of these shortcomings, though unlike (bodirsky2024complexity) they do not establish a full dichotomy.
3. Tractability for Local Languages
We now start the presentation of our results on the resilience problem by giving our main tractability result in this section. The result applies to local languages [^3] We note that local languages were also recently observed to correspond to a tractable class of RPQs for which probabilistic query evaluation is equivalent to source-to-target reliability, see (amarilli2025approximating). , a well-known subclass of the regular languages (mpri; sakarovitch2009elements), whose definition is recalled below. We show that the resilience problem is tractable for local languages, also in combined complexity, using a specific notion of automata and a product construction to reduce resilience to MinCut. We stress that local languages do not cover all cases of regular languages with tractable resilience: we will show other tractable cases in Section 7. However, in Section 5 we will give a partial converse: resilience is hard for the so-called four-legged languages, which have a specific form of counterexample to being local.
Local languages.
Intuitively, a regular language $L$ is local if membership to $L$ is described by which letter pairs can occur consecutively and which letters can start or can end words. We use an equivalent characterization (mpri, Chapter III, Section 5.1) via so-called local DFAs:
Definition 3.1 (Local DFA).
A DFA is local if, for every letter $a\in\Sigma$, all $a$-transitions in the automaton have the same target state. A language $L$ is local if there is a local DFA $A$ that recognizes $L$.
Example 3.2.
The languages $ax^{*}b$ and $ab|ad|cd$ are local, with local DFAs in Figure 2(a) and Figure 2(b).
$s_{1}$$s_{2}$$s_{3}$$a$$x$$b$
(a) Local DFA $A_{1}$ for $ax^{*}b$
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$$a$$b$$d$$c$$d$
(b) Local DFA $A_{2}$ for $ab|ad|cd$
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$$a$$b$$c$$d$$\epsilon$
(c) RO-$\epsilon$NFA $A_{3}$ for $ab|ad|cd$
Figure 2. Examples of local automata (Example 3.2) and RO-$\epsilon$NFAs (see Example 3.2 and Example 3.8). Initial states are indicated by incoming arrows, final states are doubly circled.
The main result of this section is that $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME when $L$ is local:
Theorem 3.3 (Resilience for local languages).
If $L$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. Moreover, the problem is also tractable in combined complexity: given an $\epsilon$NFA $A$ that recognizes a local language $L$, and a database $D$, we can compute $\mathtt{RES}_{\mathrm{bag}}(Q_{L},D)$ in $\tilde{O}(|A|\times|D|\times|\Sigma|)$.
Note that, in the combined complexity upper bound, the input $A$ is not necessarily a local DFA: it suffices to be given an arbitrary $\epsilon$NFA $A$ recognizing $L$, together with the promise that $L$ is local.
Before proving Theorem 3.3, we remark that it implies a PTIME bound on $\mathtt{RES}_{\mathrm{bag}}(L)$ whenever $\mathtt{reduce}(L)$ is local (as $\mathtt{RES}_{\mathrm{bag}}(L)$ is equivalent to $\mathtt{RES}_{\mathrm{bag}}(\mathtt{reduce}(L))$). For instance, $\mathtt{RES}_{\mathrm{bag}}(L_{0})$ for $L_{0}=\{a,aa\}$ is in PTIME because $\mathtt{reduce}(L_{0})=\{a\}$ is local, even though $L_{0}$ is not. In fact, to see if Theorem 3.3 applies to a language $L$, we can always focus on $\mathtt{reduce}(L)$, for the following reason:
{toappendix}
3.1. Preparatory Lemmas
Before proving the results in Section 3, we prove a few lemmas that use the notions of RO-$\epsilon$NFA (defined later in Section 3) and of letter-Cartesian languages (defined later in Definition 5.1). Specifically, we prove that the formalisms of local DFAs (Definition 3.1), of RO-$\epsilon$NFA, and of letter-Cartesian languages, are indeed equivalent. In the process, we give a general lemma about converting $\epsilon$-NFAs to a certain RO-$\epsilon$NFAs (Lemma 3.4) which we will need in some subsequent proofs.
3.1.1. Equivalence Between RO-$\epsilon$NFA and Local DFAs
First, we show that local DFAs and RO-$\epsilon$NFAs recognize the same languages. We show both directions in turn.
{claim}
Let $A$ be a local DFA. Then there is a RO-$\epsilon$NFA $A^{\prime}$ such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
Let $A=(S,I,F,\Delta)$ be the local DFA. We build from it an RO-$\epsilon$NFA $A^{\prime}=(S^{\prime},I,F,\Delta^{\prime})$ defined as follows. We let $S^{\prime}\coloneq S\cup\{s^{\prime}_{a}\mid a\in\Sigma\}$ where the states of the form $s_{a}^{\prime}$ are fresh. For every letter $a\in\Sigma$ that has a transition in $A$, letting $s_{a}$ be the common target state of all $a$-transitions in $A$, we create one $a$-transition $(s_{a}^{\prime},a,s_{a})\in\Delta^{\prime}$, and add $\epsilon$-transitions $(s,\epsilon,s_{a}^{\prime})\in\Delta^{\prime}$ from every state $s\in S$ that has an outgoing $a$-transition in $A$. The resulting automaton $A^{\prime}$ is an RO-$\epsilon$NFA by construction, and there is a clear bijection between the runs of $A$ and those of $A^{\prime}$. ∎
{claim}
Let $A$ be a RO-$\epsilon$NFA. Then there is a local DFA $A^{\prime}$ such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
Let $A$ be the RO-$\epsilon$NFA. We remove $\epsilon$-transitions in PTIME using standard techniques: first ensure that every state having an $\epsilon$-path to a final state is final; then for every letter $a\in\Sigma$ occurring in $A$, for every state $s$ having an $\epsilon$-path to the starting state $s_{a}^{\prime}$ of the unique $a$-transition $(s_{a}^{\prime},a,s_{a})$ of $A$, add an $a$-transition from $s$ to $s_{a}$; finally remove all $\epsilon$-transitions. It is clear that the resulting automaton $A^{\prime}$ is a local DFA that recognizes the same language, which concludes the proof. ∎
3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
Next, we observe that local languages are always letter-Cartesian. We first show one implication:
{claim}
If $L$ is local then it is letter-Cartesian.
Proof.
Let $L$ be a local DFA recognizing $L$. Let $x\in\Sigma$, and let $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$, and let us show that $\alpha x\delta$ is in $L$ as well. Because $L$ is local, the state $q$ that is reached after reading $\alpha x$ in the accepting run of $\alpha x\beta$ is the same state that is reached after reading $\gamma x$ in the accepting run of $\gamma x\delta$. We can thus compose a run from the initial state to $q$ reading $\alpha x$, with the run from $q$ to some final state reading $\delta$, to form an accepting run for $\alpha x\delta$. This concludes the proof. ∎
To show the second implication, let us give a lemma that encapsulates a construction of a RO-$\epsilon$NFA from an $\epsilon$NFA. We will use this construction in subsequent proofs:
Lemma 3.4.
Given an $\epsilon$NFA $A$, we can build in time $O(|A|\times|\Sigma|)$ a RO-$\epsilon$NFA $A^{\prime}$ having the following properties:
  1. (1)
    $\mathcal{L}(A)\subseteq\mathcal{L}(A^{\prime})$;
  2. (2)
    $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$ holds if and only if $\mathcal{L}(A)$ is letter-Cartesian.
To prove this result, we need some standard definitions on automata (sakarovitch2009elements, Section 1.3.1):
Definition 3.5.
Let $A=(S,I,F,\Delta)$ be an $\epsilon$NFA. We say that a state $s\in S$ is accessible if there is a path in $A$ from a state of $I$ to $s$ using transitions of $\Delta$. We say that $s$ is co-accessible if there is a path in $A$ from $s$ to a state of $F$ using transitions of $\Delta$. We say that $s$ is useful if it is both accessible and coaccessible. We say that $A$ is trimmed (or trim (sakarovitch2009elements)) if every state of $A$ is useful.
A standard result on automata (sakarovitch2009elements) is that they can be trimmed in linear time:
{claim}
Given an $\epsilon$NFA $A$, we can compute in time $O(|A|)$ an $\epsilon$NFA $A^{\prime}$ which is trimmed and such that $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$.
Proof.
We first do one graph traversal in $A$ from $I$ to compute the accessible states in $O(|A|)$. Then, we do one graph traversal from $F$ following the transitions of $\Delta$ backwards to compute the co-accessible states in $O(|A|)$. Then, we obtain $A^{\prime}$ by simply removing all states of $A$ that are not useful, and removing all transitions that involve these states. The computation of $A^{\prime}$ from $A$ is in time $O(|A|)$ overall.
It is clear that $A^{\prime}$ is then trimmed, because every remaining state $s$ is a state which was useful in $A$, and the paths witnessing it still witness that $s$ is useful in $A^{\prime}$. Further, $A^{\prime}$ recognizes the same language as $A$: the construction clearly ensures $\mathcal{L}(A^{\prime})\subseteq\mathcal{L}(A)$, and conversely there are no accepting runs in $A$ involving states that are not useful, so every accepting run in $A$ is also an accepting run of $A^{\prime}$. This concludes the proof. ∎
We can now show Lemma 3.4:
Proof of Lemma 3.4.
Let $A=(S,I,F,\Delta)$ be an $\epsilon$NFA. Write $L\coloneq\mathcal{L}(A)$. We assume by Section 3.1.2 that $A$ has been trimmed in linear time. Let us compute from $A$ the following information:
  • The set $\Sigma_{\mathrm{start}}$ of letters that can start words of $L$, i.e., the letters $a\in\Sigma$ such that there exists a word $a\alpha\in L$ for some $\alpha\in\Sigma^{*}$. This can be computed in linear time by performing a graph traversal from the initial states following the $\epsilon$-transitions, and then considering every non-$\epsilon$ transition that leaves from such states and letting $\Sigma_{\mathrm{start}}$ be the set of the letters that label such transitions.
  • The set $\Sigma_{\mathrm{end}}$ of letters that can end words of $L$, i.e., letters $a\in\Sigma$ such that $\alpha a\in L$ for some $\alpha\in\Sigma^{*}$. These can be computed in linear time in a symmetric way by doing a graph traversal from $F$, traversing $\epsilon$-transitions backwards, and taking the letters of non-$\epsilon$ transitions leaving backwards from the states reached.
  • The set $\Pi\subseteq\Sigma^{2}$ of pairs of letters that can occur consecutively in a word of $L$, formally, the pairs $(a,b)$ such that there exist $\alpha,\beta\in\Sigma^{*}$ such that $\alpha ab\beta\in L$. We compute this in time $O(|\Sigma|\times|A|)$ by taking each $a\in\Sigma$ and computing the $b\in\Sigma$ such that $(a,b)\in\Pi$. For this, we take the set of the states that are heads of $a$-transitions, and we do a linear-time graph traversal from the set of such states, following $\epsilon$-transitions, and taking all letters that label transitions that leave from states reached in the traversal.
  • The Boolean information of whether $A$ accepts the empty word, which can be computed in linear time by performing a graph traversal from the set $I$ of initial states of $A$ following the $\epsilon$-transitions and checking whether some final state of $F$ is reached.
Note that this uses the fact that the automaton is trimmed. We now define an RO-$\epsilon$NFA $A^{\prime}$ in the following way:
  1. (1)
    For every $a\in\Sigma$, create two states $s_{a}^{\mathtt{in}}$ and $s_{a}^{\mathtt{out}}$, with $(s_{a}^{\mathtt{in}},a,s_{a}^{\mathtt{out}})\in\Delta$.
  2. (2)
    For every $a\in\Sigma_{\mathrm{start}}$, make $s_{a}^{\mathtt{in}}$ an initial state.
  3. (3)
    For every $(a,b)\in\Pi$ (potentially $a=b$), add the $\epsilon$-transition $(s_{a}^{\mathtt{out}},\epsilon,s_{b}^{\mathtt{in}})$ to $\Delta$.
  4. (4)
    For every $a\in\Sigma_{\mathrm{end}}$, make $s_{a}^{\mathtt{out}}$ a final state.
  5. (5)
    If $\epsilon\in L$, create an isolated state that is both initial and final.
It is clear by definition that $A^{\prime}$ is a RO-$\epsilon$-NFA. Write $L^{\prime}\coloneq\mathcal{L}(A^{\prime})$. What remains to show is that $L\subseteq L^{\prime}$, and that we have $L=L^{\prime}$ iff $L$ is letter-Cartesian.
We start by showing that $L\subseteq L^{\prime}$. Let $\alpha\in L$. If $\alpha=\epsilon$ then $\alpha\in L^{\prime}$ by construction. Otherwise let $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. The word $\alpha$ witnesses the following:
  • We have $a_{1}\in\Sigma_{\mathrm{start}}$;
  • We have $a_{m}\in\Sigma_{\mathrm{end}}$;
  • We have $(a_{i},a_{i+1})\in\Pi$ for each $1\leq i<m$.
Thus, by construction:
  • The state $s_{a_{1}}^{\mathtt{in}}$ is initial;
  • The state $s_{a_{m}}^{\mathtt{out}}$ is final;
  • We have $(s_{a_{i}}^{\mathtt{out}},\epsilon,s_{a_{i+1}}^{\mathtt{in}})\in\Delta$ for each $1\leq i<m$.
All of this gives an accepting run of $A^{\prime}$ on $\alpha$, hence $\alpha\in L^{\prime}$ indeed.
Now, let us show the equivalence: $L$ is letter-Cartesian iff $L^{\prime}=L$. One direction is easy: if $L^{\prime}=L$, then $L$ is recognized by $A^{\prime}$, which is a RO-$\epsilon$NFA, and thus recognizes a letter-Cartesian language by Section 3.1.1 and Section 3.1.2.
For the converse direction, let us assume that $L$ is letter-Cartesian, and show that $L^{\prime}=L$. The inclusion $L\subseteq L^{\prime}$ was already shown above, so it suffices to show that $L^{\prime}\subseteq L$. Take $\alpha$ in $L^{\prime}$, let us show that it is in $L$. If $\alpha$ is the empty word then it is clear by construction that $\alpha$ is also in $L$. So assume $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. By construction of $A^{\prime}$, the fact that $\alpha$ is accepted witnesses that:
  • The state $s_{a_{1}}^{\mathtt{in}}$ is initial
  • The state $s_{a_{m}}^{\mathtt{out}}$ is final;
  • There is an $\epsilon$-transition between each $s_{a_{i}}^{\mathtt{out}}$ and $s_{a_{i+1}}^{\mathtt{in}}$ for each $1\leq i<m$
Therefore, by construction, we have:
  • $a_{1}\in\Sigma_{\mathrm{start}}$, hence there exists $\alpha^{\prime\prime}\Sigma^{*}$ such that $a_{1}\alpha^{\prime\prime}\in L$;
  • $a_{m}\in\Sigma_{\mathrm{end}}$ hence there exists $\alpha^{\prime\prime\prime}\Sigma^{*}$ such that $\alpha^{\prime\prime\prime}a_{m}\in L$;
  • For each $1\leq i<m$ we have $(a_{i},a_{i+1})\in\Pi$, hence there exist $\alpha_{i}^{\prime},\alpha_{i}^{\prime\prime}\in\Sigma^{*}$ such that $\alpha_{i}^{\prime}a_{i}a_{i+1}\alpha_{i}^{\prime\prime}$ is in $L$.
We now use the fact that $L$ is letter-Cartesian to prove the following claim (*): $\alpha_{1}^{\prime}\alpha\alpha^{\prime\prime}_{m-1}\in L$.
To do this, we show by finite induction that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ for each $1\leq i<m$. The base case of $\alpha^{\prime}_{1}a_{1}a_{2}\alpha^{\prime\prime}_{1}\in L$ is by the definition of $\alpha_{1}^{\prime}$ and $\alpha_{1}^{\prime\prime}$ above. The induction case for $11<i<m-$ 1 is as follows: from $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ (by induction hypothesis) and $\alpha_{i+1}^{\prime}a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ (by definition of $\alpha_{i+1}^{\prime},\alpha_{i+1}^{\prime\prime}$ above), we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ because $L$ is letter-Cartesian on $a_{i+1}$, which is what we wanted to show to establish the induction case. Hence, the inductive claim holds and for $i=m-1$ we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{m}\alpha^{\prime\prime}_{m-1}\in L$, which is claim (*). Now, using that $L$ is letter-Cartesian on $a_{1}$, from (*) and the word $a_{1}\alpha^{\prime\prime}\in L$ identified earlier, we deduce that $\alpha\alpha^{\prime\prime}_{m-1}$ is in $L$. Symmetrically, from this and the word $\alpha^{\prime\prime\prime}a_{m-1}\in L$ identified earlier, using the fact that $L$ is letter-Cartesian on $a_{m-1}$ we deduce that $\alpha\in L$, which concludes. ∎
From the lemma, we can now show the converse direction to Section 3.1.2, giving a self-contained proof of the following known result:
Proposition 3.6 ((sakarovitch2009elements, Chp. II, Ex. 1.8, (d))).
A language is local if, and only if, it is letter-Cartesian.
Proof.
The “only if” direction is shown in Section 3.1.2, so let us focus on the “if” direction. Let then $L$ be a letter-Cartesian language. Let us define directly from $L$ the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$, and $\Pi$ used in the proof of Lemma 3.4. (Note that we do not know yet that $L$ is regular, so we cannot take an $\epsilon$NFA recognizing $L$ and invoke Lemma 3.4 directly; we must instead define the sets directly from $L$. Further note that, in the statement of the present proposition, we do not talk about the complexity of computing these sets, and do not specify how $L$ is represented.)
From the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$, and $\Pi$, let us build the RO-$\epsilon$NFA $A^{\prime}$ as in the proof of Lemma 3.4. One can check that the same reasoning as in the proof of Lemma 3.4 implies that, because $L$ is letter-Cartesian, we have $L=\mathcal{L}(A^{\prime})$. This allows us to conclude by Section 3.1.1 that $L$ is local. ∎
3.2. Proof of Section 3.2: Reduction Preserves Locality
{lemmarep}
[Reduction preserves locality] If $L$ is local then so is $\mathtt{reduce}(L)$.
Proof.
By Proposition 3.6 we have that a language is local iff it is letter-Cartesian. So we show the equivalent claim: if $L$ is letter-Cartesian then so is $\mathtt{reduce}(L)$. Let $\alpha x\beta$ and $\gamma x\delta$ be in $\mathtt{reduce}(L)$, and let us show that $\alpha x\delta\in\mathtt{reduce}(L)$ as well. We have $\alpha x\beta\in L$ and $\gamma x\delta\in L$ and $L$ is letter-Cartesian, so we have $\alpha x\delta\in L$. Assume by contradiction that $\alpha x\delta\notin\mathtt{reduce}(L)$. Then $\alpha x\delta$ has a strict infix, call it $\tau$, that is in $L$. Notice that $\tau$ cannot be an infix of $\alpha$, because $\alpha x\beta\in\mathtt{reduce}(L)$ and $\tau$ would then be a strict infix of $\alpha x\beta$; and similarly $\tau$ cannot be an infix of $\delta$. Therefore $\tau$ must be of the form $\tau=\alpha_{1}x\delta_{1}\in L$ with $\alpha=\alpha_{2}\alpha_{1}$ and $\delta=\delta_{1}\delta_{2}$. We further have one of $\alpha_{2}\neq\epsilon$ and $\delta_{2}\neq\epsilon$, because $\tau$ is a strict infix of $\alpha x\delta$. Assume that $\alpha_{2}\neq\epsilon$ (the case of $\delta_{2}\neq\epsilon$ is symmetrical). Then we have $\alpha_{1}x\delta_{1}\in L$, and $\alpha x\beta\in L$, so since $L$ is letter-Cartesian we have $\alpha_{1}x\beta\in L$. But as $\alpha_{2}\neq\epsilon$ we know that $\alpha_{1}x\beta$ is a strict infix of $\alpha x\beta$, contradicting the fact that $\alpha x\beta\in\mathtt{reduce}(L)$. This concludes the proof. ∎
One related question is whether the criterion of Theorem 3.3 can easily be tested: can we easily test whether a given language $L$ is local? If $L$ is given as a DFA, then the task is in PTIME:
{toappendix}
3.3. Proof of Section 3.3: Locality Testing
{propositionrep}
[Locality testing] Given a DFA $A$, it is in PTIME to determine if $\mathcal{L}(A)$ is local.
Proof.
Given the DFA $A$, we build the RO-$\epsilon$NFA $A^{\prime}$ from Lemma 3.4. By this result and Proposition 3.6, we have $\mathcal{L}(A)\subseteq\mathcal{L}(A^{\prime})$, and $\mathcal{L}(A)$ is local iff $\mathcal{L}(A)=\mathcal{L}(A^{\prime})$. So to check whether $\mathcal{L}(A)$ is local, it is enough to check whether $\mathcal{L}(A^{\prime})\subseteq\mathcal{L}(A)$. For this, we complement $A$ in PTIME to build a DFA $A^{\prime\prime}$ with $\mathcal{L}(A^{\prime\prime})=\Sigma^{*}\setminus\mathcal{L}(A)$, and we must now check whether there is a word of $\mathcal{L}(A^{\prime})$ which is not in $\mathcal{L}(A)$, i.e., which is in $\mathcal{L}(A^{\prime\prime})$. So it suffices to check whether $\mathcal{L}(A)\cap\mathcal{L}(A^{\prime\prime})$ is empty or not. We can do this in PTIME using a product construction to build an $\epsilon$NFA recognizing $\mathcal{L}(A)\cap\mathcal{L}(A^{\prime\prime})$, and determining whether its language is empty or not by a reachability test. This concludes the proof. ∎
Note that Section 3.3 is about testing if the language $L=\mathcal{L}(A)$ recognized by $A$ is local (i.e., some local DFA for $L$ exists); it is not about testing if $A$ is a local DFA (which is easy in linear time).
If the input language $L$ is given as an NFA, we do not know the complexity of testing if $L$ is local: we show in Section 10 that it is PSPACE-hard to test if $L$ satisfies a property related to being local. Another question that remains open is to test whether the reduction of a given language $L$ is local: we can construct an automaton $A^{\prime}$ representing $\mathtt{reduce}(L)$ from an automaton $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, Proposition 6)).
Read-Once automata.
The proof of Theorem 3.3 is not difficult and is a simple reduction to the MinCut problem. To present it, we need a variant of local automata, which we name Read-Once (RO) automata. Intuitively, RO automata strengthen the requirement of local automata by requiring that for each letter $a\in\Sigma$ there is at most one transition labeled with $a$; but in exchange RO automata are not required to be deterministic and allow for (arbitrarily many) $\epsilon$-transitions. Formally:
Definition 3.7 (RO-$\epsilon$NFA).
A read-once $\epsilon$NFA (RO-$\epsilon$NFA for short) is an $\epsilon$NFA $A=(S,I,F,\Delta)$ such that for every letter $a\in\Sigma$ there is at most one transition of the form $(s,a,s^{\prime})$ in $\Delta$.
Example 3.8 (Example 3.2 continued).
The local DFA $A_{1}$ in Figure 2(a) for $ax^{*}b$ is an RO-$\epsilon$NFA, but the local DFA $A_{2}$ in Figure 2(b) for $ab|ad|cd$ is not. An RO-$\epsilon$NFA for this language is $A_{3}$ (Figure 2(c)).
We show that RO-$\epsilon$NFAs recognize precisely the local languages (and the presence of $\epsilon$-transitions is essential for this, see Section 11):
{toappendix}
3.4. Proof of Section 3.4: RO-$\epsilon$NFA = Local Languages
{lemmarep}
[RO-$\epsilon$NFA = local languages] A language $L$ is local iff it is recognized by an RO-$\epsilon$NFA. Moreover, given an $\epsilon$NFA $A$ recognizing a local language $L=\mathcal{L}(A)$, we can build in $O(|A|\times|\Sigma|)$ an RO-$\epsilon$NFA recognizing $L$.
Proof.
The first sentence follows from Section 3.1.1 and Section 3.1.1. The second sentence follows immediately from Proposition 3.6 and Lemma 3.4. ∎
The point of RO-$\epsilon$NFAs is that we can use them to solve resilience with a product construction:
Proof of Theorem 3.3.
Given an $\epsilon$-NFA $A^{\prime}$ for the local language $L$, we compute an RO-$\epsilon$NFA $A=(S,I,F,\Delta)$ for $L$ in $O(|A^{\prime}|\times|\Sigma|)$ using Section 3.4. Let $D$ be the input bag database with fact multiplicities given by $\mathtt{mult}$, and let $V$ be the domain of $D$. Define a network $\mathcal{N}_{D,A}$ as follows:
  • The set of vertices is $(V\times S)\cup\{t_{\text{source}},t_{\text{target}}\}$, with the fresh vertices $t_{\text{source}}$ and $t_{\text{target}}$ being respectively the source and target of the network.
[…]
  • There is an edge with capacity +$\infty$+ ∞ from $(v,s)$ to $(v,s^{\prime})$ for all $v\in V$ and $s,s^{\prime}\in S$ such that $(s,\epsilon,s^{\prime})\in\Delta$.
  • There is an edge with capacity +$\infty$+ ∞ from $t_{\text{source}}$ to every vertex of the form $(v,s)$ with $s\in I$, and from every vertex of the form $(v,s)$ with $s\in F$ to $t_{\text{target}}$.
Clearly $\mathcal{N}_{D,A}$ is built in $O(|A|\times|D|)$ from $A$ and $D$. Moreover, there is a one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of $\mathcal{N}_{D,A}$ that have finite cost: intuitively, selecting facts of $D$ in the contingency set corresponds to cutting edges in $\mathcal{N}_{D,A}$, and ensuring that the remaining facts do not satisfy $Q_{L}$ corresponds to ensuring that no source to target paths remain. Crucially, as $A$ is read-once, the correspondence between facts of $D$ and edges of $\mathcal{N}_{D,A}$ with finite capacity is one-to-one. Further, the sum of multiplicities of the facts of a contingency set is equal to the cost of the corresponding cut, so minimum cuts correspond to minimum contingency sets. We can thus solve MinCut on $\mathcal{N}_{D,A}$ in $\tilde{O}(|\mathcal{N}_{D,A}|)$ (henzinger2024deterministic) to compute the answer to $\mathtt{RES}_{\mathrm{bag}}(L)$ in $D$. ∎
4. Hardness Techniques and First Results
In this section, we move on to cases where computing resilience is intractable. From now on we only study languages that are reduced, because showing hardness for a reduced language $L$ is sufficient to imply the same for any language $L^{\prime}$ such that $\mathtt{reduce}(L^{\prime})=L$.
We start the section by showing the NP-hardness of $\mathtt{RES}_{\mathrm{set}}(aa)$, by an explicit reduction from the minimum vertex cover problem. We then generalize this reduction and define the notion of (hardness) gadgets as databases that can be used to encode graphs and reduce from minimum vertex cover. We show that whenever a language $L$ admits a gadget then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard. We finally illustrate these techniques to show the hardness of resilience for the specific language $axb|cxd$.
The notion of gadgets introduced in this section will be used in the next two sections to show the hardness of entire language families. Specifically, we will generalize in Section 5 the hardness of $\mathtt{RES}_{\mathrm{set}}(axb|cxd)$ to that of the so-called four-legged languages; and generalize in Section 6 the hardness of $\mathtt{RES}_{\mathrm{set}}(aa)$ to that of all finite languages featuring a word with a repeated letter.
4.1. Hardness of $aa$
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20, Prop. 10) by a direct reduction from 3-SAT. We give a self-contained proof where we reduce instead from minimum vertex cover: this proof serves as an intuitive presentation of the more general constructions that come next.
Proposition 4.1.
$\mathtt{RES}_{\mathrm{set}}(aa)$ is NP-hard.
In other words, given a directed graph $G$ and integer $k\in\mathbb{N}$, it is NP-hard to decide if we can remove $k$ edges of $G$ such that $G$ no longer contains any path of length 2. We show Proposition 4.1 by reducing from the minimum vertex cover problem. Recall that a vertex cover of an undirected graph $G=(V,E)$ is a subset of vertices $V^{\prime}\subseteq V$ such that every edge of $E$ is incident to a vertex of $V^{\prime}$, i.e., for each $e\in E$, we have $e\cap V^{\prime}\neq\emptyset$. The vertex cover number of $G$ is the minimum cardinality of a vertex cover of $G$. Given an undirected graph $G$ and an integer $k\in\mathbb{N}$, the minimum vertex cover problem asks if the vertex cover number of $G$ is at most $k$. It is an NP-complete problem (karp1972reductibility).
(a) Completed gadget for $aa$
(b) Graph of $aa$-matches of Figure 3(a)
(c) Completed gadget for $axb|cxd$
(d) Condensed hypergraph of $axb|cxd$-matches of Figure 3(c)
Figure 3. Illustration of hardness gadgets and corresponding condensed hypergraphs of matches. Figures 3(a) and 3(b) illustrates Proposition 4.1, while Figures 3(c) and 3(d) illustrates Proposition 4.8.
Before showing the reduction, we show that the vertex cover number of an undirected graph $G$ can be computed from the vertex cover number of any odd-length subdivision of $G$. Formally, for $\ell\in\mathbb{N}$, $\ell\geq 1$, an $\ell$-subdivision of $G$ is an undirected graph $G^{\prime}$ obtained by replacing each edge of $G$ by a path of length $\ell$ (i.e., with $\ell-1$ fresh internal vertices, that are disjoint across edges). We claim:
{toappendix}
4.2. Proof of Section 4.2: Hardness of Vertex Cover on Subdivisions
{propositionrep}
[Hardness of vertex cover on subdivisions] Let $\ell\in\mathbb{N}$ be an odd integer. Then for any undirected graph $G=(V,E)$, letting $G^{\prime}$ be an $\ell$-subdivision of $G$, then the vertex cover number of $G^{\prime}$ is $k+m(\ell-1)/2$, where $k$ is the vertex cover number of $G$ and $m$ is the number of edges of $G$.
Proof.
It suffices to show that $G$ has a vertex cover of cardinality at most $k$ iff $G^{\prime}$ has a vertex cover of cardinality at most $m(\ell-1)/2$.
We first show the first direction. Let $G=(V,E)$ and let $C$ be a vertex cover of $V$ of cardinality at most $k$. We pick a subset $C^{\prime}$ of the vertices of $G^{\prime}$ by taking the vertices of $C$ and picking vertices of the subdivisions in the following way. For each edge $\{u,v\}$ of $G$, letting $u-v_{1}-\cdots v_{\ell-1}-v$ be the corresponding path in $G$, we distinguish three cases:
  • If $u\in C$ and $v\notin C$ then we add to $C^{\prime}$ the vertices $v_{2},\ldots,v_{\ell-1}$
  • If $v\in C$ and $u\notin C$ then we add to $C^{\prime}$ the vertices $v_{1},\ldots,v_{\ell-2}$.
  • If $u\in C$ and $v\in C$ we do one of the above (it does not matter which).
The case where $u\notin C$ and $v\notin C$ is impossible because $C$ is a vertex cover of $G$.
Now, for each edge of $G$ we select $(\ell-1)/2$ vertices in $C^{\prime}$, so in total $C^{\prime}$ has the prescribed cardinality. To see that it is a vertex cover, note that all edges of $G^{\prime}$ occurs in a path of the form $u-v_{1}-\cdots v_{\ell-1}-v$, and the vertices of $C^{\prime}$ selected above form a vertex cover of the edges of such paths. Hence, $C^{\prime}$ is a vertex cover of $G^{\prime}$.
We next show the second direction, which is somewhat more challenging. Given a vertex cover $C^{\prime}$ of $G^{\prime}$ of cardinality at most $k+m(\ell-1)/2$, we will first modify $C^{\prime}$ without increasing its cardinality to guarantee the following property (*): for each edge $(u,v)\in E$, considering the corresponding path $u-v_{1}-\cdots v_{\ell-1}-v$ of $G^{\prime}$, then there are precisely $(\ell-1)/2$ of the intermediate vertices of the path that are in $C^{\prime}$, formally $C^{\prime}\cap\{v_{1},\ldots,v_{\ell-1}\}=(\ell-1)/2$.
For this, we process each path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, and consider each vertex $v_{1},\ldots,v_{\ell-1}$ in order. If for $i>1$ we have that $v_{i}\in C^{\prime}$ and the previous vertex $v_{i-1}$ was already in $C^{\prime}$, then we remove $v_{i}$ from $C^{\prime}$ and add instead $v_{i+1}$ to $C^{\prime}$. Note that the cardinality of $C^{\prime}$ does not increase, and may strictly decrease if $v_{i+1}$ was already in $C^{\prime}$. At the end of this process, given that we start with a vertex cover of $C^{\prime}$, we see that $C^{\prime}$ contains precisely $(\ell-1)/2$ vertices of the path: if $v_{1}\in C^{\prime}$ then we keep the $v_{i}$’s for which $i$ is odd, and otherwise the $v_{i}$’s for which $i$ is even.
We have processed $C^{\prime}$ to still be a vertex cover, and have cardinality $m(\ell-1)/2+k^{\prime}$ for some $k^{\prime}\leq k$, with $k^{\prime}=C^{\prime}\cap V$. Let us define $C\coloneq C^{\prime}\cap V$. It has cardinality at most $k$, and we claim that $C$ is a vertex cover. Indeed, assuming by contradiction that some edge $(u,v)$ of $G$ is not covered by $C$, so that $u\notin C$ and $v\notin C$, and considering the path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, we know that $C^{\prime}$ (after the processing of the previous paragraph) was a vertex cover of $G^{\prime}$, but if $u\notin C$ and $v\notin C$ then it cannot be the case that precisely $(\ell-1)/2$ of the $v_{i}$’s are in $C^{\prime}$ and cover all the edges of the path, a contradiction.
We have thus shown both directions of the claimed equivalence, which concludes the proof. ∎
Let us now show the hardness of resilience for the language $aa$:
{toappendix}
4.3. Proof of Proposition 4.1
Proof of Proposition 4.1.
{toappendix}
$v$$u$$w$$v$$u$$w$
(a) Undirected graph $G$ (top) and arbitrary orientation $G^{\prime}$ of $G$ (bottom)
(b) Database $\Xi$ constructed to encode $G^{\prime}$ with the gadget whose completion is pictured in Figure 3(a). All edges are labeled by the letter $a$. Letting $e=(u,v),e^{\prime}=(v,w)$ and $e^{\prime\prime}=(w,u)$, we write, e.g., $i^{\prime\prime}$ instead of $t_{e^{\prime\prime},i}$ for brevity.
Figure 4. Illustration of the reduction in the proof of Proposition 4.1.
[…]
The resilience of $aa$ on $\Xi$ is then equal to the vertex cover number of a 5-subdivision [^4] A simpler construction can achieve a 3-subdivision, but our construction will generalize better for subsequent definitions. of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-path. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(b)). Thus, the graph of $aa$-matches of $\Xi$ is (isomorphic to) a 5-subdivision of $G$, and we conclude by Section 4.2. ∎
4.4. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(a) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, as in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd chain. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
Definition 4.2 (Pre-gadget).
[…]
Example 4.3.
[…]
We next explain how we use gadgets to encode directed graphs (see Figure 4 in Section 4.3):
Definition 4.4 (Encoding a graph with a pre-gadget).
Given a directed graph $G$ and a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$, the encoding of $G$ with $\Gamma$ is the database $\Xi$ obtained from $G$ as follows:
  1. (1)
[…]
  2. (2)
    For each edge $e=(u,v)$ in $G$, add to $\Xi$ a copy $D_{e}$ of $D$ defined by the following isomorphism:
    • domain element $t_{\textup{in}}$ is renamed to $t_{u}$ and domain element $t_{\textup{out}}$ is renamed to $t_{v}$
    • all other elements are renamed to fresh elements called the internal elements of $D_{e}$.
Let us observe the following property about the encoding, which is the reason why we require that in-elements and out-elements never occur as the head of facts in pre-gadgets (before completion).
{toappendix}
4.5. Proof of Section 4.5: Paths in Encodings
{claimrep}
[Paths in encodings] Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a pre-gadget, and let $\Xi$ be the encoding of $G$ with $\Gamma$. Then no walk of $\Xi$ passes through two internal elements belonging to different copies of $D$.
Proof.
Consider a walk in $\Xi$ and consider the first element used in the walk which is an internal element of a copy $D_{e}$ of $D$. The only elements shared between $D_{e}$ and the rest of $\Xi$ are the images of the elements $t_{\textup{in}}$ and $t_{\textup{out}}$, i.e., elements of the form $t_{v}$ for $v$ a vertex of $D$. However, by our requirement on $D$, there is no fact in $D$ having $t_{\textup{in}}$ or $t_{\textup{out}}$ as head. Hence, the remainder of the walk can never reach the images of $t_{\textup{in}}$ and $t_{\textup{out}}$, so it can never reach an internal element of another copy. ∎
For now, our definition of pre-gadgets is independent of the query. Let us now formalize what we require from query matches in gadgets, in order to reduce from the vertex cover problem. We do this by introducing the hypergraph of matches, generalizing the graph of matches from the proof of Proposition 4.1, along with a condensation operation to simplify it.
4.6. Condensed Hypergraph of Matches
Let $L$ be a language and $D$ a database. A match (or witness) of $L$ on $D$ is a set of edges corresponding to an $L$-walk (as defined in Section 2): formally, each $L$-walk $e_{1},\ldots,e_{m}$ defines a match which is the set $\{e_{1},\ldots,e_{m}\}$, noting that several $L$-walks may define the same match. Then:
Definition 4.5 (Hypergraph of matches).
The hypergraph of matches of $D$ and $L$, denoted $\mathcal{H}_{L,D}$, is the hypergraph whose nodes are the facts of $D$ and whose hyperedges are the matches of $L$ on $D$.
Recall that a hitting set (also called hypergraph transversal (gottlob2004hypergraph)) of a hypergraph $\mathcal{H}=(V,E)$ is a subset $V^{\prime}\subseteq V$ such that $V^{\prime}\cap e\neq\emptyset$ for every $e\in E$. Thus, hitting sets are a generalization of vertex covers to hypergraphs. It is clear from the definition of resilience that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of the hypergraph of matches $\mathcal{H}_{L,D}$. We now introduce two rewriting rules, called the condensation rules, that can be applied to any hypergraph $\mathcal{H}=(V,E)$ to simplify it without affecting the minimum size of hitting sets. For $v\in V$, write $E(v)=\{e\in E\mid v\in e\}$. The condensation rules are:
Edge-domination.:
If there are $e,e^{\prime}\in E$ with $e\neq e^{\prime}$ and $e\subseteq e^{\prime}$, we can remove $e^{\prime}$ to obtain $\mathcal{H}^{\prime}=(V,E\setminus\{e^{\prime}\})$. Intuitively, hitting sets must intersect $e$ so they automatically intersect $e^{\prime}$.
Node-domination.:
If there are $v,v^{\prime}\in V$ such that $E(v)\subseteq E(v^{\prime})$, we can remove $v$ to obtain $\mathcal{H}^{\prime}=(V\setminus\{v\},\{e\setminus\{v\}\mid e\in E\})$. Intuitively, we can always replace $v$ by $v^{\prime}$ in hitting sets.
We say that $\mathcal{H}^{\prime}$ is a condensed hypergraph of $\mathcal{H}$ if $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by applying some sequence of condensation rules (note that we do not require that $\mathcal{H}^{\prime}$ cannot be further simplified using the rules). One can easily see that the rules preserve the minimum size of a hitting set:
{toappendix}
4.7. Proof of Section 4.7: Condensation Preserves Minimum Size of Hitting Sets
{claimrep}
[Condensation preserves minimum size of hitting sets] Let $\mathcal{H}^{\prime}$ be a condensed hypergraph of $\mathcal{H}$. Then the minimum size of a hitting set of $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are the same.
Proof.
We show the claim by induction on the length on the sequence of applications of the condensation rules: the base case, corresponding to the empty sequence, is trivial.
For the induction case, let us first assume that $\mathcal{H}^{\prime}$ is obtained from $\mathcal{H}=(V,E)$ by applying the edge-domination rule. We then immediately see that any $V^{\prime}\subseteq V$ is a hitting set of $\mathcal{H}$ iff it is a hitting set of $\mathcal{H}^{\prime}$, so the minimal size of a hitting set in $\mathcal{H}$ and $\mathcal{H}^{\prime}$ is the same.
Second, let us assume that $\mathcal{H}^{\prime}=(V^{\prime},E^{\prime})$ is obtained from $\mathcal{H}=(V,E)$ by applying the node-domination rule on two vertices $v$ and $v^{\prime}$ with $E(v)\subseteq E(v^{\prime})$. Write $V^{\prime}=V\setminus\{v\}$ and $E^{\prime}=\{e\setminus\{v\}\mid e\in E\}$. We show both directions of the claim. First, any hitting set of $\mathcal{H}^{\prime}$ is also a hitting set of $\mathcal{H}$, so the minimum size of a hitting set of $\mathcal{H}$ is at most the minimum size of a hitting set of $\mathcal{H}^{\prime}$. Second, let us consider a hitting set $S^{\prime}$ of $\mathcal{H}$. If $v\notin S^{\prime}$ then $S^{\prime}$ is a hitting set of $\mathcal{H}^{\prime}$. Otherwise, if $v\in S^{\prime}$, then $(S^{\prime}\setminus\{v\})\cup\{v^{\prime}\}$ is a hitting set of $\mathcal{H}^{\prime}$ of cardinality at most $|V^{\prime}|$. In both cases $\mathcal{H}^{\prime}$ has a hitting set of cardinality at most $|S^{\prime}|$, so that the minimum size of a hitting set of $\mathcal{H}^{\prime}$ is at most the minimum set of a hitting set of of $\mathcal{H}$. Hence, the equality is proven, which concludes the induction step and finishes the proof. ∎
By Section 4.7, we know that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of any condensed hypergraph of $\mathcal{H}_{L,D}$, so we can always apply condensation rules in hardness proofs.
We can now define the requirement for a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ to be a gadget [^5] Note that a notion generalizing gadgets has been introduced under the name of Independent Join Paths in (makhija2024unified), where it is used to show the hardness of resilience of conjunctive queries.: we require that the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on the completion $D^{\prime}$ of $\Gamma$ can be condensed into an odd path. Formally, let $\mathcal{H}=(V,E)$ be a hypergraph and $u,v\in V$. We call $\mathcal{H}$ an odd path from $u$ to $v$ if all hyperedges of $\mathcal{H}$ have size 2 and the corresponding graph is an odd path from $u$ to $v$; i.e., there is a sequence $u=w_{1},\ldots,w_{2k}=v$ such that $E=\{\{w_{i},w_{i+1}\}\mid 1\leq i<2k\}$.
Definition 4.6 (Gadget).
A gadget for language $L$ is a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ that satisfies the following condition: letting $D^{\prime}$ be the completion of $\Gamma$ (see Definition 4.2) where we add two $a$-facts $F_{\textup{in}}$ and $F_{\textup{out}}$ from fresh elements to $t_{\textup{in}}$ and $t_{\textup{out}}$ respectively, there exists a condensation of the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on $D^{\prime}$ which is an odd path from $F_{\textup{in}}$ to $F_{\textup{out}}$.
Example 4.7.
The hardness pre-gadget whose completion is pictured in Figure 3(a) is a gadget for $aa$, as can be seen on the hypergraph of matches of Figure 3(b) (with no condensation rules applied).
We claim that the existence of a gadget implies that the resilience problem is hard, namely:
{toappendix}
4.8. Proof of Section 4.8: Gadgets Imply Hardness
{propositionrep}
[Gadgets imply hardness] Let $L$ be a reduced language. If there is a gadget $\Gamma$ for $L$, then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
The proof follows the same idea as that of Proposition 4.1. We reduce from the vertex cover problem: let $G^{\prime}$ be an undirected graph and $G$ be a directed graph obtained by choosing an arbitrary orientation for the edges of $G^{\prime}$. We build in linear time the encoding $\Xi$ of $G$ by $\Gamma$. We use the oracle to $\mathtt{RES}_{\mathrm{set}}(L)$ to obtain $r\coloneq\mathtt{RES}_{\mathrm{set}}(Q_{L},\Xi)$, which is equal to the minimum size of a hitting set of the hypergraph $\mathcal{H}_{L,\Xi}$ of matches of $L$ on $\Xi$. Using Section 4.5 and the fact that $\Gamma$ is a gadget, we show that there exists a condensation $\mathcal{H}^{\prime}$ of $\mathcal{H}_{L,\Xi}$ that has only hyperedges of size 2 and which (when seen as an undirected graph) is an odd-length subdivision $G^{\prime\prime}$ of $G^{\prime}$. By Section 4.7, $r=\mathtt{RES}_{\mathrm{set}}(Q_{L},\Xi)$ is equal to the minimum size of a hitting set of $\mathcal{H}^{\prime}$, i.e., $r$ is exactly the vertex cover number of $G^{\prime\prime}$, from which we recover the vertex cover number of $G^{\prime}$ via Section 4.2.
Hence, from now on, all hardness results will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo) to automatically check some correctness properties on our gadgets.
{toappendix}
To prove this, we reduce from the vertex cover problem. Let $G^{\prime}$ be an undirected graph, and let $G$ be an arbitrary orientation of $G$. We use the gadget to encode the input graph. The definitions that we impose on gadgets are designed to ensure that the encoding of $G$ has a condensed hypergraph of matches $\mathcal{H}^{\prime}$ which is an odd-length subdivision of the original graph $G^{\prime}$, so that by Section 4.2 computing a minimum vertex cover on $G^{\prime}$ reduces to computing a minimal hitting set of $\mathcal{H}^{\prime}$.
An odd subdivision of an undirected graph is an $\ell$-subdivision for some odd $\ell\in\mathbb{N}$. The technical claim used in the proof of Section 4.8 is the following:
{claim}
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a gadget for $L$, and let $\Xi$ be the encoding of $G$ using $\Gamma$. Then $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph that has only hyperedges of size $2$ and which is (when seen as an undirected graph) an odd subdivision of $G^{\prime}$, the undirected graph corresponding to $G$.
Proof.
Thanks to Section 4.5, one can check that the hypergraph of matches $\mathcal{H}_{L,\Xi}$ is precisely the union of the hypergraphs of matches of the various copies $D_{e}$ of $D$, but where in- and out-elements have been merged as in Definition 4.4. Then, considering the sequence of condensation rules to be done to obtain the condensed hypergraph of matches of the completion $D^{\prime}$ of $D$ that is an odd path as in Definition 4.6, we can do this sequence of simplifications on every copy $D_{e}$ of the gadget $D$ (noticing that by Definition 4.6 the source and target facts are not affected by the condensation rules), thus indeed obtaining a condensed hypergraph of matches of $\mathcal{H}_{L,\Xi}$ that is an odd subdivision of $G^{\prime}$. ∎
We can now prove Section 4.8 from Section 4.8:
Proof of Section 4.8.
We reduce from the vertex cover problem. Let $\Gamma$ be a gadget for $L$. Let $G^{\prime}$ be an undirected graph, and pick an arbitrary orientation $G$ of $G^{\prime}$. Construct the encoding $\Xi$ of $G$ by $\Gamma$: clearly this can be done in linear time. Use the oracle to compute the resilience of $Q_{L}$ on $\Xi$, written $r$, which is equal to the minimum size of a hitting set in $\mathcal{H}_{L,\Xi}$. By Section 4.8, $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph $\mathcal{H}^{\prime}$ that is an odd subdivision of $G^{\prime}$ and moreover by Section 4.7 we have that the size of a minimum hitting set of $\mathcal{H}^{\prime}$ is also equal to $r$. We can then conclude using Section 4.2 (noticing that the length of the odd subdivision is a constant that depends only on $L$). ∎
4.9. Hardness for $axb|cxd$
We illustrate the use of gadgets by building one for the language $axb|cxd$:
Proposition 4.8.
$\mathtt{RES}_{\mathrm{set}}(axb|cxd)$ is NP-hard.
Proof.
{toappendix}
4.10. Hypergraph of Matches for $axb|cxd$ (Proposition 4.8)
[…]
Figure 5. The hypergraph of matches of $axb|cxd$ on the completed gadget in Figure 3(c)
We use the gadget $\Gamma$ for the language $L=axb|cxd$ whose completion $D^{\prime}$ is depicted in Figure 3(c). Indeed, $\Gamma$ is clearly a pre-gadget. Further, we explain in Section 4.10 why the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on $D^{\prime}$ (shown in Figure 5 in Section 4.10) can be condensed to the hypergraph of Figure 3(d), which is an odd path. Thus, $\Gamma$ is a gadget so Section 4.8 concludes. ∎
5. Hardness of Four-Legged Languages
In the previous section, we have presented hardness gadgets, and shown the hardness of resilience for the two specific languages $aa$ and $axb|cxd$. In this section we generalize the latter result to a subclass of the reduced non-local languages, called the four-legged languages: these languages intuitively feature a specific kind of counterexample to locality. We show that every four-legged language admits a gadget, so that resilience for such languages is always NP-hard by Section 4.8. We give two consequences of this result: it implies that resilience is hard for all non-star-free regular languages; and it also implies a dichotomy on resilience for languages featuring a neutral letter.
5.1. Four-Legged Languages
Recall the notion of local languages (Definition 3.1) from Section 3, which hinges on recognizability by certain automata (namely, local DFAs). We now give a known equivalent language-theoretic characterization of local languages, that we call being letter-Cartesian:
Definition 5.1 (Letter-Cartesian (sakarovitch2009elements)).
A language $L$ is letter-Cartesian if the following implication holds: for every letter $x\in\Sigma$ and words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$, then $\alpha x\delta\in L$.
Note that, by exchanging the roles of $\alpha,\beta$ and $\gamma,\delta$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$ then also $\gamma x\beta\in L$. Hence the name letter-Cartesian: for each letter $x\in\Sigma$, the words of $L$ containing an $x$ can be expressed as a Cartesian product between what can appear left of an $x$ and what can appear right of an $x$. Formally, we have $L=\bigcup_{x\in\Sigma}L_{x}xR_{x}$ for some $L_{x},R_{x}\subseteq\Sigma^{*}$. It is known that letter-Cartesian languages are precisely the local languages: see (sakarovitch2009elements, Chp. II, Ex. 1.8, (d)) for a statement without proof, or Proposition 3.6 in Section 3.1 for a self-contained proof.
Example 5.2.
The local languages of Example 3.8 are letter-Cartesian. Take now $L=aa$. With $x\coloneq a$, $\alpha\coloneq a$, $\beta\coloneq\epsilon$, $\gamma\coloneq\epsilon$, $\delta\coloneq a$, we see that $L$ is not letter-Cartesian (as $aaa\notin L$), hence it is not local.
If a language $L$ is not letter-Cartesian, by definition there are $x\in\Sigma$ and $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We then define the four-legged languages as those reduced languages that feature such a violation in which none of the words $\alpha$, $\beta$, $\gamma$, and $\delta$ is empty. Formally:
Definition 5.3 (Four-legged).
A language $L$ is four-legged if it is reduced and has a letter $x\in\Sigma$ (the body) and four non-empty words $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ (the legs) with $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$.
Example 5.4.
The languages $axb|cxd$ (from Proposition 4.8) and $axb|cxd|cxb$ are non-local and four-legged. The languages $aa$ (from Proposition 4.1) and $ab|bc$ are non-local but not four-legged.
Note that four-legged languages are never local, because they are not letter-Cartesian. The main result of this section is that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard when $L$ is four-legged; formally:
{toappendix}
5.2. Proof of Section 5.2: Hardness of Four-Legged Languages
{theoremrep}
[Hardness of four-legged languages] If $L$ is four-legged then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
Let $L$ be a four-legged language, with body $x\in\Sigma$ and legs $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ so that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We first show that, up to changing the legs, we can ensure that no infix of $\alpha x\delta$ is in $L$. Then we distinguish two cases, depending on whether some infixes of $\gamma x\beta$ are in $L$ or not, and exhibit a gadget for each case. See Section 5.2 for details, with the gadgets and condensed hypergraphs of matches pictured in Figures 6 and 7.
{toappendix}
To show Section 5.2, we first show an intermediate lemma: for reduced languages that are four-legged, we can assume without loss of generality that the legs are stable, i.e., no infix of $\alpha x\delta$ is in $L$:
Definition 5.5.
Let $L$ be a four-legged language, and let $x\in\Sigma$ be a body and let $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ be legs that witness that $L$ is four-legged, i.e., $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We say that the legs $\alpha,\beta,\gamma,\delta$ are stable if, in addition, there is no infix of $\alpha x\delta$ which is in $L$.
{lemmarep}
Any reduced language which is four-legged with body $x\in\Sigma$ also has stable legs with body $x$.
Proof.
Let $L$ be a reduced language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $\eta^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $\eta^{\prime}$ is in $L$) then we are done. Otherwise, as $\eta^{\prime}\notin L$, it means that there is a strict infix $\eta$ of $\eta^{\prime}$ that is in $L$. We know that $\eta$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is reduced. Similarly, $\eta$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $\eta=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $\eta=\eta^{\prime}$ but we assumed that $\eta$ is a strict infix of $\eta^{\prime}$.
We distinguish two cases:
  • We have $\delta_{2}\neq\epsilon$. Then we claim that $\alpha\coloneq\gamma^{\prime}$, $\beta\coloneq\delta^{\prime}$, $\gamma\coloneq\alpha_{1}$ and $\delta\coloneq\delta_{1}$ are stable legs. Indeed, observe first that none of $\alpha,\beta,\gamma,\delta$ is empty. Moreover we have $\alpha x\beta=\gamma^{\prime}x\delta^{\prime}\in L$ and $\gamma x\delta=\alpha_{1}x\delta_{1}\in L$. Last, no infix of $\alpha x\delta=\gamma^{\prime}x\delta_{1}$ can be in $L$ because $\gamma^{\prime}x\delta_{1}\delta_{2}=\gamma^{\prime}x\delta^{\prime}$ is in $L$ and $\delta_{2}\neq\epsilon$ so any infix of $\alpha x\delta$ is a strict infix of a word of $L$; but $L$ is reduced so no strict infix of a word of $L$ is a word of $L$, hence no infix of $\alpha x\delta$ is a word of $L$.
  • We have $\alpha_{2}\neq\epsilon$. Then we do the symmetric choice of taking $\alpha\coloneq\alpha_{1},\beta\coloneq\delta_{1}$, $\gamma\coloneq\alpha^{\prime}$, and $\delta\coloneq\beta^{\prime}$ and claim that they are stable legs. Instead, none of $\alpha,\beta,\gamma,\delta$ is empty. Further, $\alpha x\beta=\alpha_{1}a\delta_{1}$ and $\gamma x\delta=\alpha^{\prime}x\beta^{\prime}$ are in $L$. Moreover we have $\alpha x\delta=\alpha_{1}x\beta^{\prime}$ and all its infixes are infixes of $\alpha^{\prime}x\beta^{\prime}$ which are strict because $\alpha_{2}\neq\epsilon$; as $\alpha^{\prime}x\beta^{\prime}\in L$ and $L$ is reduced, no infix of $\alpha x\delta$ is a word of $L$.
Hence, in either case, we can define stable legs for $L$ with body $x$, which concludes the proof. ∎
For completeness, we point out that Section 5.2 does not hold if $L$ is not reduced:
{remark}
Take $L=\{x,axb,cxd\}$, which is four-legged, with the only possible body being $x$ and legs $(\alpha,\beta,\gamma,\delta)=(a,b,c,d)$; but these legs are not stable because $x$ is a infix of $\alpha x\delta=axd$ that is in the language.
Figure 6. Gadget for Case 1 of Section 5.2.
Figure 7. Gadget for Case 2 of Section 5.2.
We can now prove Section 5.2:
Proof of Section 5.2.
Let $L$ be a reduced and four-legged language. Using Section 5.2, let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be four stable legs, i.e., we have that $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but no infix of $\alpha^{\prime}x\delta^{\prime}$ is in $L$. We then distinguish two cases:
Case 1.:
The first case is when, furthermore, no infix of $\gamma^{\prime}x\beta^{\prime}$ is in $L$. Then in this case, let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$, which is well defined since $\alpha^{\prime}\neq\epsilon$) and likewise write $\beta^{\prime}=\beta b$ and $\gamma^{\prime}=c\gamma$ and $\delta^{\prime}=\delta d$ (note that some of the letters $a$, $b$, $c$, and $d$ may be equal). In this case we claim that the gadget depicted in Figure 6 is a valid gadget. Edges in the gadgets labelled with words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ represent chains labelled with the letters in the words. If any of the words is empty, then the head node of the edge is merged with the tail node. Such a drawing allows us to show hardness of any language that can be obtained by replacing the $\alpha,\beta,\gamma,\delta$ with any words in $\Sigma^{*}$. This is possible since, for any language that meets the conditions of case 1, a condensed hypergraph of matches can be obtained that does not contain any vertices corresponding to a word (i.e., a Greek symbols). We check this programmatically by exhaustively verifying that all directed paths in the completed gadget are either accounted for by the condensed hypergraph of matches or are not in the language. Thus, we know that for any such language, a condensed hypergraph of matches will be an odd chain as shown in Figure 6, and hence the gadget is a valid gadget.
Case 2.:
The second case is when there are some infixes of $\gamma^{\prime}x\beta^{\prime}$ that are in $L$. Observe that such infixes must intersect non-trivially both $\gamma^{\prime}$ and $\beta^{\prime}$. Indeed, infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\gamma^{\prime}$ are strict infixes of $\alpha^{\prime}x\beta^{\prime}\in L$ so they are not in $L$ because $L$ is reduced, and likewise infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\beta^{\prime}$ are strict infixes of $\gamma^{\prime}x\delta^{\prime}\in L$ so again they are not in $L$. Let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$) and $\beta^{\prime}=b\beta$ ($b$ is the first letter of $\beta^{\prime}$) and $\gamma^{\prime}=c_{1}\gamma_{1}$ ($c_{1}$ is the first letter of $\gamma^{\prime}$) and $\gamma^{\prime}=\gamma_{2}c_{2}$ ($c_{2}$ is the last letter of $\gamma^{\prime}$) and $\delta^{\prime}=\delta d$ ($d$ is the last letter of $\delta^{\prime}$). Then we claim that the gadget depicted in Figure 7 is a valid gadget. Indeed, we can programmatically verify this as well by exhaustively checking all directed paths. We observe that all paths (1) either contain or are subsumed by $\alpha^{\prime}x\beta^{\prime}$ or $\gamma^{\prime}x\delta^{\prime}$ matches, (2) are an infix of $\alpha^{\prime}x\delta^{\prime}$ (and hence can never be a match) or (3) contain $c_{2}xb$ as an infix. We notice that any infix of $\gamma^{\prime}a\beta^{\prime}$ that is a match must contain $c_{2}xb$ as an infix, and we know that such a match exists. For any such match, the other letters of the match can be eliminated in the condensed hypergraph of matches, as they are subsumed (via the edge domination rule) by the $c_{2}xb$ thus the node domination rule can remove the other letters of the $\gamma^{\prime}a\beta^{\prime}$-infix match. Thus, it suffices to consider $\alpha^{\prime}x\beta^{\prime}$, $\gamma^{\prime}x\delta^{\prime}$, and $c_{2}xb$ as the only matches in the condensed hypergraph of matches. When we do so, we can see that there exists a condensed hypergraph of matches that is an odd chain as shown in Figure 7, and hence the gadget is valid.
This concludes the proof of Section 5.2. ∎
Note that Section 5.2 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not in NP). Next, we spell out two consequences of the result.
5.3. Implications of Hardness for Four-Legged Languages
{toappendix}
5.4. Reduced Languages of Star-Free Languages are Star-Free
We formally show the following result stated in the main text:
{claim}
Let $L$ be a star-free regular language. Then $\mathtt{reduce}(L)$ is star-free.
Of course, the converse statement is not true, e.g., $(aa)^{*}$ is not star-free but its reduced language $\epsilon$ is star-free. Let us prove the result:
Proof of Section 5.4.
We use an alternative characterization of star-free languages (mpri): they are those languages that can be constructed inductively by applying the operations of concatenation and the Boolean operations (including complementation) on base languages consisting of the empty language, the singleton language containing only the empty word, and the singleton languages formed of single-letter words for each letter of the alphabet $\Sigma$.
We then use the fact that, for any language $L$, we can express $\mathtt{reduce}(L)$ as follows:
$\mathtt{reduce}(L)=L\setminus(\Sigma^{+}L\Sigma^{*}\cup\Sigma^{*}L\Sigma^{+})$
Note that $\Sigma^{*}$ is the complement of the empty set, and $\Sigma^{+}$ is the concatenation of $\Sigma^{*}$ with the union of all letters of $\Sigma$, so both are star-free languages. Hence, if $L$ is a star-free language, then thanks to the fact that star-free languages are closed under concatenation and Boolean operations, so is $\mathtt{reduce}(L)$, which concludes the proof. ∎
5.5. Proof of Section 5.5: Hardness of Resilience for Non-Star-Free Languages
Hardness for non-star-free languages.
We say that a regular language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri)), and if a language $L$ is star-free then so is $\mathtt{reduce}(L)$ (see Section 5.4). We can show that being star-free is a necessary condition to enjoy tractable resilience, because non-star-free reduced regular languages are four-legged and hence resilience for them is NP-hard by Section 5.2:
{lemmarep}
[Hardness of non-star-free languages] Let $L$ be a regular language which is reduced and non-star-free. Then $L$ is four-legged.
Proof.
We use the equivalent characterization of star-free languages as those languages that are counter-free (mcnaughton1971counter). Let $A$ be a DFA which recognizes $L$, and let $k^{\prime}>0$ be greater than the number of states of $A$. Let $\rho,\sigma,\tau\in\Sigma^{*}$ and $m\geq k$ be witnesses to the fact that $L$ is not counter-free for $k=k^{\prime}$, i.e., exactly one of $\rho\sigma^{k}\tau$ and $\rho\sigma^{m}\tau$ is in the language. In particular this implies that $\sigma\neq\epsilon$; let us write $\sigma=x\sigma^{\prime}$.
By the pigeonhole principle, there must be $0\leq i<j\leq k^{\prime}$ such that the partial runs of $\rho\sigma^{i}$ and $\rho\sigma^{j}$ finish at the same state when read by $A$. Let $\omega\coloneq j-i$. We know that we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+\omega}\tau\in L$, because exchanging the prefixes $\rho\sigma^{i}$ and $\rho\sigma^{j}$ does not change the state that we reach. Iterating the argument, we get the following fact, dubbed (*): for any $p\geq 0$ we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+p\omega}\tau\in L$. For similar reasons, we get the fact dubbed (**): for any $p\geq 0$ we have $\rho\sigma^{m}\tau\in L$ iff $\rho\sigma^{m+p\omega}\tau\in L$.
Let $(\phi,\psi)=(k,m)$ if $\rho\sigma^{k}\tau\in L$ and $\rho\sigma^{m}\tau\notin L$, and $(\phi,\psi)=(m,k)$ if it is the opposite. Recall that $m,k>0$ so $\phi>0$ and $\psi>0$. Further, by (*) and (**) we know the following fact, dubbed ($\dagger$): we have $\rho\sigma^{\phi+p\omega}\tau\in L$ for all $p\geq 0$, and $\rho\sigma^{\psi+p\omega}\tau\notin L$ for all $p\geq 0$. Let $p$ be large enough so that $\phi+p\omega-1>\psi$. Now let us take:
  • $\alpha=\rho\sigma^{2\omega-1}$,
  • $\beta=\sigma^{\prime}\sigma^{\phi}w$,
  • $\gamma=\rho\sigma^{\phi+p\omega-1-\psi}$ and
  • $\delta=\sigma^{\prime}\sigma^{\psi}w$.
Note that all of $\alpha,\beta,\gamma,\delta$ are non-empty because they all contain $v\neq\epsilon$ at a power which is greater than 0. We claim that they witness that $L$ is four-legged with body $x$. Indeed:
  • $\alpha x\beta=\rho\sigma^{2\omega-1}x\sigma^{\prime}\sigma^{\phi}\tau=\rho\sigma^{2\omega+\phi}\tau$
  • $\gamma x\delta=\rho\sigma^{\phi+p\omega-1-\psi}x\sigma^{\prime}\sigma^{\psi}\tau=\rho v^{\phi+p\omega}\tau$
  • $\alpha x\delta=\rho\sigma^{2\omega-1}x\sigma^{\prime}\sigma^{\psi}\tau=\rho\sigma^{2\omega+\psi}\tau$.
So by ($\dagger$) we have that:
  • $\alpha x\beta\in L$ because $\rho\sigma^{\phi}\tau\in L$
  • $\gamma x\delta\in L$ because $\rho\sigma^{\phi}\tau\in L$
  • $\alpha x\delta\in L$ because $\rho\sigma^{\psi}\tau\notin L$
This establishes that $L$ is four-legged. As it is reduced, we can conclude by Proposition 5.2. ∎
{toappendix}
5.6. Proof of Section 5.6: Dichotomy for Languages with Neutral Letters
Dichotomy for languages with a neutral letter.
A second consequence of Section 5.2 concerns languages $L$ with a neutral letter, i.e., a letter $e\in\Sigma$ such that inserting or deleting $e$ anywhere does not change the membership of words to $L$. Formally, $e$ is neutral for $L$ if for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\beta\in L$ iff $\alpha e\beta\in L$. In formal language theory, one often assumes the existence of a neutral letter as a technical assumption on languages (koucky2005bounded; barrington2005first). Under this assumption, we can show:
{propositionrep}
[Dichotomy for languages with neutral letters] Let $L$ be a language with a neutral letter. If $\mathtt{reduce}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME, otherwise $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
{proofsketch}
If $\mathtt{reduce}(L)$ is local then we have tractability thanks to Theorem 3.3. Otherwise, we can show that $\mathtt{reduce}(L)$ is either four-legged or contains a word of the form $aa$ (or both), so that at least one of Proposition 4.1 and Section 5.2 applies.
{toappendix}
To prove this result, we first show the claim mentioned in the proof sketch: for languages featuring a neutral letter, whenever the reduced language is non-local then it must either be four-legged or contain a word with a repeated letter, specifically a word of the form $aa$ for $a\in\Sigma$:
{lemmarep}
Let $L$ be a language with a neutral letter such that $\mathtt{reduce}(L)$ is not local. Then one of the following two cases holds (possibly both):
  • $\mathtt{reduce}(L)$ is four-legged;
  • there is $x\in\Sigma$ such that $xx\in\mathtt{reduce}(L)$.
To prove this result, let us prove the following intermediate result:
{claim}
Let $L$ be a language with a neutral letter $e$ and let $L^{\prime}=\mathtt{reduce}(L)$. Then for any nonempty words $\theta,\eta\in\Sigma^{+}$, if $\theta\eta\in L^{\prime}$ then $\theta e\eta\in L^{\prime}$.
Proof.
We have $\theta\eta\in L^{\prime}$ so $\theta\eta\in L$, and $e$ is neutral for $L$ so we have $\theta e\eta\in L$. Let us proceed by contradiction and assume that $\theta e\eta\notin L^{\prime}$. This means that there is a strict infix $\tau$ of $\theta e\eta$ which is in $L$. We first claim that $\tau$ must include the $e$. Indeed, if it does not, then there are two cases:
  • $\tau$ is an infix of $\theta$. Then it is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the fact that $L^{\prime}$ is reduced, so this case is impossible.
  • $\tau$ is an infix of $\eta$. This is impossible for the same reason as the previous case.
Now, we know that $\tau$ is a strict infix of $\theta e\eta$ which includes the $e$. Let $\tau^{\prime}$ be obtained from $\tau$ by removing the $e$. We have $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$. What is more, it is easily seen that $\tau^{\prime}$ is an infix of $\theta\eta$, and it is a strict infix because $\tau$ includes the $e$ so from $|\tau|<|\theta e\eta|$ (because $\tau$ is a strict infix) we deduce $|\tau^{\prime}|=|\tau|-1<|\theta e\eta|-1=|\theta\eta|$. So $\tau^{\prime}\in L$ is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the assumption that $L^{\prime}$ is reduced. ∎
We can now prove Section 5.6:
Proof of Section 5.6.
[…]
Now, because $L^{\prime}$ is reduced, it is easy to check that we must have the following claim, dubbed ($\star$): we have $\alpha^{\prime}\beta^{\prime}\neq\epsilon$ and $\gamma^{\prime}\delta^{\prime}\neq\epsilon$ and $\alpha^{\prime}\gamma^{\prime}\neq\epsilon$ and $\beta^{\prime}\delta^{\prime}\neq\epsilon$. Indeed:
  • If $\alpha^{\prime}\beta^{\prime}=\epsilon$ then $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ is equal to $x$ which is an infix of $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$, so by reducedness it must be the case that $\gamma^{\prime}\delta^{\prime}=\epsilon$, but now $x=\alpha^{\prime}x\delta^{\prime}=x$ and we assumed $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$, contradiction.
  • If $\gamma^{\prime}\delta^{\prime}=\epsilon$, the reasoning is similar: in this case $\gamma^{\prime}x\delta^{\prime}$ is equal to $x$, which is an infix of $\alpha^{\prime}x\beta^{\prime}$, so by reducedness we must have $\alpha^{\prime}\beta^{\prime}=\epsilon$, but then $x=\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ and $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ is a contradiction.
Thus, we have established ($\star$).
We now start proving the statement. Let us first show that (call this $\dagger$) if both $\alpha^{\prime}$ and $\delta^{\prime}$ are empty then the result holds. Indeed, in that case we have $x\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}x\in L^{\prime}$. By ($\star$) (first two points) we have that $\beta^{\prime}$ and $\gamma^{\prime}$ are both nonempty. By Section 5.6 we deduce from $x\beta^{\prime}\in L^{\prime}$ that $xe\beta^{\prime}\in L^{\prime}$, and we deduce from $\gamma^{\prime}x\in L^{\prime}$ that $\gamma^{\prime}ex\in L^{\prime}$. Now, take $e$ as a candidate choice of body, and $\alpha=x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x$, as a candidate choice of legs: all of these are non-empty. We have $\alpha e\beta\in L^{\prime}$ and $\gamma e\delta\in L^{\prime}$. So, considering the word $\alpha e\delta=xex$, there are two possibilities: either $xex\notin L^{\prime}$ or $xex\in L^{\prime}$. In the first case, we conclude with our choice of body and legs that $L^{\prime}$ is four-legged, which allows us to conclude. In the second case, we have $xx\in L$ because $e$ is neutral for $L$. Further, we have $xx\in L^{\prime}$ because the strict infixes of $xx$ are the empty word and $x$ which are both strict infixes of the word $x\beta^{\prime}\in L^{\prime}$: we conclude because $L^{\prime}$ is reduced. The fact that $xx\in L^{\prime}$ allows us to conclude: indeed, we can then do the exact same reduction from vertex cover as in Proposition 4.1 (noticing that we do not use other letters than ’$x$’).
The same reasoning shows that (call this $\dagger^{\prime}$) if both $\beta^{\prime}$ and $\gamma^{\prime}$ are empty then the result holds. Indeed, in this case we have $\alpha^{\prime}x\in L^{\prime}$ and $x\delta^{\prime}\in L^{\prime}$, and we have $\alpha^{\prime}$ and $\delta^{\prime}$ both nonempty by ($\star$) (first two points). Applying Section 5.6 to each gives us that $\alpha^{\prime}ex\in L^{\prime}$ and $xe\delta^{\prime}\in L^{\prime}$, and again we conclude that $L^{\prime}$ is four-legged or that $xex\in L^{\prime}$ from which we show $xx\in L^{\prime}$ in the same way.
We have ruled out the case of $\alpha^{\prime}$ and $\delta^{\prime}$ being both empty, and of $\beta^{\prime}$ and $\gamma^{\prime}$ being both empty. Combining with ()($\star$)( ⋆ ), one gets that either $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty, or $\beta^{\prime}$ and $\delta^{\prime}$ are both non-empty. Indeed, assume that the first term in the disjunction is false, i.e., that it is not the case that $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty. If we have $\alpha^{\prime}=\epsilon$, then $\delta^{\prime}$ is non-empty from ($\dagger$), and $\beta^{\prime}$ is non-empty by ()($\star$)( ⋆ ) (first point). If we have $\gamma^{\prime}=\epsilon$, then $\beta^{\prime}$ is non-empty by ($\dagger^{\prime}$), and $\delta^{\prime}$ is non-empty by ()($\star$)( ⋆ ) (second point).
We finish the proof by distinguishing these two cases:
  • Assume that $\alpha^{\prime}\neq\epsilon$ and $\gamma^{\prime}\neq\epsilon$. We have $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ so by Section 5.6 we have $\alpha^{\prime}ex\beta^{\prime}\in L^{\prime}$, and we have $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$ so by Section 5.6 we have $\gamma^{\prime}ex\delta^{\prime}\in L^{\prime}$. We then claim that $\alpha^{\prime}ex\delta^{\prime}\notin L^{\prime}$, implying that $L^{\prime}$ is four-legged (taking $e$ as the body, $\alpha=\alpha^{\prime}$, $\beta=x\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x\delta^{\prime}$, which are all not empty). Indeed, assume by way of contradiction that $\alpha^{\prime}ex\delta^{\prime}\in L^{\prime}$. We would then have $\alpha^{\prime}ex\delta^{\prime}\in L$, so $\alpha^{\prime}x\delta^{\prime}\in L$ (as $e$ is neutral). As $\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ by assumption, there exists a strict infix $\tau$ of $\alpha^{\prime}x\delta^{\prime}$ that is in $L^{\prime}$. Now, adding $e$ to $\tau$ would give a strict infix $\tau^{\prime}$ of $\alpha^{\prime}ex\delta^{\prime}$, with $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$, a contradiction because $L^{\prime}$ is reduced.
  • Assume that $\beta^{\prime}\neq\epsilon$ and $\delta^{\prime}\neq\epsilon$. The reasoning is similar: using Section 5.6 we have $\alpha^{\prime}xe\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}xe\delta^{\prime}\in L^{\prime}$ and we obtain that $L^{\prime}$ is four-legged by taking $e$ as the body, $\alpha=\alpha^{\prime}x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}x$ and $\delta=\delta^{\prime}$.
This concludes the proof. ∎
For completeness we give here examples of languages for which only one of the two cases in the statement of Section 5.6 holds. Let $L_{1}=e^{*}be^{*}ce^{*}|e^{*}de^{*}fe^{*}$ (for which $e$ is neutral), with $\mathtt{reduce}(L_{1})=be^{*}c|de^{*}f$. Then $\mathtt{reduce}(L_{1})$ is four-legged (taking $x=e$, $\alpha=b$, $\beta=c$, $\gamma=d$, $\delta=f$), hence $\mathtt{reduce}(L_{1})$ is not local, but it does not have a word of the form $xx$ for $x\in\Sigma$. Now take $L_{2}=e^{*}(a|c)e^{*}(a|d)e^{*}$, with $\mathtt{reduce}(L_{2})=(a|c)e^{*}(a|d)$. Then $\mathtt{reduce}(L_{2})$ is not local (taking $x=a$, $\alpha=c$, $\beta=\epsilon$, $\gamma=\epsilon$, $\delta=d$), it contains the word $aa$, but one can check that it is not four-legged.
We can then use Section 5.6 to prove Section 5.6:
Proof of Section 5.6.
If $\mathtt{reduce}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME by Theorem 3.3. Otherwise, by Section 5.6 either $\mathtt{reduce}(L)$ is four-legged, in which case $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard by Section 5.2, or $aa\in\mathtt{reduce}(L)$ for $a\in\Sigma$. In this case one can easily adapt the proof of Proposition 4.1: one can check that the reduction only builds instances using the letter $a$, and the restriction of $\mathtt{reduce}(L)$ to words using only the letter $a$ must be exactly $aa$ given that $\mathtt{reduce}(L)$ is reduced. This concludes the proof. ∎
The dichotomy above relies on the presence of a neutral letter, so it does not apply to all languages. In particular, it does not apply at all to finite languages, so we focus on them from now on.
6. Hardness of Finite Languages Having Repeated Letters
We now conclude the presentation of our main hardness results by showing that resilience is hard for a class of languages that generalizes $aa$ (Proposition 4.1). We say that a word $\alpha$ has a repeated letter if we have $\alpha=\beta a\gamma a\delta$ for some $a\in\Sigma$ and $\beta,\gamma,\delta\in\Sigma^{*}$. We show that, for reduced and finite languages, resilience is NP-hard whenever the language contains a word with a repeated letter:
{toappendix}
6.1. Proof of Section 6.1: Hardness with Repeated Letters
{theoremrep}
[Hardness with repeated letters] Let $L$ be a reduced finite language that contains a word with a repeated letter. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-complete.
{proofsketch}
The proof is rather technical. We use the finiteness of $L$ to select a word with a repeated letter of the form $\beta a\gamma a\delta$ where the gap between repeated letters is maximal. A long case distinction then allows us to deduce that either $L$ is four-legged (in which case the result follows from Section 5.2), or we can show that one of several gadgets applies.
{toappendix}
The proof of Section 6.1 is technical, so we sketch it here before giving the details in the rest of the section. We first make a preliminary observation: the tractability of resilience is preserved under the mirror operation on regular languages. We then start from a word featuring a repeated letter, and pick it to optimize certain criteria (maximizing the gap between the repeated letters, and then maximizing the total length of the word): we call this a maximal-gap word, and the existence of such words is the only place where we use the fact that $L$ is finite.
We then consider such a maximal-gap word $\beta a\gamma a\delta$, and consider various cases. We first exclude the case where $\beta$ and $\delta$ are both non-empty, by showing that $L$ is then four-legged. Then, thanks to the mirror operation, we assume without loss of generality that $\delta$ is empty. We then consider the word $\gamma a\gamma$ and study whether it has infixes in $L$. If it does not, we easily build a gadget. If it does, we can now show that $\beta$ must be empty or otherwise the language is four-legged. We then study the choice of infix $\alpha^{\prime}$ of $\gamma a\gamma$ and distinguish two cases by considering the suffix of the left $\gamma$ and the prefix of the right $\gamma$: either they overlap or they don’t. We build suitable gadgets for each case.
The mirror operation.
We first make a straightforward observation on the mirror of languages. The mirror of a word $\alpha=a_{1}\cdots a_{n}\in\Sigma^{*}$ is $\alpha^{\mathrm{R}}=a_{n}\cdots a_{1}\in\Sigma^{*}$ (in particular $\epsilon^{\mathrm{R}}=\epsilon$). The mirror of a language $L\subseteq\Sigma^{*}$ is $L^{\mathrm{R}}\coloneq\{\alpha^{\mathrm{R}}\mid\alpha\in L\}$. Observe that $(L^{\mathrm{R}})^{\mathrm{R}}=L$. It is immediate by symmetry that a language $L$ is reduced iff $L^{\mathrm{R}}$ is reduced. Further, it is also immediate that a language and its mirror have the same complexity for resilience:
{propositionrep}
Let $L$ be any language. There are PTIME reductions between $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$ and $\mathtt{RES}_{\mathrm{set}}(L)$, and between $\mathtt{RES}_{\mathrm{bag}}(L^{\mathrm{R}})$ and $\mathtt{RES}_{\mathrm{bag}}(L)$.
The point of this claim will be to eliminate symmetric cases, up to replacing the language by its mirror. We only use the claim about set semantics in the sequel (given that we will show hardness of resilience in set semantics, from which hardness for bag semantics follows): the claim for bag semantics in Section 6.1 is given just for completeness.
Proof of Section 6.1.
We give the proof for set semantics, the proof for bag semantics is the same mutatis mutandis. Given a database instance $D$ for $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$, simply create a database instance $D^{\mathrm{R}}$ by reversing the direction of all arrows of $D$. It is then obvious that a subset $D^{\prime}\subseteq D$ satisfies $Q_{L^{\mathrm{R}}}$ iff the corresponding subset $(D^{\prime})^{\mathrm{R}}$ of $D^{\mathrm{R}}$, which has the same cardinality, satisfies $Q_{L}$. Hence, the answer to $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$ on $D$ and to $\mathtt{RES}_{\mathrm{set}}(L)$ on $D^{\mathrm{R}}$ are identical, giving the forward PTIME reduction. The converse reduction is obtained by applying the previous argument to reduce $(L^{\mathrm{R}})^{\mathrm{R}}=L$ to $L^{\mathrm{R}}$. ∎
Maximal-gap words
Let $L$ be a reduced finite language that contains a word with a repeated letter, and let us show that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard, to establish Section 6.1. We will pick a word with a repeated letter for the rest of the proof, but we will choose such a word carefully. Here is the precise notion we will need:
Definition 6.1.
For $L$ a finite language, we say that a word $\alpha=\beta a\gamma a\delta$ that features a repeated letter is a maximal-gap word if:
  • The gap between the repeated letters is maximal, i.e., there is no word $\alpha^{\prime}=\beta^{\prime}a^{\prime}\gamma^{\prime}a^{\prime}\delta^{\prime}$ in $L$ such that $|\gamma^{\prime}|>|\gamma|$
  • Among such words, the length of $\alpha$ is maximal, i.e., there is no word $\alpha^{\prime}=\beta^{\prime}a^{\prime}\gamma^{\prime}a^{\prime}\delta^{\prime}$ in $L$ such that $|\gamma^{\prime}|=|\gamma|$ and $|\alpha^{\prime}|>|\alpha|$.
It is clear that any finite language that contains a word with a repeated letter must have maximal-gap words: this is the only place where we use the finiteness of the language. Hence, in the rest of the proof, we fix $\alpha=\beta a\gamma a\delta$ to be an arbitrary such word. We will do case disjunctions on this word in the rest of the proof.
Reducing to the case $\beta=\epsilon$.
Let us first assume that $\beta$ and $\delta$ are both non-empty. In this case we conclude because $L$ is four-legged. Indeed:
{claimrep}
Let $L^{\prime}$ be a reduced language containing a maximal-gap word $\beta a\gamma a\delta$ such that $\beta\neq\epsilon$ and $\delta\neq\epsilon$. Then $L^{\prime}$ is four-legged.
Proof.
Consider the words $\alpha=(\beta a\gamma)a\delta$ and $\alpha=\beta a(\gamma a\delta)$. Note that $\beta a\gamma$, $\delta$, $\beta$, and $\gamma a\delta$ are all nonempty, so they can be used as legs. So consider the word $\alpha^{\prime}=(\beta a\gamma)a(\gamma a\delta)$. The word $\alpha^{\prime}$ contains two $a$’s separated by $\gamma a\gamma$, which is longer than $\gamma$. Hence, $\alpha^{\prime}$ is not in in $L^{\prime}$, as it would contradict the fact that $\alpha$ is maximal-gap (the first criterion of the definition of maximal-gap suffices). So we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
If $\beta$ and $\delta$ are both nonempty, then applying Section 6.1 and Section 5.2 with $L^{\prime}\coloneq L$ concludes. Hence, in the rest of the argument, we assume that one of $\beta$ and $\delta$ is empty. Further, without loss of generality, up to replacing $L$ by $L^{\mathrm{R}}$ using Section 6.1, we assume that $\beta=\epsilon$. Hence, in the rest of the proof, we can assume that $\beta=\epsilon$: we have a maximal-gap word $\alpha=a\gamma a\delta\in L$ and this word is maximal-gap.
Reducing to the case where some infix of $\gamma a\gamma$ is in $L$.
Continuing the proof, we will now consider the word $\gamma a\gamma$ and study whether some infixes of this word are in $L$. (Intuitively, when designing a gadget for $L$, such paths will appear, and so the construction of the gadget will be different depending on whether these paths constitute query matches or not.) Let us first assume that no infix of $\gamma a\gamma$ is in $L$: this case can easily be ruled out by designing a gadget for $L$ based on the word $\alpha=a\gamma a\delta$:
{lemmarep}
Let $L^{\prime}$ be a reduced language containing a word $a\gamma a\delta$ where no infix of $\gamma a\gamma$ is in $L^{\prime}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 8. Gadget for the proof of Section 6.1 in the case when $\delta=\epsilon$ (left), and condensed hypergraph of matches (right).
Figure 9. Gadget for Section 6.1 in the case where $\delta\neq\epsilon$ (left), and condensed hypergraph of matches (right).
Proof.
We first present the case where $\delta=\epsilon$, and then present the case where $\delta\neq\epsilon$.
In the case where $\delta=\epsilon$, we use the completed gadget pictured on Figure 8 (left), where the edge labeled $\gamma$ stands for a (possibly empty) path. Note that, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as their head (except the two facts of the gadget completion, pictured in orange), so indeed it is a pre-gadget. The condensed hypergraph of matches is shown on Figure 8 (right), and it is an odd chain, so indeed the instance pictured on Figure 8 (left) is (the completion of) a gadget and we can conclude by Section 4.8. Note how we use the assumption that the matches on the completed gadget are the ones indicated: directed paths including two $a$’s must include $a\gamma a$ as an infix and so they are dominated, and directed paths including only one $a$ cannot be a match of the query.
In the case where $\delta\neq\epsilon$, we use the (completed) gadget pictured on Figure 9 (left), again with the edge labeled $\gamma$ standing for a possibly empty path. Again, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as head (except for the facts added in the completion), so again we have a pre-gadget. The condensed hypergraph of matches is shown on Figure 9 (right). Now, to show that it is actually like this, first observe that the facts of the $\delta$-paths are always dominated by the preceding $a$: there is no match of the query within $\delta$ (as it is a strict infix of $a\gamma a\delta$), so all matches using $\delta$ must use the preceding $a$, and we can remove the facts of $\delta$-paths in the condensation via the node domination rule. We are then essentially back to the preceding case. More precisely, again, no directed path containing only one copy of the $a$-edges of the gadget can be a word of the language, because these paths are either infixes of $\gamma a\delta$ (hence strict infixes of $a\gamma a\delta\in L$) or they are infixes of $\gamma a\gamma$ and so they are not in the language by assumption. Further, for any two consecutive $a$’s in the gadget, there is a $a\gamma a\delta$ match containing them. Further, after observing that $\delta$ facts are eliminated by node domination, we note that all query matches that are infixes of a word of the form $a(\gamma a)^{*}$ must contain two consecutive $a$’s, so they are subsumed by the matches corresponding to $a\gamma a\delta$, which after node domination contain just two consecutive $a$’s and the $\gamma$ between them. (The facts corresponding to $\gamma$ are then eliminated by node domination.)
Hence, we indeed have a condensed hypergraph of matches corresponding to Figure 9 (right), and we can conclude by Section 4.8. ∎
Applying this lemma with $L^{\prime}\coloneq L$ concludes. So in the rest of the proof we can assume that there is an infix of $\gamma a\gamma$ which is in $L$.
Reducing to the case $\delta=\epsilon$.
The maximal-gap word $\alpha=a\gamma a\gamma\delta$, together with the infix of $\gamma a\gamma$ which is in $L$, now allows us to show that $L$ is four-legged unless $\delta=\epsilon$. To show this, we need to discuss the structure of the infix of $\gamma a\gamma$:
{claim}
Let $L^{\prime}$ be a reduced language containing the word $\alpha=a\gamma a\delta$ and containing an infix of $\gamma a\gamma$. Then $L^{\prime}$ contains a word of the form $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ where $\gamma_{1}$ is a non-empty suffix of $\gamma$ and $\gamma_{2}$ is a non-empty prefix of $\gamma$.
Proof.
Consider the infix of $\gamma a\gamma$ which exists by hypothesis. This infix must include the middle $a$, as otherwise it is an infix of $\gamma$, which is a strict infix of $\alpha\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is reduced.
So write the infix as $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ being a (not necessarily strict) suffix of $\gamma$ and $\gamma_{2}$ being a (not necessarily strict) prefix of $\gamma$. All that remains is to show that $\gamma_{1}$ and $\gamma_{2}$ are non-empty.
First, we must have $\gamma_{1}\neq\epsilon$, otherwise $\alpha^{\prime}$ is an infix of $a\gamma$ which is a strict infix of $a\gamma a\delta\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is reduced.
Second, we must have $\gamma_{2}\neq\epsilon$ otherwise $\alpha^{\prime}$ is an infix of $\gamma a$ which is a strict infix of $a\gamma a\delta$ and we conclude in the same way. ∎
Hence, applying Section 6.1 to $L^{\prime}\coloneq L$, we know that $L$ contains a maximal-gap word $\alpha=a\gamma a\delta$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ and $\gamma_{2}$ being respectively a non-empty suffix of $\gamma$ and a non-empty prefix of $\gamma$. We now claim that, if $\delta\neq\epsilon$, then $L$ is four-legged, allowing us to conclude:
{claimrep}
Let $L^{\prime}$ be a reduced language containing a word $a\gamma a\delta$ such that $\delta\neq\epsilon$ and a word $\gamma_{1}a\gamma_{2}$ such that $\gamma_{1}\neq\epsilon$, $\gamma_{2}\neq\epsilon$, and $\gamma_{1}$ is a suffix of $\gamma$. Then $L^{\prime}$ is four-legged.
Proof.
Indeed, consider $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ and $\alpha=(a\gamma)a\delta$. The words $\delta$, $a\gamma$, $\gamma_{1}$, and $\gamma_{2}$ are all non-empty, so they can be used as legs. So consider the word $\alpha^{\prime\prime}=\gamma_{1}a\delta$ and show that $\alpha^{\prime\prime}\notin L^{\prime}$, so that $L^{\prime}$ is four-legged.
Indeed, the word $\alpha^{\prime\prime}$ is an infix of $\gamma a\delta$ (because $\gamma_{1}$ is a suffix of $\gamma$), which is a strict infix of $a\gamma a\delta$ which is in $L^{\prime}$. Thus, as $L^{\prime}$ is reduced, $\alpha^{\prime\prime}\notin L^{\prime}$, so we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
Applying Section 6.1 with $L^{\prime}\coloneq L$, we can conclude unless $\delta=\epsilon$, which we assume in the rest of the proof.
Distinguishing two cases depending on whether $\gamma_{1}$ and $\gamma_{2}$ overlap.
To summarize, we now know that $L$ contains a maximal-gap word $\alpha=a\gamma a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$ for $\gamma_{1}$ a non-empty suffix of $\gamma$ and $\gamma_{2}$ a non-empty prefix of $\gamma$. We distinguish two cases, intuitively based on whether $\gamma_{1}$ and $\gamma_{2}$ cover $\gamma$ in a way that overlaps or not:
Non-overlapping case.:
The non-overlapping case is when $|\gamma_{1}|+|\gamma_{2}|\leq|\gamma|$, so that we can write $\gamma=\gamma_{2}\eta\gamma_{1}$ for some $\eta\in\Sigma^{*}$.
Overlapping case.:
The overlapping case is when we have $|\gamma_{1}|+|\gamma_{2}|>|\gamma|$, so that we can write $\gamma=\eta^{\prime\prime}\eta\eta^{\prime}$ such that $\gamma_{1}=\eta\eta^{\prime}$ and $\gamma_{2}=\eta^{\prime\prime}\eta$, with the “overlap” $\eta$ being nonempty.
Showing hardness in the overlapping case.
Let us first deal with the overlapping case. We have $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, meaning $\alpha^{\prime}=\eta\eta^{\prime}a\eta^{\prime\prime}\eta\in L$, and we know that the “overlap” $\eta$ is nonempty. We then show:
{claim}
If the words $\eta^{\prime}$ and $\eta^{\prime\prime}$ are not both empty, then $L$ is four-legged.
Proof.
Let us first assume that $\eta^{\prime}$ is nonempty and show that $L$ is four-legged in this case. Let us write $\eta^{\prime}=x\sigma$. Substituting this into $\alpha^{\prime}$ and $\alpha$, we have $\alpha^{\prime}=\eta x(\sigma a\eta^{\prime\prime}\eta)\in L$ and $\alpha=(a\eta^{\prime\prime}\eta)x(\sigma a)\in L$. Now, taking $x$ as body, let us consider the cross-product word $\alpha^{\prime\prime}=\eta x\sigma a=\eta\eta^{\prime}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta\eta^{\prime}a\in L$. Hence, as $L$ is reduced, $\alpha^{\prime\prime}\notin L$, so that we can conclude that $L$ is four-legged unless $\eta^{\prime}$ is empty.
Let us substitute $\eta^{\prime}=\epsilon$ in $\alpha$ and $\alpha^{\prime}$. We know that $\alpha^{\prime}=\eta a\eta^{\prime\prime}\eta\in L$ and $\alpha=a\eta^{\prime\prime}\eta a\in L$. Let us now assume that $\eta^{\prime\prime}$ is nonempty and show that $L$ is four-legged in that case as well. Let us write $\eta^{\prime\prime}=x\sigma$. We have $\alpha=ax(\sigma\eta a)\in L$ and $\alpha^{\prime}=(\eta a)x(\sigma\eta)\in L$. Now, taking $x$ as body again, let us consider the cross-product word $\alpha^{\prime\prime}=ax\sigma\eta=a\eta^{\prime\prime}\eta$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta a\in L$. Hence, as $L$ is reduced, $\alpha^{\prime\prime}\notin L$, so we can conclude again that $L$ is four-legged unless $\eta^{\prime\prime}$ is empty.
We conclude the claim: $L$ is four-legged unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$. ∎
From Section 6.1, we can conclude unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$, so we assume this in the rest of the proof of hardness for the overlapping case. We now know that $\alpha^{\prime}=\eta a\eta\in L$ and $\alpha=a\eta a\in L$.
Let us now re-use the fact that $\alpha$ was chosen to be a maximal-gap word of $L$. Now, as $\eta$ is non-empty, observe that $\alpha^{\prime}=\eta a\eta$ contains repeated letters across the two copies of $\eta$. These letters are spaced at least by the length of $\eta$ (and possibly more, if $\eta$ contains repeated letters – even though this cannot happen by the maximal-gap condition). In any case, $\alpha^{\prime}$ achieves a gap between two repeated letters which is at least that of $\alpha$: so $\alpha$ and $\alpha^{\prime}$ are tied for the first criterion in the definition of maximal-gap. Hence, by the second criterion, since $\alpha$ is maximal-gap, we know that $\alpha^{\prime}$ cannot be strictly longer than $\alpha$. So from $|\alpha^{\prime}|\leq|\alpha|$, as $|\alpha^{\prime}|=2|\eta|+1$ and $|\alpha|=|\eta|+2$, we conclude $|\eta|\leq 1$. As $\eta$ is nonempty, we have $|\eta|=1$. This letter may be different from $a$ or not. So we can conclude the overlapping case by showing two hardness results:
{claim}
Let $L^{\prime}$ be any reduced language containing $aba$ and $bab$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is hard.
Figure 10. Gadget used in the proof of Section 6.1
Proof.
We use the gadget whose completion is pictured on Figure 10 (left). It is obviously a pre-gadget, and one can check that the hypergraph of matches can be reduced to the odd chain of Figure 10 (right), so that it is indeed a gadget and we can conclude by Section 4.8. More precisely, observe that there are three kinds of directed paths in the gadget:
  • Paths of length 2 or less: all of these form a word which is a strict infix of $aba$ or $bab$, so they are not in the language by reducedness.
  • Paths of length 3: all of these form the word $aba$ or $bab$, so they are in the language.
  • Paths of length 4 or more: all of these have a path of length 3 as infix, and these always form words of the language by the previous point, so they are all subsumed by edge domination.
[…]
{claim}
Let $L$ be any reduced language containing the word $aaa$. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Figure 11. Gadget used in the proof of Section 6.1
Proof.
We use the gadget whose completion is pictured on Figure 11 (left) (which incidentally turns out to be identical to that of Figure 3(a) in Proposition 4.1). This is indeed a pre-gadget, and the hypergraph of matches can be condensed to Figure 11 (right), so it is indeed a gadget. Note that we use the fact that all directed paths in the gadget form words of the form $a^{*}$: by reducedness of the language, the matches precisely correspond to the words of length 3. ∎
Showing hardness in the non-overlapping case.
We conclude the proof of Section 6.1 by covering the non-overlapping case. Recall that, in this case, we know that the language $L$ contains a maximal-gap word $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, and recall that we already know that $\gamma_{1}$ and $\gamma_{2}$ are both non-empty.
We show:
{claim}
The language $L$ is four-legged unless $\gamma_{1}$ and $\gamma_{2}$ both have length 1.
Proof.
As $\gamma_{1}$ is nonempty, let us write $\gamma_{1}=\chi^{\prime}x$. If $\gamma_{1}$ has length 2 or more, then $\chi^{\prime}$ is nonempty, and we have $\alpha^{\prime}=\chi^{\prime}x(a\gamma_{2})\in L$ and $\alpha=(a\gamma_{2}\eta\chi^{\prime})xa\in L$. Let us take $x$ as body and consider the cross-product word $\alpha^{\prime\prime}=\chi^{\prime}xa=\gamma_{1}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$. So, as $L$ is reduced, we have $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged. Hence, it suffices to focus on the case where $\gamma_{1}=x$ has length 1.
Now, as $\gamma_{2}$ is nonempty, let us write $\gamma_{2}=y\chi^{\prime\prime}$. If $\gamma_{2}$ has length 2 or more, then $\chi^{\prime\prime}$ is nonempty, and we have $\alpha=ay(\chi^{\prime\prime}\eta\gamma_{1}a)\in L$ and $\alpha^{\prime}=(\gamma_{1}a)y\chi^{\prime\prime}\in L$. Let us take $y$ as body and consider the cross-product word $\alpha^{\prime\prime}=ay\chi^{\prime\prime}=a\gamma_{2}$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha\in L$, so as $L$ is reduced we know that $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged.
Hence, we conclude that $L$ is four-legged unless $\gamma_{1}$ and $\gamma_{2}$ both have length 1. ∎
Thanks to Section 6.1, we know that $\alpha=ax\eta ya\in L$ and $\alpha^{\prime}=yax\in L$, for some letters $x,y\in\Sigma$ and some word $\eta\in\Sigma^{*}$. So all that remains is to show:
{claim}
Let $L^{\prime}$ be a reduced language that contains words $ax\eta ya$ and $yax$ for some $a,x,y\in\Sigma$ and $\eta\in\Sigma^{*}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 12. Gadget used in the proof of Section 6.1
To show this claim, we make an auxiliary claim:
{claim}
Let $L^{\prime}$ be a reduced language that contains a word of the form $aab$ for $a,b\in\Sigma$ with $a\neq b$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Proof.
[…]
We can now conclude the proof of Section 6.1:
Figure 13. Gadget used in the proof of Section 6.1
Proof of Section 6.1.
We first rule out the case where $y=a$. Then $L^{\prime}$ contains $yax=aax$, and we conclude directly by Section 6.1 or Section 6.1 depending on whether $x=a$ or not.
We next rule out the case where $x=a$. Then, using Section 6.1, let us show hardness for $L^{\prime\mathrm{R}}$. This language contains $xay$, and as $x=a$ we can apply Section 6.1 or Section 6.1 again depending on whether $y=a$ or not, and conclude.
Hence, in the sequel, we assume that $x$ and $y$ are both distinct from $a$. (We may have $x=y$ or not.)
We then use the gadget whose completion is pictured in Figure 13 (left), with a condensed hypergraph of matches shown on figure Figure 13 (right). The same construction, and the same condensed hypergraph of matches, is used no matter whether $x$ and $y$ are the same letter or not.
To see why the condensed hypergraph of matches is as pictured on Figure 13, let us describe the directed paths in the completion of the gadget. The ones forming words of the form $ax\eta ya$, and $yax$, are guaranteed to be matches: the ones forming strict infixes of these are guaranteed not be matches because the language is reduced; and the ones containing one of these as an infix are guaranteed to be eliminated by node domination. In particular, paths that start within an $\eta$-edge, or end within an $\eta$-edge, or both, are not relevant. Indeed, observe that $\eta$-edges are always followed by (some prefix of) $yax\eta$. So a match starting within an $\eta$-edge (possibly using all facts of the edge) must include the subsequent $yax$ (because $\eta ya$ is a strict infix of $ax\eta ya$ and so is not in the language), and the match is then subsumed by the $yax$-match. Likewise, $\eta$-edges are always preceded by (some suffix of $\eta yax$), so matches ending within an $\eta$-edge (possibly using all facts of the edge) must include the preceding $yax$ (because $ax\eta$ is a strict infix of $ax\eta ya$) and then they are subsumed by the $yax$ match.
For this reason, to build the condensed hypergraph of matches, it suffices to consider paths that start and end at the edges labeled $a$ and $x$ and $y$. One can check that such paths either are infixes of $ax\eta ya$ or contain $yax$ as an infix, so their membership to $L$ is determined by the fact that $L$ contains $yax$ and $ax\eta ya$ and is reduced. Building the corresponding condensed hypergraph of matches and further applying condensation rules gives Figure 13 (right), which is an odd chain, so we conclude by Section 4.8. ∎
Applying Section 6.1 with $L^{\prime}\coloneq L$, we conclude that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard in the remaining cases of the non-overlapping case. This concludes the case distinction and concludes the proof of Section 6.1.
Connections to UCQs and self-joins.
Observe that finite languages $L$ correspond to the case of RPQs that are in fact unions of conjunctive queries (UCQs). The UCQs obtained from RPQs have a specific form: they are posed on an arity-two signature corresponding to the alphabet $\Sigma$, and their constituent conjunctive queries (CQs) are directed paths corresponding to words of the language $L$. In this sense, repeated letters in words correspond to self-joins in CQs. Thus, Section 6.1 implies that, for those UCQs that correspond to RPQs, the resilience problem is NP-hard unless every CQ is self-join-free. This contrasts with the setting of general UCQs, where resilience is sometimes tractable even in the presence of self-joins (DBLP:conf/pods/FreireGIM20). Specifically, in that setting, the following Boolean CQ on an arity-two signature is known to be in PTIME (DBLP:conf/pods/FreireGIM20) despite having two self-joins: $Q{\,:\!\!-\,}A(x),B(x),R(x,y),R(z,y),A(z),C(z)$.
Connections to other language classes.
We now explain how Section 6.1 relates to the languages discussed so far. We first perform a sanity check: such languages are never local.
{toappendix}
6.2. Proof of Section 6.2
{lemmarep}
Let $L$ be a finite language containing a word with a repeated letter. Then $L$ is not local.
{toappendix}
We point out that Section 6.2 does not in fact require the assumption that the language is reduced, and we do not use it in the proof:
Proof of Section 6.2.
As $L$ is finite and contains a word with a repeated letter, we can take a word with a repeated letter whose length is maximum among words of $L$ with a repeated letter. (Notice the similarity with the first criterion of the maximal-gap words of Definition 6.1.) Let $\alpha=\beta a\gamma a\delta\in L$ be such a word. Assume by contradiction that $L$ is local. Then, by Proposition 3.6, $L$ is letter-Cartesian. Applying the definition to $\alpha$ and $\alpha$, we must have that the following word is also in $L$:
$\beta a\gamma a\gamma a\delta\in L.$
This word is a word of $L$ which contains a repeated letter and is strictly longer than $\alpha$. This contradicts the maximality of $\alpha$, and concludes the proof. ∎
Note that the finiteness hypothesis is essential, e.g., $L=ax^{*}b$ contains words with repeated letters but it is local (see Example 3.2). Of course it is also essential that $L$ is reduced, e.g., $L=a|aa$ contains a word with a repeated letter but its reduction $\mathtt{reduce}(L)=a$ is local.
Further notice that there is no converse to Section 6.2: there are finite reduced non-local languages that feature no words with repeated letters and for which resilience is hard. This is for instance the case of the four-legged language $axb|cxd$ of Proposition 4.8. Thus, the languages covered by Section 6.1 and Section 5.2 are incomparable: they each generalize Proposition 4.1 and Proposition 4.8, respectively, without implying the other result. (However, the proof of Section 6.1 relies on Section 5.2.) These two results also do not cover all intractable cases, as we will now see.
7. Additional Complexity Results
We have shown the tractability of resilience for local languages (Theorem 3.3), and shown hardness for four-legged languages (Section 5.2) and for finite reduced languages featuring a word with a repeated letter (Section 6.1). Our results imply a dichotomy over languages with a neutral letter (Section 5.6), but not in general. In this section, we classify some languages not covered by these results, and highlight remaining open cases. We focus on languages that are finite and reduced.
We first illustrate two cases of finite languages that are tractable but not local. One are those we call bipartite chain languages, such as $ab|bc$. For them, we can compute resilience via a variant of the product construction of Theorem 3.3, simply by “reversing” some matches. Another is a class of languages (including for instance $abc|be$) which we show tractable via submodular minimization. These are the only tractable case where we do not know how to reduce to a MinCut problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable, and finally exhibit queries that remain unclassified.
Tractability of Bipartite Chain Languages
Let us define the chain languages:
Definition 7.1 (Chain languages).
A language $L$ is a chain language if:
  1. (1)
    No word in $L$ contains a repeated letter.
  2. (2)
    Each word of $L$ of length at least 2 may only share its first or last letter with other words in $L$: formally, for each word $\alpha=a\gamma b$ in $L$, no letter of $\gamma$ occurs in another word of $L$.
Note that chain languages are always finite, because they contain at most $|\Sigma|$ words of length $>2$: indeed, each word of length 3 or more contains intermediate letters which cannot occur in another word. We will focus on chain languages satisfying an extra condition, called being bipartite:
Definition 7.2 (Endpoint graph, BCL).
The endpoint graph of a language $L$ is the undirected graph $(\Sigma,E)$ having as vertices the letters of $\Sigma$, and having an edge $\{a,b\}$ for $a\neq b$ in $\Sigma$ iff $a$ and $b$ are the endpoints of some word of $L$ of length at least $2$ (i.e., $L$ contains a word of the form $a\alpha b$ or $b\alpha a$).
A chain language is a bipartite chain language (BCL) if its endpoint graph is bipartite.
Example 7.3 (Figure 14).
The languages $ab|bc$ (from Example 5.4) and $axyb|bztc|cd|dea$ are BCLs; the language $ab|bc|ca$ is a chain language but not a BCL. None of these languages are local.
Figure 14. Illustrations of the endpoint graphs for the three languages mentioned in Example 7.3.
Notice that the definition of BCLs is symmetric: any word $\alpha$ of $L$ can be replaced by the mirror word of $\alpha$ and the resulting language is still a BCL. Also notice that the class of BCLs is incomparable to local languages (even finite local languages): indeed, $ax^{*}b$ and $axb|axc$ are local languages but not BCLs, and Example 7.3 shows BCLs that are not local. Further note that BCLs containing no word of length less than 2 are always reduced. Further, we have the immediate analogue of Section 3.2: the reduced language of a BCL is always a BCL. More generally:
{toappendix}
7.1. Proof of Section 7.1
{lemmarep}
For any BCL $L$, any $L^{\prime}\subseteq L$ is also a BCL.
{toappendix}
We first show that being a chain language is closed under taking subsets:
Lemma 7.4.
For any chain language $L$, any $L^{\prime}\subseteq L$ is also a chain language.
Proof.
Let $L^{\prime}\subseteq L$. We check the conditions of chain languages:
  • $L^{\prime}$ cannot contain a word with a repeated letter
  • There cannot be a word $\alpha$ of $L^{\prime}$ sharing an intermediate letter with another word $\beta$ of $L^{\prime}$ without $\alpha$ and $\beta$ being counterexamples for $L$ being a chain language.
This establishes that $L^{\prime}$ is a chain language. ∎
We are now ready to prove Section 7.1:
Proof of Section 7.1.
Let $L$ be a BCL and take $L^{\prime}\subseteq L$. By Lemma 7.4, we know that $L^{\prime}$ is a chain language. Further, if the endpoint graph of $L$ is bipartite, then the same is true of $L^{\prime}$, because the endpoint graph of $L^{\prime}$ is a subgraph of that of $L$. Hence, $L^{\prime}$ is a BCL. ∎
We can show that resilience is tractable for all BCLs, even in combined complexity:
{toappendix}
7.2. Proof of Section 7.2: Bipartite Chain Languages are Tractable
{propositionrep}
[Bipartite chain languages are tractable] If $L$ is a BCL then $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. Moreover, the problem is also tractable in combined complexity: given an $\epsilon$NFA $A$ that recognizes a BCL language $L$, and a database $D$, we can compute $\mathtt{RES}_{\mathrm{bag}}(Q_{L},D)$ in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$.
{proofsketch}
We partition the letters of $\Sigma$ based on the bipartition of the endpoint graph. We then build a flow network like for Theorem 3.3, but we reverse some words depending on the sides of the partition in which their endpoints fall.
{toappendix}
To prove Section 7.2, we first show that an $\epsilon$NFAs recognizing a chain language can be efficiently converted to an explicit representation of the finite language that it recognizes:
Lemma 7.5.
Given an $\epsilon$NFA $A$ such that the language $\mathcal{L}(A)$ recognized by $A$ is a chain language, we can compute in time $O(|\Sigma|^{2}\times|A|)$ the finite language $L$ as an explicit list of words.
To prove this, we need an auxiliary claim that intuitively corresponds to the setting where we remove the first and last letters from a chain language:
{claim}
Let $A$ be an $\epsilon$NFA recognizing a finite language where no words contain a repeated letter and where no words share the same letter, i.e., for every word $\alpha\in\mathcal{L}(A)$ all letters of $\alpha$ are distinct, and for any two words $\alpha$ and $\beta$ with $\alpha\neq\beta$ the set of letters of $\alpha$ and of $\beta$ are distinct. Then $\mathcal{L}(A)$ is finite, and we can compute in $O(|A|)$ the finite language $\mathcal{L}(A)$ as an explicit list of words.
Proof.
The finiteness of the language is simply because it has cardinality at most $|\Sigma|+1$: except for the empty word, every word of the language eliminates a letter that cannot be used elsewhere. We assume that $A$ is trimmed (see Definition 3.5); this can be enforced in linear time by Section 3.1.2. We perform two graph traversals to compute in linear time the set $S_{l}$ of states having a path from some initial state via $\epsilon$-transitions, and the set $S_{r}$ of states having a path to some final state via $\epsilon$-transitions.
We first check if $S_{l}\cap S_{r}\neq\emptyset$: if so we add $\epsilon$ to the list of recognized words.
Now, we claim the following (*): for any state $s$ of $A$ which is not in $S_{r}$, there is at most one word $\alpha$ labeling the runs from the initial states of $A$ to $s$. Indeed, assume by way of contradiction that there are two such words $\alpha\neq\beta$. As $A$ is trimmed, let $\gamma$ be a word labeling a run from $s$ to a final state of $A$: as $s\notin S_{r}$, we know that $\gamma$ is non-empty. So $A$ accepts $\alpha\gamma$ and $\beta\gamma$, which are different words because $\alpha\neq\beta$, and which share letters (the ones in $\gamma$), contradicting the assumption on $\mathcal{L}(A)$.
Hence, we compute the list of the non-empty words accepted by $A$ by doing a graph exploration from $S_{l}$, maintaining for each state a word that corresponds to its witness word for the claim (*). The witness word of the states of $S_{l}$ is $\epsilon$. For any other state $s$ that is reached in the traversal, then either it is reached by an $\epsilon$-transition from a state $s^{\prime}$ and its witness word is the same as that of $s^{\prime}$, or it is reached by an $x$-transition for some letter $x$ from a state $s^{\prime}$ and its witness word is $\alpha x$ where $\alpha$ is the witness word of $s^{\prime}$. We store the witness words as pointers to a trie data structure:
  • The root of the trie corresponds to the word $\epsilon$ and the nodes of $S_{l}$ point to the root;
  • If we traverse a transition $s^{\prime}\xrightarrow{\epsilon}s$ then the pointer for $s$ is the same as the pointer for $s^{\prime}$;
[…]
When we reach a state of $S_{r}$ then we add a mark the corresponding node in the trie. Note that, as $A$ is trimmed, all leaves of the trie are marked.
At the end of this process, the trie has linear size because the graph traversal is in linear time. Further, les us show that the trie obeys a very strong structural restriction: the least common ancestor of any two marked nodes is the root (**). The claim (**) immediately implies that the trie has the following structure (***): a union of disjoint branches (i.e., the only node with more than one child is the root), with the marked nodes being precisely the set of leaves. Let us now show claim (**). Assume by way of contradiction that there are two marked nodes $n_{1}$ and $n_{2}$ whose lowest common ancestor $n$ is not the root. Then this witnesses that there is a non-empty word $\alpha$ (corresponding to the label of the path from the root to $n$) such that $\alpha\beta$ and $\alpha\gamma$ are accepted by the automaton, for two distinct $\beta$ and $\gamma$ corresponding to paths from $n$ to $n_{1}$ and $n_{2}$ respectively. However, $\alpha\beta$ and $\alpha\gamma$ are two distinct words which share the letters of $\alpha$, a contradiction again.
Thanks to the structure of the trie (***), we can finish by computing explicitly the labels of the paths from the root of the trie to the leaves. Claim (***) implies that this has size linear in the size of the trie, hence the overall process runs in $O(|A|)$. This concludes the proof. ∎
We can now prove our claim on $\epsilon$NFA recognizing chain languages (Lemma 7.5) from Section 7.2:
Proof of Lemma 7.5.
Let $A$ be the $\epsilon$NFA recognizing a language $L\coloneq\mathcal{L}(A)$ which is asserted to be a chain language. We assume that $A$ is trimmed (see Definition 3.5); this can be enforced in linear time by Section 3.1.2.
We first deal in a special way with words of length 00 and $1$ that are accepted by $A$. We first test in linear time as in the proof of Section 7.2 whether $A$ accepts $\epsilon$, and if so, add $\epsilon$ to the list of accepted words.
We then compute in linear time in $A$ the set $\Sigma_{1}$ of letters $a$ such that the single-letter word $a$ is accepted by $L$. We compute these by first computing in linear time the states $S_{l}$ accessible from initial states via $\epsilon$-transitions, computing in linear time the states $S_{r}$ accessible from final states by traversing $\epsilon$-transitions backwards, then going over all transitions with source in $S_{l}$ and target in $S_{r}$ and letting $\Sigma_{1}$ be the set of words which label such transitions. We add all the words of $\Sigma_{1}$ to the list of accepted words.
Hence, after this preprocessing in $O(|A|)$, we can focus on the words of length 2 or more that are accepted by $A$. We now consider every pair of letters $(a,b)\in\Sigma^{2}$ (possibly $a=b$), and compute explicitly the list of words accepted by $A$ that start with $a$ and end with $b$. Let $S_{l,a}$ be the set of states defined in the following way: we take all $a$-labeled transitions whose tail is in $S_{l}$, and let $S_{l,a}$ be the set of all states that are heads of such transitions. Further, let $S_{r,b}$ be the set of states defined in the following way: we take all $b$-labeled transitions whose head in is $S_{r}$, and let $S_{r,b}$ be the set of all states that are tails of such transitions. The sets $S_{l,a}$ and $S_{r,b}$ are computed in time $O(|A|)$. We will now compute explicitly the set of words $\alpha$ such that $a\alpha b\in L$. This clearly amounts to computing the words accepted by the $\epsilon$NFA $A_{a,b}$ obtained from $A$ by changing the initial states to be the states of $S_{l,a}$ and changing the final states to be the states of $S_{r,b}$. Hence, we can compute the list of words in time $O(|A|)$ by Section 7.2. The algorithm has an overall complexity of $O(|A|\times|\Sigma|^{2})$, which concludes the proof. ∎
We can now prove Section 7.2 from Lemma 7.5:
Proof of Section 7.2.
Thanks to Lemma 7.5, we compute in time $O(|A|\times|\Sigma|^{2})$ the language $L\coloneq\mathcal{L}(A)$ as an explicit list of words.
We first process $L$ in linear time to eliminate the words of length 00 or $1$. If $L$ contains $\epsilon$, then resilience is trivial; so in the sequel we assume that $L$ does not contain $\epsilon$. For every letter $a\in\Sigma$ such that $L$ contains the single-letter word $a$, then we know that in the input database $D$ we must remove every $a$-fact for the query to be false; so we can preprocess $D$ to remove all such facts and remove the corresponding one-letter words from $L$. Hence, at the end of this preprocessing, in time $O(|A|+|D|)$, we can assume that $L$ consists of words of length 2 or more.
We now point out that $L$ is reduced. This is by definition of a chain language and from the assumption that the words have length at least 2. Indeed, if $\alpha\in L$ is a strict infix of $\beta\in L$ then $\beta$ must have length at least 3, so there are letters in $\beta$ that are neither the first nor the last letter, by definition of $L$ being a chain language these letters cannot occur in a word of $L\setminus\{\beta\}$, so they do not occur in $\alpha$, so $\alpha$ being an infix of $\beta$ implies that $\alpha$ has length $1$, contradicting the assumption that the words of $L$ have length 2 or more.
We call an endpoint letter a letter of $\Sigma$ which occurs as the first or last letter of a word of $L$, i.e., has some incident edges in the endpoint graph $G$ of $L$. The set of such letters can be computed in linear time from the explicit list. We can also build explicitly the endpoint graph $G$ of $L$ from the explicit list, in linear time in the list. We know by assumption that $G$ is bipartite: we compute a witnessing bipartition in linear time by the obvious greedy algorithm. Let us assign each endpoint letter of $L$ to either the source or target partition. Now, the words of $L$ can be partitioned between the forward words, in which the first letter is assigned to the source partition and the last letter to the target partition; and the reversed words, where the opposite occurs.
Then we construct a flow network as follows:
  • We add two fresh vertices $t_{\text{source}}$ and $t_{\text{target}}$ which are the source and the target respectively of the network.
  • For every fact $a(v,v^{\prime})\in D$ we define two vertices $\mathrm{start}_{a,v,v^{\prime}}$ and $\mathrm{end}_{a,v,v^{\prime}}$.
  • For every letter $a$, there is an edge with capacity $\mathtt{mult}(a(v,v^{\prime}))$ from $\mathrm{start}_{a,v,v^{\prime}}$ to $\mathrm{end}_{a,v,v^{\prime}}$ for every $v,v^{\prime}$ such that we have $a(v,v^{\prime})\in D$.
  • For every forward word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity +$\infty$+ ∞ from $\mathrm{end}_{a,v,v^{\prime}}$ to $\mathrm{start}_{a,v^{\prime},v^{\prime\prime}}$
  • For every reversed word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity +$\infty$+ ∞ from $\mathrm{end}_{b,v^{\prime},v^{\prime\prime}}$ to $\mathrm{start}_{a,v,v^{\prime}}$.
  • There is an edge with capacity +$\infty$+ ∞ from $t_{\text{source}}$ to every vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is an endpoint assigned to the source partition, and from every vertex of the form $\mathrm{end}_{a,v,v^{\prime}}$ to $t_{\text{target}}$ where $a$ is an endpoint in the target partition.
The explicit list of words has size $O(|A|\times|\Sigma|^{2})$, and the construction runs in time $O(|D|^{2}\times|A|\times|\Sigma|^{2})$.
We can see that this construction contains a single edge for every fact, whose capacity is the multiplicity of said fact. Every witness of the instance corresponds to a path from source to target by construction. Indeed, for each forward word, we have a path from $t_{source}$ to the edge corresponding to its first letter, then following each letter in succession until we reach the last letter and then $t_{target}$. Further, for each backward word, we have a path from $t_{source}$ to the edge corresponding to its last letter, then following each letter in succession in reverse order, until we reach the first letter and $t_{target}$.
We must also show that there is no path from source to target in the network that does not correspond to a witness. Any path from $t_{source}$ to $t_{target}$ must start with a vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is assigned to the source partition (in particular it is not a letter that occurs in the middle of any word). We then move to a vertex $\mathrm{start}_{b,v^{\prime},v^{\prime\prime}}$ and there are two cases: either $a$ was the first letter of a forward word $\alpha$ and $ab$ occurs as an infix in that word, necessarily at the beginning because no middle letters are repeated, or $a$ was the last letter of a backward word $\alpha$ and $ba$ occurs in that word, again necessarily at the end because no middle letters are repeated. We describe the first case (corresponding to forward words), the second (corresponding to backward words) is symmetric. From the condition that intermediate letters of words do not occur in other words, whenever we reach a vertex, then the next vertex must correspond to the next letter of $\alpha$. Eventually we reach the last letter of $\alpha$ and the corresponding node has no outgoing edge in the flow network except the edge to $t_{\text{target}}$, so we terminate. Thus, all paths in this network correspond to witnesses.
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $Q_{L}$ for $D$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, as in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
As for chain languages $L$ that are not bipartite, we conjecture that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard. We leave this open, but show hardness for the specific non-bipartite chain language of Example 7.3:
{toappendix}
7.3. Proof of Section 7.3
{propositionrep}
The resilience problem for the language $ab|bc|ca$ is NP-hard.
Proof.
Figure 15. Gadget used in the proof of Section 7.3
We claim that the gadget shown in Figure 15 is valid for $ab|bc|ca$. Indeed, the gadget has the following properties: We can see that it meets all the criteria of being a pre-hardness gadget. We can also generate a condensed hypergraph of matches that forms an odd chain, and hence this gadget is a valid hardness gadget for $ab|bc|ca$. ∎
{toappendix}
7.4. Proof of Section 7.4: Tractability via Submodularity
Tractability due to submodularity.
We now present our second case of tractable languages. All PTIME cases discussed in this paper, as well as those discussed in previous work on resilience for conjunctive queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; makhija2024unified), are established by reducing to MinCut problems in a flow graph constructed in linear-time data complexity. But we now show that, in our setting, some languages are tractable for a different reason, by reducing to submodular function minimization. This implies a PTIME algorithm (MCCORMICK2005321), albeit with a worse polynomial degree than when reducing to MinCut.
{propositionrep}
[Tractability via submodularity] Let $\alpha=a_{1}\cdots a_{n}$ be a word where all letters are distinct, and let $a_{n+1}$ be a letter not occurring in $\alpha$. The problem $\mathtt{RES}_{\mathrm{bag}}(\alpha|a_{n-1}a_{n+1})$ is in PTIME.
Hence, the query $abc|be$, or the query $abcd|ce$ from Figure 1, enjoy tractable resilience.
{proofsketch}
We define a function on sets $Y$ of domain elements whose value is the resilience when we commit to removing precisely the $a_{n+1}$-facts whose tail is in $Y$. We show that this function is submodular, relying on a known result on the modularity of maximal flows (megiddo1974optimal, Lemma 3.2(ii)), which implies that we can find the set $Y$ minimizing the function and compute the resilience.
{toappendix}
We first recall the standard definition of submodular set functions:
Definition 7.6 (Submodular set function).
A submodular set function is a function of the form $f\colon 2^{V}\rightarrow\mathbb{R}$ that satisfies the following property: for every $A\subseteq S\subseteq V$ and $x\in S\setminus B$, we have that $f(A\cup\{x\})-f(A)\geq f(B\cup\{x\})-f(B)$.
Our result then hinges on a lemma about the resilience of singleton languages:
{lemmarep}
Let $\alpha=a_{1}\cdots a_{m}$ be a non-empty word without repeated letters, let $L=\{\alpha\}$ be the corresponding singleton language, and consider the query $L$. Let $D$ be a bag database (seen as a multiset). Then the function $f\colon 2^{\mathrm{Adom}(D)}\rightarrow\mathbb{N}$ which maps $Y\subseteq\mathrm{Adom}(D)$ to $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus\{a_{m}(u,v)\in D\mid u\in Y\})$ is a submodular set function.
In other words, if $\alpha$ has no repeated letters, then $\mathtt{RES}_{\mathrm{bag}}(\{\alpha\})$ is a submodular function of the set $Y$ of domain elements on which we remove all outgoing $a_{m}$-edges. Let us prove the lemma, using a known result from the study of multi-target flow problems (megiddo1974optimal, Lemma 3.2(ii)).
Proof of Definition 7.6.
To prove this, we formulate $\mathtt{RES}_{\mathrm{bag}}(L)$ as a single-source multi-target flow problem and use the existing submodularity result of (megiddo1974optimal).
Since $L$ is a singleton language without repeated letters, we know through the proof of Theorem 3.3 that it can be encoded as a MinCut problem on a product network construction. We can modify the product network construction to have multiple target nodes, with one target for each element of $\mathrm{Adom}(D)$. More precisely, let $q_{0},\ldots,q_{m}$ be the successive states of the obvious DFA for the language $\{\alpha\}$. The network as previously constructed consists of one target node $t_{\text{target}}$, and edges with capacity $\infty$ from nodes of the form $(v,q_{m})$ to $t_{\text{target}}$, then for each fact $f=a_{m}(u,v)$ there is an edge of weight $\mathtt{mult}(D,f)$ from $(u,q_{m-1})$ to $(v,q_{m})$. The nodes of the form $(v,q_{m})$ have no other incoming or outgoing edges, and the nodes of the form $(u,q_{m-1})$ have no other outgoing edges.
It is then clear that we can modify the product network to be a multi-target network and to have targets corresponding to each element of $\mathrm{Adom}(D)$: for each $u\in\mathrm{Adom}(D)$, letting $\kappa$ be the total number of facts of the form $a_{m}(u,v)$ in the (bag) database $D$, replace the outgoing edges from $(u,q_{m-1})$ with one outgoing edge with capacity $\kappa$ going to a fresh target node $t_{u}$. (In particular, when there are no facts of the form $a_{m}(u,v)$, then the node $(u,q_{m-1})$ had no outgoing edges and we add a useless outgoing edge of capacity zero to a fresh target node $t_{u}$. It is clear that this does not change the value of the flow.
The construction ensures that there is a bijection between the set of targets $t_{u}$ of the modified flow graph and $\mathrm{Adom}(D)$. It is also clear that, for any $Y\subseteq\mathrm{Adom}(D)$, the resilience $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus\{a_{m}(u,v)\in D\mid u\in Y\})$ is the value of the flow in the network where we remove the targets $\{t_{a}\mid a\in Y\}$. Now, for every subset $T$ of the targets $\{t_{a}\mid a\in\mathrm{Adom}(D)\}$, let $mc(T)$ be the cost of a minimum cut that separates the source from all targets of $T$. (In particular, the cost of a minimum cut of the original product network is $mc(C)$.) It is then known (megiddo1974optimal, Lemma 3.2(ii)), that for a flow graph with a single source and multiple targets, the function $mc(T)$ is a submodular set function. It is also known that if $f\colon 2^{V}\to\mathbb{N}$ is a submodular function then so is $g\colon 2^{V}\to\mathbb{N}$ defined by $g(A)\coloneq f(V\setminus A)$ (Clark2016, Lemma 1.5). Thus, the function that maps $Y\subseteq\mathrm{Adom}(D)$ to $mc(\mathrm{Adom}(D)\setminus Y)$ is also submodular. As we explained, $mc(\mathrm{Adom}(D)\setminus Y)$ is precisely the resilience of $L$ on $D$ with the facts of the form $\bigcup_{u\in Y}a_{m}(u,\_)$ removed. Thus, through the connection between the value of the flow network and the output to the resilience problem $\mathtt{RES}_{\mathrm{bag}}(L,D\setminus X)$, we can conclude that the desired function is submodular, as we claimed. ∎
We then use this lemma to show Section 7.4
Proof of Section 7.4.
Let $\alpha=a_{1}\cdots a_{n}$ and $\alpha^{\prime}=a_{n-1}a_{n+1}$ where all letters $a_{1},\ldots,a_{n+1}$ are distinct, and consider the finite reduced language $L=\alpha|\alpha^{\prime}$. The query $Q_{L}$ is equivalent to the following UCQ:
$\exists x_{1}\cdots x_{n+5}[\big{(}A_{1}(x_{1},x_{2})\wedge\cdots\wedge A_{n}(x_{n},x_{n+1})\big{)}\vee\big{(}A_{n-1}(x_{n+2},x_{n+3})\wedge A_{n+1}(x_{n+4},x_{n+5})\big{)}].$
For each value $v$ of the domain of the input database $D$, let $a_{i}(v,\_)$ be the multiset of edges of the form $\bullet\xrightarrow{a_{i}}v$ having $v$ as head, and similarly define $a_{i}(\_,v)$ to be the multiset of edges labeled $a_{i}$ with $v$ as source. We would like to express the function $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)$ in terms of the function $\mathtt{RES}_{\mathrm{bag}}(\alpha)$ and apply Definition 7.6.
To do this, we first observe that in any contingency set $S$, for any value $v$ in the domain, we must have either all edges of the form $a_{n-1}(\_,v)$ in $S$ or all edges of the form $a_{n+1}(v,\_)$ in $S$. (In particular, if one of these two sets is empty then we can vacuously take in $S$ the empty set of its edges.) Indeed, we can see that all edges of the form $a_{n-1}(\_,v)$ and $a_{n+1}(v,\_)$ create an $a_{n-1}a_{n+1}$-walk. Thus, if both sets are not fully included, then the facts of the nonempty set $a_{n-1}(\_,v)\setminus S$ together with those of the nonempty set $a_{n+1}(v,\_)\setminus S$ create an $a_{n-1}a_{n+1}$-walk in $D\setminus S$, contradicting the fact that $S$ is a contingency set. So indeed, for each value $v\in\mathrm{Adom}(D)$, any contingency set must contain all edges of at least one of the two sets $a_{n-1}(\_,v)$ or $a_{n+1}(v,\_)$.
Now we can express the resilience of the language $\alpha|a_{n-1}a_{n+1}$ in terms of the resilience of the language $\alpha$. For each value $v$ which is the middle vertex of an $a_{n-1}(\_,v)a_{n+1}(v,\_)$-walk, we can decide whether to include all edges from $a_{n-1}(\_,v)$ or all edges from $a_{n+1}(v,\_)$ in the contingency set and compute the resilience of the query $\alpha$ over the database with these edges removed. This ensures that there can be no matches of the query disjunct $a_{n-1}a_{n+1}$. Thus, the resilience of the query $\alpha|a_{n-1}a_{n+1}$ can be computed as:
(1) $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}\Big{\{}\sum_{v\in Z}|a_{n-1}(\_,v)|+\sum_{v\not\in Z}|a_{n+1}(v,\_)|+\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n-1}(\_,v))\Big{\}}$
where cardinalities denote multiset cardinalities. In other words, the resilience is computed by minimizing over subsets $Z$ of the domain elements: for $v\in Z$ we remove all incoming $a_{n-1}$-edges, for $v\notin Z$ we remove all outgoing $a_{n+1}$-edges, and then we find the resilience problem for the query $\alpha$ on the database where the $a_{n-1}$-edges leading into vertices of $Z$ have been removed.
Let $D_{n-1,Z}\coloneq D\setminus\bigcup_{v\in Z}a_{n-1}(\_,v))$, which occurs is the expression above, and let $D_{n,Z}\coloneq D\setminus\bigcup_{v\in Z}a_{n}(v,\_))$. We make the following claim, dubbed (*): $\mathtt{RES}_{\mathrm{bag}}(\alpha,D_{n-1,Z})=\mathtt{RES}_{\mathrm{bag}}(\alpha,D_{n,Z})$, i.e., the resilience of $D$ with all edges of form $a_{n-1}(\_,v)$ removed for all $v\in Z$ is the same as the resilience of $D$ with all edges of form $a_{n}(v,\_)$ removed for all $v\in Z$. Indeed, to show (*), first observe that minimum contingency sets for $\alpha$ on $D_{n-1,Z}$ never include edges of $a_{n}(v,\_)$ for $v\in Z$, because such edges never occur in $\alpha$-walks given that $v$ has no incoming $a_{n-1}$-edge; symmetrically, minimum contingency sets for $\alpha$ on $D_{n,Z}$ never include edges of $a_{n-1}(\_,v)$ for $v\in Z$, because such edges never occur in $\alpha$-walks given that $v$ has no outgoing $a_{n}$-edge. Now, to finish the proof of (*), letting $D^{\prime}\coloneq D\setminus\left(\bigcup_{v\in Z}a_{n-1}(\_,v)\cup\bigcup_{v\in Z}a_{n}(v,\_)\right)$, it suffices to show that any subset $S\subseteq D^{\prime}$ is a contigency set for $D_{n-1,Z}$ iff it is one for $D_{n,Z}$. Indeed, for one direction, taking such an $S$ such that $D_{n-1,Z}\setminus S$ does not satisfy $Q_{\alpha}$, it suffices to show that adding back the facts $\bigcup_{v\in Z}$ and removing those of $D\setminus\bigcup_{v\in Z}a_{n}(v,\_)$ cannot create a query match. But the added facts cannot be part of an $\alpha$-walk, because this walk would necessarily have $v\in Z$ as its last vertex, but there are no $a_{n}$-edges leaving out of $v$ in $D_{n,Z}$. The other direction is symmetrical. Hence, we have established (*).
From (*), we can thus restate Equation 1 as:
(2) $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}\Big{\{}\sum_{v\in Z}|a_{n-1}(\_,v)|+\sum_{v\not\in Z}|a_{n+1}(v,\_)|+\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n}(\_,v))\Big{\}}$
We next show that Equation 2 is a submodular minimization problem, and can be solved in PTIME. In other words, having fixed the input (multiset) database $D$, for $\mathtt{RES}_{\mathrm{bag}}(\alpha|\alpha^{\prime},D)=\min_{Z\subseteq\mathrm{Adom}(D)}f(D,Z)$ with $f$ defined as the expression inside the minimization in Equation 2 above, we would like to show that $f(D,Z)$ is submodular. The sum of submodular functions is a submodular function, so it suffices to show that each term of the sum is a submodular function in $Z$.
We can show that the function $f_{1}(Z)=\sum_{v\in Z}|a_{n-1}(\_,v)|$ is submodular because for each $Z\subseteq\mathrm{dom}(D)$ and $v\in\mathrm{dom}(D)\setminus Z$, we have $f_{1}(Z\cup\{v\})-f_{1}(Z)=|a_{n-1}(\_,v)|$. Further, $f_{2}(Z)=\sum_{v\not\in Z}|a_{n+1}(v,\_)|$ is submodular for the same reason: for each $Z\subseteq\mathrm{dom}(D)$ and $v\in\mathrm{dom}(D)\setminus Z$, we have $f_{2}(Z\cup\{v\})-f_{2}(Z)=-|a_{n+1}(v,\_)|$. Finally, since $\alpha$ is a non-empty word without repeated letters, it follows from Definition 7.6 that $f_{3}(Z)=\mathtt{RES}_{\mathrm{bag}}(\alpha,D\setminus\bigcup_{v\in Z}a_{n}(v,\_))$ is submodular over $\mathrm{Adom}(D)$, since $\alpha$ is a chain query.
In this way, we have shown that the resilience of the query $\alpha|\alpha^{\prime}$ can be phrased as a submodular minimization problem, and hence can be computed in PTIME. ∎
Interestingly, this proof technique does not apply even if we add a single letter to right of the first word - in particular, as we pointed out in the main text, we do not know if the resilience of the query $abce|be$ can be computed in PTIME. We also do not know what happens when the second word is longer, i.e., we do not know if the resilience of the query $abc|bef$ can be computed in PTIME. Note that some variants of the query are four-legged (e.g., $abc|ebf$, which up to renaming is $axb|cxd$, for which resilience is NP-hard by Proposition 4.8), and others are not four-legged but covered by Section 7.5.
Additional NP-hard cases and open problems.
Already in the case of finite reduced languages, there are gaps remaining between the tractable cases (local languages, and languages covered by Section 7.2 or Section 7.4) and the hard cases (languages containing a word with a repeated letter, and four-legged languages) that we have identified so far. We show below that some of these unclassified languages are NP-hard by building specific gadgets:
{toappendix}
7.5. Proof of Section 7.5
{propositionrep}
$\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$ and $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$ are NP-hard.
Proof.
Figure 16. Gadget used in the proof of Section 7.5 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$
Figure 17. Gadget used in the proof of Section 7.5 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$
We claim that the gadget shown in Figure 16 is valid for $abcd|be|ef$ and the gadget shown in Figure 17 is valid for $abcd|bef$. We can see that the gadgets meet all the criteria of a pre-hardness gadget, and show that there exists a condensed hypergraph of matches of the pre-gadgets for these queries that is an odd chain. These properties are checked programmatically, and we conclude that these gadgets are valid hardness gadgets for $abcd|be|ef$ and $abcd|bef$. ∎
We leave the question open for other languages, including $abcd|be$, $abc|bcd$, or $abc|bef$. Many infinite languages also remain unclassified, such as $ax^{*}b|xd$ for instance.
8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results show no difference between the two settings, unlike for CQs (makhija2024unified). We leave open the question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard. We also leave open the question of classifying the complexity of resilience in intermediate settings, e.g., bag semantics where only weights $1$, and $\infty$ are allowed; or the setting where some relations are indicated as exogenous in the signature, amounting to giving a weight of $\infty$ to all their facts.
Our work does not fully characterize the tractability boundary and leaves some cases open. At a high level, we still do not understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty. We also do not know if the submodularity approach is needed for queries like $abc|be$, or whether tractability can be obtained in easier ways. [^6] After we had submitted this paper for review, an argument was pointed out by Martín Muñoz (cstheory_martin) implying that Section 7.4 can be shown using a flow encoding instead of submodular function minimization. Further, his argument seems to imply that the tractability of resilience extends to more queries, such as $abcd|be$. We leave the exploration of these consequences to a future version of the paper.
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity; 10.1145/1868237.1868241), or multicuts (garg1994multiway) and skew multicuts (kratsch2015fixed). Note that this non-Boolean setting corresponds to the problem of deletion propagation with source side effects (buneman2002propagation), which was studied before resilience: this problem asks about finding a minimum contingency set to remove a specific query answer. The case of removing multiple tuples has also been investigated (kimelfeld2013multituple). We do not know whether a dichotomy in one of these settings implies a dichotomy for our problem, or vice-versa. We also do not understand how our results relate to VCSPs (bodirsky2024complexity), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $1,\infty$).
Last, resilience can be studied for recursive queries beyond RPQs, in particular two-way RPQs or Datalog, though this would require new techniques (unlike RPQs, these queries are not “directional”).
9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages, and to the reviewers for their valuable feedback. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.