### v3

# Resilience for Regular Path Queries: Towards a Complexity Classification

Abstract.
The resilience problem for a query and an input set or bag database is to compute the minimum number of facts to remove from the database to make the query false. In this paper, we study how to compute the resilience of Regular Path Queries (RPQs) over graph databases. Our goal is to characterize the regular languages $L$ for which it is tractable to compute the resilience of the existentially-quantified RPQ built from $L$.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter. We last show tractability for some classes of non-local languages, namely the so-called bipartite chain languages and one-dangling languages, and highlight some remaining obstacles towards a full dichotomy.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXX
1. Introduction
[…]
Figure 1. Overview of our results on the complexity of resilience (all results apply both to bag and set semantics), with some examples of languages. All languages are infix-free. For legibility, we omit Proposition 5.7.
This paper studies the resilience problem which asks, for a Boolean query $Q$ and a given database $D$ (in set or bag semantics) what is a minimum-cardinality set of tuples to remove from $D$ so that $Q$ is no longer satisfied. (In particular: if $Q$ does not hold on $D$ then resilience is zero; if $Q$ holds on $D$ and all subinstances of $D$ then resilience is infinite.) The notion of resilience is motivated by the need to identify how “robust” a query is, especially when the facts of the input database may be incorrect. Specifically, resilience quantifies how many tuples of $D$ must be removed before $Q$ no longer holds, so that higher resilience values mean that $Q$ is less sensitive to tuples of $D$ being removed.
The resilience problem has been introduced for Boolean conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15, ; DBLP:conf/pods/FreireGIM20, ; bodirsky2024complexity, ; makhija2024unified, ). The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified, )). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
This leads to a natural question: Can one give a characterization of the queries for which resilience can be tractably computed? Further, can one show a dichotomy result to establish that resilience is always either in PTIME or NP-complete, depending on the query? This question has remained elusive, already in the case of conjunctive queries (CQs).
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, the only other work which studies this problem is the recent paper of Bodirsky et al. (bodirsky2024complexity, ), in which the authors use CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for Boolean RPQs (and also for the broader class of so-called two-way RPQs aka 2RPQs). However, their lower bounds do not apply to databases in set semantics. In addition, their dichotomy criterion, while decidable, does not give an intuitive language-theoretic characterization of the tractable cases. We discuss this in more detail at the end of Section 2.
In contrast with CQs, RPQs are in certain ways harder to analyze, and in other ways easier. On the one hand, RPQs feature the union operator and the recursion operator; and they inherently correspond to queries with self-joins, because the same relation symbols (corresponding to alphabet letters) may occur multiple times in query matches. On the other hand, RPQs do not allow all features of CQs: they ask for a simple pattern shape (namely, a walk, which importantly is always directed), and they are posed over graph databases (in which all facts have arity two).
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem, allowing multiple sources and sinks without loss of generality, is easily seen to be equivalent to the resilience problem in bag semantics for the RPQ $ax^{*}b$ on the alphabet $\{a,b,x\}$, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20, ), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
Contributions and outline.
After giving preliminary definitions and stating the problem in Section 2, we start in Section 3 by giving our main tractability result (Theorem 3.13): resilience is tractable for RPQs defined from so-called local languages (mpri, ; sakarovitch2009elements, ). Those languages are intuitively those that are defined by specifying the possible initial labels of a path, the possible final labels of a path, and the consecutive pairs of labels that are allowed. We show that, for RPQs defined with a local language $L$, we can compute resilience by reducing to the MinCut problem via a natural product construction. The construction applies to queries both in set and bag semantics, and it is also tractable in combined complexity, running in time $\tilde{O}(|A|\times|\Sigma|\times|D|)$ where $A$ is an automaton representing $L$, where $\Sigma$ is the signature, and where $D$ is the input database.
We then turn to intractability results in Sections 4, 5 and 6. In all these results, we assume that we focus on the infix-free sublanguage where redundant words have been eliminated, e.g., resilience for $abbc|bb$ and $bb$ is the same because the corresponding (existentially-quantified) Boolean RPQs are equivalent.
We define in Section 4 the notion of a hardness gadget for a language $L$, namely, a specific database to be used as a building block in a reduction. These gadgets are the analogue, in our setting, of the Independent Join Paths (IJPs) used in earlier work (makhija2024unified, ) to show hardness for CQs. When an RPQ $Q$ has a hardness gadget $\Gamma$, we can reduce the minimum vertex cover problem on an undirected graph $G$ to a resilience computation (in set semantics) for $Q$ on a database built using $\Gamma$, by replacing the edges of $G$ by a copy of $\Gamma$. We illustrate in Section 4 how our gadgets can be used to show hardness for two specific RPQs: $axb|cxd$ (Proposition 4.13), and $aa$ (Proposition 4.1).
We then generalize these two hardness results in turn to apply to a broader class of languages. First, in Section 5, we introduce the notion of a four-legged language, as a language featuring a specific kind of counterexample to being local. More precisely, the local languages of Section 3 have the following known equivalent characterization (sakarovitch2009elements, ): for any letter $x$ and words $\alpha,\beta,\gamma,\delta$, if a language $L$ contains the words $\alpha x\beta$ and $\gamma x\delta$ but does not contain the “cross-product” word $\alpha x\delta$, then $L$ is not local. A four-legged language is a non-local language that features a counterexample of this form in which all the words $\alpha,\beta,\gamma,\delta$ are required to be non-empty. This covers, for instance, $axb|cxd$, but not $ab|bc$. We show (Theorem 5.3) that resilience is intractable for any four-legged language. This result implies hardness for many languages: it covers all non-star-free languages (Lemma 5.6), a well-known class of infinite regular languages; and, along with the tractability of local languages, it implies a dichotomy for languages having a neutral letter (Proposition 5.7).
Second, in Section 6, we show that resilience is hard for any finite infix-free language containing a word with a repeated letter (Theorem 6.1). This generalizes the case of $aa$, and shows that self-joins always imply hardness in the following sense: finite RPQs (those equivalent to UCQs), once made infix-free, have intractable resilience as soon as one of their constituent CQs features a self-join. This contrasts with the setting of CQs, where characterizing the tractable queries with self-joins is still open.
Note that our two hardness results (Theorem 5.3 and Theorem 6.1) are incomparable: languages with a repeated letter are non-local, but they are not all four-legged (e.g., $aa$); conversely, some four-legged languages have no words with a repeated letter (e.g., $axb|cxd$). Assuming finiteness in Theorem 6.1 is also crucial, as $ax^{*}b$ features words with repeated letters and still enjoys tractable resilience. The proof of Theorem 6.1 is our main technical achievement: it relies on Theorem 5.3 and uses different gadgets depending on which additional words the language contains.
Section 7 then explores how our results do not classify all RPQs. We mostly focus on finite RPQs: the remaining open cases are then the RPQs which have no words with repeated letters, are not local, but are not four-legged. We introduce the class of chain languages (e.g., $axb|byc$) to cover some of these RPQs, and show tractability (and reducibility to MinCut) assuming a bipartiteness condition. We then introduce the class of one-dangling languages (e.g., $abc|be$), for which we show tractability (via a more elaborate reducibility to MinCut). We finish the section by showing that some other unclassified RPQs are intractable. We conclude the paper in Section 8.
Conference version. This paper is an extended version of an earlier conference paper (amarilli2025resilience, ). Compared to the conference version, all our tractability results in the current paper use a unified reduction to flow. Some of the new proofs are considerably simpler, and we currently do not have any example of a query with tractable resilience for which tractability does not follow by reducing to an appropriate network flow problem. Our current results support our hypothesis that flow is all one needs for tractable cases, even those that are currently open. In more detail, we have improved Proposition 7.9 (compared to (amarilli2025resilience, , Prp. 7.7)) by showing a tractability result that covers a larger family of languages (dubbed one-dangling languages) and uses a reduction to a flow problem. This yields tractability in combined complexity and gives a better exponent than an earlier proof via submodular optimization from (amarilli2025resilience, ). An overview of our results in the present paper is in Figure 1. Compared with (amarilli2025resilience, , Figure 1), the languages $abcd|be$ and $ax^{*}b|xd$ have now been classified as tractable thanks to the new Proposition 7.9, and new unclassified languages $ab^{*}c|ba$ and $ab^{*}d|ac^{*}d|bc$ were added. The present paper also includes the complete proofs for all results, which had been omitted from the conference article.
2. Preliminaries and Problem Statement
Formal languages.
Let $\Sigma$ be an alphabet, consisting of letters. A word over $\Sigma$ is a finite sequence of elements of $\Sigma$. We write $\epsilon$ for the empty word. We write $\Sigma^{*}$ for the set of words over $\Sigma$, and write $\Sigma^{+}\coloneq\Sigma^{*}\setminus\{\epsilon\}$. We use lowercase Latin letters such as $a,b,x,y$ to denote letters in $\Sigma$, and lowercase Greek letters like $\alpha,\beta,\gamma,\tau$ to denote words in $\Sigma^{*}$. The concatenation of two words $\alpha,\beta$ is written $\alpha\beta$. We write $|\alpha|$ for the length of a word. A word $\alpha$ is an infix of another word $\beta$ if $\beta$ is of the form $\delta\alpha\gamma$, and it is a strict infix if moreover $\delta\gamma\neq\epsilon$. If $\beta=\alpha\gamma$ then $\alpha$ is a prefix of $\beta$, and a strict prefix if $\gamma\neq\epsilon$. Likewise if $\beta=\delta\alpha$ then $\alpha$ is a suffix of $\beta$, and a strict suffix if $\delta\neq\epsilon$.
[…]
We also use regular expressions to denote regular languages: we omit their formal definition, see (sakarovitch2009elements, ). We write $|$ for the union operator and * for the Kleene star operator.
Databases, queries, resilience.
A graph database on the alphabet $\Sigma$, or simply database, is a set of $\Sigma$-labeled edges of the form $D\subseteq V\times\Sigma\times V$, where $V$ is the set of nodes of $D$: we also call $V$ the domain and write it as $\mathrm{Adom}(D)$. Notice that a graph database can equivalently be seen as a relational database on an arity-two signature that directly corresponds to the alphabet $\Sigma$. We also call facts the edges $(v,a,v^{\prime})$ of $D$, and write them as $v\xrightarrow{a}v^{\prime}$. We call $v$ the source (or tail) of $v\xrightarrow{a}v^{\prime}$ and $v^{\prime}$ its target (or head). A Boolean query is a function $Q$ that takes as input a database $D$ and outputs $Q(D)\in\{0,1\}$. We write $D\models Q$, and say that $Q$ is satisfied by $D$, when $Q(D)=1$, and write $D\not\models Q$ otherwise. In this paper we study the resilience of queries:
Definition 2.1 (Resilience).
Let $D$ be a database and $Q$ a Boolean query. The resilience of $Q$ on $D$, denoted $\mathtt{RES}_{\mathrm{set}}(Q,D)$, is defined [^1] By convention, if $Q$ is satisfied by all subsets of $D$, then the minimum is defined to be $\infty$. as $\mathtt{RES}_{\mathrm{set}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}|D^{\prime}|$.
In other words, the resilience is the minimum number of facts to remove from $D$ so that $Q$ is not satisfied. A subset of facts $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$ is called a cut (or a contingency set) of $D$ for $Q$. Notice that if $D\not\models Q$, then $\emptyset$ is a cut and we have [^2] We deviate from the original definition in (DBLP:journals/pvldb/FreireGIM15, ) which left resilience undefined for the case of $D\not\models Q$. $\mathtt{RES}_{\mathrm{set}}(Q,D)=0$.
Regular Path Queries.
From a language $L$ over $\Sigma$ we now define a query over graph databases that tests the existence of walks labeled by words of $L$. Given a database $D$, a walk in $D$ is a (possibly empty) sequence of consecutive edges $v_{1}\xrightarrow{a_{1}}v_{2}$, $v_{2}\xrightarrow{a_{2}}v_{3}$, $\ldots$, $v_{k}\xrightarrow{a_{k}}v_{k+1}$ in $D$. Note that the $a_{i}$ and $v_{i}$ are not necessarily pairwise distinct. We say that the walk is labeled by the word $\alpha=a_{1}\cdots a_{k}$, and call it an $\alpha$-walk. We also say that the walk contains a word $\beta\in\Sigma^{*}$ if $\beta$ is an infix of $\alpha$. An $L$-walk in $D$ is then an $\alpha$-walk for some $\alpha\in L$. We define the Boolean query $Q_{L}$ as: $Q_{L}(D)=1$ iff there is some $L$-walk in $D$. In particular, if $\epsilon\in L$ then $Q_{L}(D)=1$ on every database $D$ (including $D=\emptyset$). A regular path query (RPQ) is then a Boolean query of the form $Q_{L}$ for $L$ a regular language. Note that, in all of the present paper, we evaluate RPQs with walk semantics (i.e., arbitrary paths with repetitions), as opposed to other works such as (martens2019dichotomies, ) which evaluate them with simple path or trail semantics: we leave open the question of studying resilience for RPQs in simple path or trail semantics.
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the infix-free language of $L$ is $\mathtt{IF}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{IF}(L)}$ are the exact same queries. The infix-free language of $L$ can also be called the factor-free sublanguage of $L$ in (barcelo2019boundedness, ), and is called the reduced sublanguage in our earlier conference paper (amarilli2025resilience, ). It is easy to show that if $L$ is regular then so is $\mathtt{IF}(L)$. We say that $L$ is infix-free if $L=\mathtt{IF}(L)$; this is also called being an infix code (ito1991outfix, ).
Bag semantics.
We also consider the resilience problems on bag graph databases, as opposed to set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. In the input, the multiplicities are expressed as integers written in binary. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set of $D$ for $Q$ is then again a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
Networks and cuts.
Our resilience algorithms work by reducing to the problem of finding a minimum cut of a flow network (Cormen:2009dz, ). Recall that a flow network consists of a directed graph $\mathcal{N}=(V,t_{\text{source}},t_{\text{target}},E,c)$ with $V$ being the vertices, $t_{\text{source}},t_{\text{target}}\in V$ being special vertices called the source and the target respectively, $E\subseteq V\times V$, and $c:E\to\mathbb{R}_{+}\cup\{+\infty\}$ giving the capacity of each edge. For $E^{\prime}\subseteq E$, we define the cost of $E^{\prime}$ as $c(E^{\prime})\coloneq\sum_{e\in E^{\prime}}c(e)$, and denote $\mathcal{N}\setminus E^{\prime}$ as the network in which all edges from $E^{\prime}$ are removed. A cut is a set of edges $E^{\prime}\subseteq E$ such that there is no directed path from $t_{\text{source}}$ to $t_{\text{target}}$ in $\mathcal{N}\setminus E^{\prime}$. The min-cut problem, written MinCut, takes as input a flow network and returns the minimum cost of a cut. From the max-flow min-cut theorem (ford1956maximal, ) and Menger’s theorem (menger:1927, ), we know that the min-cut problem can be solved in PTIME. More precisely, practical algorithms can run in $O(|V|^{2}\sqrt{|E|})$ (ahuja1997computational, ), and recent theoretical algorithms run in $\tilde{O}(|\mathcal{N}|)$ (henzinger2024deterministic, ), where $|\mathcal{N}|\coloneq|V|+|E|$ and where $\tilde{O}$ hides polylogarithmic factors.
Problem statement.
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-complete (makhija2024unified, , Table 1) (even when the multiplicities of the input bag database are written in unary). We also note that the definition given here corresponds to a decision problem of determining whether the resilience is no greater than the value $k$. However, thanks to self-reducibility, the function problem of computing a witness contingency set reduces in polynomial time to this decision problem: see (bodirsky2025complexity, , Remark 8.4) for a self-contained argument.
Our definitions correspond to the data complexity perspective, where we have one computational problem per language $L$ and where the complexity is measured as a function of $|D|$ only. We mostly focus on data complexity, but most of our upper bounds also apply in combined complexity, i.e., when both the database $D$ and an automaton $A$ representing $L$ are given as input.
It is easy to see that, for any regular language $L$, the problem $\mathtt{RES}_{\mathrm{bag}}(L)$, hence $\mathtt{RES}_{\mathrm{set}}(L)$, is in NP. Indeed, given a bag database $D$, it suffices to guess a subdatabase $D^{\prime}\subseteq D$, and verify in PTIME that $\sum_{f\in D^{\prime}}\mathtt{mult}(f)$ is small enough and that $D^{\prime}$ is a contingency set, i.e., $D\setminus D^{\prime}\not\models Q$. The latter point, i.e., the nonexistence of an $L$-walk in $D\setminus D^{\prime}$, can easily be checked in PTIME by a product construction followed by a reachability test (see, e.g., (DBLP:journals/siamcomp/MendelzonW95, , Lemma 3.1)).
However, the NP upper bound is not always tight: there are languages $L$ (e.g., $L=a|b$) such that $\mathtt{RES}_{\mathrm{bag}}(L)$ is in PTIME. The goal of our paper is to characterize the languages $L$ for which $\mathtt{RES}_{\mathrm{set}}(L)$ and $\mathtt{RES}_{\mathrm{bag}}(L)$ are in PTIME, and give efficient algorithms for them.
Connection to VCSPs.
The recent work of Bodirsky et al. (bodirsky2024complexity, ) uses earlier results on Valued Constraint Satisfaction Problems (VCSP) to obtain the following dichotomy on $\mathtt{RES}_{\mathrm{bag}}$:
Theorem 2.2 ((bodirsky2024complexity, , Remark 8.12)).
Let $L$ be a regular language. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ is either in PTIME or it is NP-complete.
We point out that this result can be extended to so-called two-way RPQs, and to the case where some relations are declared exogenous (i.e., they cannot be part of a contingency set). Hence, as discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages. However, it also has four important shortcomings: 1  The VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity, ), using (kolmogorov2019testing, ), can be computationally checked but takes time exponential in a representation of the so-called dual of $L$. 2  In tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that the first two difficulties also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds, for instance in (hirai2021minimum, ). 3  The VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and the database are given as input. 4  The VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$. Our results improve on some of these shortcomings. Specifically, all our intractability results already apply to the $\mathtt{RES}_{\mathrm{set}}$ problem, they apply to combined complexity (when the query is given as an automaton along with the database), and they feature a polynomial complexity with a low degree (and in fact up to polylogarithmic factors the data complexity is always linear except for one result, Proposition 7.6, where it is quadratic). However, unlike (bodirsky2024complexity, ), our results do not establish a full dichotomy.
3. Tractability for Local Languages
We now start the presentation of our results on the resilience problem by giving our main tractability result in this section. The result applies to local languages [^3] We note that local languages were also recently observed to correspond to a tractable class of RPQs for which probabilistic query evaluation is equivalent to source-to-target reliability, see (amarilli2025approximating, ). , a well-known subclass of the regular languages (mpri, ; sakarovitch2009elements, ), whose definition we recall below. We show that the resilience problem is tractable for local languages, also in combined complexity, using a specific notion of automata and a product construction to reduce resilience to MinCut. We stress that local languages do not cover all cases of regular languages with tractable resilience: we will show other tractable cases in Section 7. However, in Section 5 we will give a partial converse: resilience is hard for the so-called four-legged languages, which have a specific form of counterexample to being local.
The remainder of this section is structured as follows: in Section 3.1 we define what are local languages and provide an equivalent characterization that we will use later. We then show in Section 3.2 that resilience is PTIME for local languages.
3.1. Local Languages
We define here local languages and show some useful properties. Intuitively, a language $L$ is local if membership to $L$ is described by which letter pairs can occur consecutively and which letters can start or can end words. To define them formally, we use the notion of so-called local DFAs (mpri, , Chapter III, Section 5.1):
Definition 3.1 (Local DFA).
A DFA is local if, for every letter $a\in\Sigma$, all $a$-transitions in the automaton have the same target state. A language $L$ is local if there exists a local DFA $A$ that recognizes $L$.
Example 3.2.
The languages $ax^{*}b$ and $ab|ad|cd$ are local, with local DFAs in Figure 2(a) and Figure 2(b).
$s_{1}$$s_{2}$$s_{3}$aaxxbb
(a) Local DFA $A_{1}$ for $ax^{*}b$
$s_{1}$$s_{2}$$s_{3}$$s_{4}$$s_{5}$aabbddccdd
(b) Local DFA $A_{2}$ for $ab|ad|cd$
[…]
Letter-Cartesian languages.
We give a known equivalent language-theoretic characterization of local languages, that we call being letter-Cartesian and that will be useful later:
Definition 3.3 (Letter-Cartesian).
A language $L$ is letter-Cartesian if the following implication holds: for every letter $x\in\Sigma$ and words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$, then $\alpha x\delta\in L$.
Note that, by exchanging the roles of $\alpha,\beta$ and $\gamma,\delta$, if $\alpha x\beta\in L$ and $\gamma x\delta\in L$ then also $\gamma x\beta\in L$. Hence the name letter-Cartesian: for each letter $x\in\Sigma$, the words of $L$ containing an $x$ can be expressed as a Cartesian product between what can appear left of an $x$ and what can appear right of an $x$. Formally, we have $L=\bigcup_{x\in\Sigma}L_{x}xR_{x}$ for some $L_{x},R_{x}\subseteq\Sigma^{*}$.
Example 3.4.
The local languages of Example 3.2 are letter-Cartesian. Take now $L=aa$. With $x\coloneq a$, $\alpha\coloneq a$, $\beta\coloneq\epsilon$, $\gamma\coloneq\epsilon$, $\delta\coloneq a$, we see that $L$ is not letter-Cartesian (as $aaa\notin L$), and hence, as we will see, it is not local.
It is known that letter-Cartesian languages are precisely the local languages:
Proposition 3.5.
A language is local if, and only if, it is letter-Cartesian.
This result has been claimed in (sakarovitch2009elements, , Chp. II, Ex. 1.8, (d)) without proof (as an exercise). For completeness, and to provide more intuition on local languages to the reader, we provide a proof here. One direction is easy to show: local languages are always letter-Cartesian.
Claim 3.6.
If $L$ is local then it is letter-Cartesian.
Proof.
Let $A$ be a local DFA recognizing $L$. Let $x\in\Sigma$, and let $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$, and let us show that $\alpha x\delta$ is in $L$ as well. Because $L$ is local, the state $q$ that is reached in $A$ after reading $\alpha x$ in the accepting run of $\alpha x\beta$ is the same state that is reached after reading $\gamma x$ in the accepting run of $\gamma x\delta$. We can thus compose a run from the initial state to $q$ reading $\alpha x$, with the run from $q$ to some final state reading $\delta$, to form an accepting run for $\alpha x\delta$. This concludes the proof. ∎
We now prove the converse, thus establishing Proposition 3.5:
Claim 3.7.
If $L$ is letter-Cartesian then it is local.
To establish this claim, we use a construction from (berstel1996local, , Proposition 2.1). We define the local overapproximation of a language $L$ as the local DFA constructed as follows:
Definition 3.8.
Let $L$ be a language. We start by defining the following:
  • The set $\Sigma_{\mathrm{start}}$ of letters that can start words of $L$, i.e., the letters $a\in\Sigma$ such that there exists a word $a\alpha\in L$ for some $\alpha\in\Sigma^{*}$.
  • The set $\Sigma_{\mathrm{end}}$ of letters that can end words of $L$, i.e., letters $a\in\Sigma$ such that $\alpha a\in L$ for some $\alpha\in\Sigma^{*}$.
  • The set $\Pi\subseteq\Sigma^{2}$ of pairs of letters that can occur consecutively in a word of $L$, formally, the pairs $(a,b)$ such that there exist $\alpha,\beta\in\Sigma^{*}$ such that $\alpha ab\beta\in L$.
We now define the DFA $A$ in the following way. We have an initial state $q_{0}$, which is final iff $\epsilon\in L$. For every $a\in\Sigma$, we have a state $q_{a}$, which is final iff $a\in\Sigma_{\mathrm{end}}$. Then, for every $a\in\Sigma_{\mathrm{start}}$ we add an $a$-transition from $q_{0}$ to $q_{a}$, and for every $(a,b)\in\Pi$ we add a $b$-transition from $q_{a}$ to $q_{b}$.
The DFA $A$ built from $L$ is local by definition, and the following is immediate:
Claim 3.9.
We have $\mathcal{L}(A)\supseteq L$.
We now claim that, for letter-Cartesian languages, the local overapproximation is in fact equal to the original language, namely:
Claim 3.10.
If $L$ is letter-Cartesian, then $\mathcal{L}(A)=L$.
This immediately implies 3.7: if $L$ is letter-Cartesian, then it is local, because $L$ is equal to $\mathcal{L}(A)$ for $A$ the local overapproximation, and $A$ is a local DFA. So all that remains is to show 3.10.
Proof of 3.10.
We must show that if $L$ is letter-Cartesian then $\mathcal{L}(A)=L$. By 3.9, it suffices to show that $\mathcal{L}(A)\subseteq L$.
Take $\alpha$ in $\mathcal{L}(A)$, and let us show that it is in $L$. If $\alpha$ is the empty word then it is clear by construction that $\alpha$ is also in $L$. So assume $\alpha=a_{1}\cdots a_{m}$ with $m\geq 1$. By construction of $A$, the fact that $\alpha$ is accepted witnesses that:
  • $a_{1}\in\Sigma_{\mathrm{start}}$, hence there exists $\alpha^{\prime\prime}\in\Sigma^{*}$ such that $a_{1}\alpha^{\prime\prime}\in L$;
  • $a_{m}\in\Sigma_{\mathrm{end}}$ hence there exists $\alpha^{\prime\prime\prime}\in\Sigma^{*}$ such that $\alpha^{\prime\prime\prime}a_{m}\in L$;
  • For each $1\leq i<m$ we have $(a_{i},a_{i+1})\in\Pi$, hence there exist $\alpha_{i}^{\prime},\alpha_{i}^{\prime\prime}\in\Sigma^{*}$ such that $\alpha_{i}^{\prime}a_{i}a_{i+1}\alpha_{i}^{\prime\prime}$ is in $L$.
We now use the fact that $L$ is letter-Cartesian to prove the following claim (*): $\alpha_{1}^{\prime}\alpha\alpha^{\prime\prime}_{m-1}\in L$.
To do this, we show by finite induction that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ for each $1\leq i<m$. The base case of $\alpha^{\prime}_{1}a_{1}a_{2}\alpha^{\prime\prime}_{1}\in L$ is by the definition of $\alpha_{1}^{\prime}$ and $\alpha_{1}^{\prime\prime}$ above. The induction case for 1<i<m11<i<m-1 is as follows: from $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}\alpha^{\prime\prime}_{i}\in L$ (by induction hypothesis) and $\alpha_{i+1}^{\prime}a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ (by definition of $\alpha_{i+1}^{\prime},\alpha_{i+1}^{\prime\prime}$ above), we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{i+1}a_{i+2}\alpha^{\prime\prime}_{i+1}\in L$ because $L$ is letter-Cartesian on $a_{i+1}$, which is what we wanted to show to establish the induction case. Hence, the inductive claim holds and for i=m1i=m-1 we deduce that $\alpha_{1}^{\prime}a_{1}\cdots a_{m}\alpha^{\prime\prime}_{m-1}\in L$, which is claim (*). Now, using that $L$ is letter-Cartesian on $a_{1}$, from (*) and the word $a_{1}\alpha^{\prime\prime}\in L$ identified earlier, we deduce that $\alpha\alpha^{\prime\prime}_{m-1}$ is in $L$. Symmetrically, from this and the word $\alpha^{\prime\prime\prime}a_{m}\in L$ identified earlier, using the fact that $L$ is letter-Cartesian on $a_{m}$ we deduce that $\alpha\in L$, which concludes. ∎
We have concluded the proof of Proposition 3.5.
Testing if a language is local.
We can use the local overapproximation (Definition 3.8) to show that it is PTIME to determine whether a language is local. Indeed, the previous results imply:
Claim 3.11.
Let $L$ a language, and $A$ its local overapproximation. Then we have $\mathcal{L}(A)=L$ if and only if $L$ is local.
Proof.
If $L$ is local, then $L$ is letter-Cartesian by Proposition 3.5 so by 3.10 we have $\mathcal{L}(A)=L$. Conversely, if $\mathcal{L}(A)=L$ then $L$ is recognized by a local DFA, so it is local. ∎
This result gives us a way to check whether the language recognized by an automaton is local. First, in the case of DFA, we can check this tractably:
Proposition 3.12 (Locality testing for DFAs).
Given a DFA $A$, it is in PTIME to determine if $\mathcal{L}(A)$ is local.
Proof.
Let $L\coloneq\mathcal{L}(A)$. Build the local overapproximation $A^{\prime}$ of $L$ using Definition 3.8, which is clearly in PTIME. By 3.11, it suffices to check whether $L=\mathcal{L}(A^{\prime})$. It is well-known that testing equivalence of two DFAs is PTIME, so this concludes the proof. ∎
Note that Proposition 3.12 is about testing if the language $L=\mathcal{L}(A)$ recognized by $A$ is local (i.e., some local DFA for $L$ exists); it is not about testing if the specific DFA $A$ is a local DFA (which is easy to do in linear time simply by following the definition). It is also not about the complexity of testing whether the infix-free language of $\mathcal{L}(A)$ is local, which can be the case even if $\mathcal{L}(A)$ is not local. We can do this by building an automaton $A^{\prime}$ representing $\mathtt{IF}(L)$ from a DFA $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, , Proposition 6)). We leave the precise complexity of this task open.
By contrast, if the input language $L$ is given as an NFA instead of a DFA, then testing if $L$ is local is PSPACE-complete (amarilli2025locality, ). This holds even under the assumption that $L$ is infix-free, so that the same lower bound applies to the problem of testing whether the infix-free language of an input NFA is local.
3.2. Tractability of Resilience for Local Languages
Having introduced local languages and presented their properties, we can now turn to the main result of this section, namely, that $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME when $L$ is local:
Theorem 3.13 (Resilience for local languages).
[…]
Before proving Theorem 3.13, we remark that it implies a PTIME bound on $\mathtt{RES}_{\mathrm{bag}}(L)$ whenever $\mathtt{IF}(L)$ is local (as $\mathtt{RES}_{\mathrm{bag}}(L)$ is equivalent to $\mathtt{RES}_{\mathrm{bag}}(\mathtt{IF}(L))$). For instance, $\mathtt{RES}_{\mathrm{bag}}(L_{0})$ for $L_{0}=\{a,aa\}$ is in PTIME because $\mathtt{IF}(L_{0})=\{a\}$ is local, even though $L_{0}$ is not. In fact, to see if Theorem 3.13 applies to a language $L$, we can always focus on $\mathtt{IF}(L)$, thanks to the following lemma (whose proof is deferred to LABEL:apx:qwerty):
Lemma 3.14 (Infix-free sublanguages preserve locality).
If $L$ is local then so is $\mathtt{IF}(L)$.
The proof of Theorem 3.13 is not difficult and is a simple reduction to the MinCut problem. To present it, we need a variant of local automata, which we name Read-Once (RO) automata. Intuitively, RO automata strengthen the requirement of local automata by requiring that for each letter $a\in\Sigma$ there is at most one transition labeled with $a$; but in exchange RO automata are not required to be deterministic and allow for (arbitrarily many) $\epsilon$-transitions. Formally:
[…]
Example 3.16 (Example 3.2 continued).
[…]
Proof.
We first prove the first sentence of the claim.
[…]
The second sentence follows from the construction given in 3.7, noticing that the sets $\Sigma_{\mathrm{start}}$, $\Sigma_{\mathrm{end}}$ and $\Pi$ can be computed in PTIME from $A$. ∎
[…]
Proof of Theorem 3.13.
[…]
  • The set of vertices is $(V\times S)\cup\{t_{\text{source}},t_{\text{target}}\}$, with the fresh vertices $t_{\text{source}}$ and $t_{\text{target}}$ being respectively the source and target of the network.
  • There is an edge with capacity $\mathtt{mult}(v\xrightarrow{a}v^{\prime})$ from $(v,s)$ to $(v^{\prime},s^{\prime})$ for all $v,v^{\prime}\in V$ and $s,s^{\prime}\in S$ and $a\in\Sigma$ such that $v\xrightarrow{a}v^{\prime}\in D$ and $(s,a,s^{\prime})\in\Delta$
  • There is an edge with capacity $\infty$ from $(v,s)$ to $(v,s^{\prime})$ for all $v\in V$ and $s,s^{\prime}\in S$ such that $(s,\epsilon,s^{\prime})\in\Delta$.
  • There is an edge with capacity $\infty$ from $t_{\text{source}}$ to every vertex of the form $(v,s)$ with $s\in I$, and from every vertex of the form $(v,s)$ with $s\in F$ to $t_{\text{target}}$.
Clearly $\mathcal{N}_{D,A}$ is built in $O(|A|\times|D|)$ from $A$ and $D$. Moreover, there is a one-to-one correspondence between the contingency sets of $D$ for $Q_{L}$ and the cuts of $\mathcal{N}_{D,A}$ that have finite cost: intuitively, selecting facts of $D$ in the contingency set corresponds to cutting edges in $\mathcal{N}_{D,A}$, and ensuring that the remaining facts do not satisfy $Q_{L}$ corresponds to ensuring that no source to target paths remain. Crucially, as $A$ is read-once, the correspondence between facts of $D$ and edges of $\mathcal{N}_{D,A}$ with finite capacity is one-to-one. Further, the sum of multiplicities of the facts of a contingency set is equal to the cost of the corresponding cut, so minimum cuts correspond to minimum contingency sets. We can thus solve MinCut on $\mathcal{N}_{D,A}$ in $\tilde{O}(|\mathcal{N}_{D,A}|)$ (henzinger2024deterministic, ) to compute the answer to $\mathtt{RES}_{\mathrm{bag}}(L)$ in $D$. ∎
4. Hardness Techniques and First Results
In this section, we move on to cases where computing resilience is intractable. From now on we only study languages that are infix-free, because showing hardness for an infix-free language $L$ is sufficient to imply the same for any language $L^{\prime}$ such that $\mathtt{IF}(L^{\prime})=L$.
We start the section by showing the NP-hardness of $\mathtt{RES}_{\mathrm{set}}(aa)$, by an explicit reduction from the minimum vertex cover problem. We then generalize this reduction and define the notion of (hardness) gadgets as databases that can be used to encode graphs and reduce from minimum vertex cover. We show that whenever a language $L$ admits a gadget then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard. We finally illustrate these techniques to show the hardness of resilience for the specific language $axb|cxd$.
The notion of gadgets introduced in this section will be used in the next two sections to show the hardness of entire language families. Specifically, we will generalize in Section 5 the hardness of $\mathtt{RES}_{\mathrm{set}}(axb|cxd)$ to that of the so-called four-legged languages; and generalize in Section 6 the hardness of $\mathtt{RES}_{\mathrm{set}}(aa)$ to that of all finite languages featuring a word with a repeated letter.
4.1. Hardness of aaaa
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20, , Prop. 10) by a direct reduction from 3-SAT. We give a self-contained proof where we reduce instead from minimum vertex cover: this proof serves as an intuitive presentation of the more general constructions that come next.
Proposition 4.1.
$\mathtt{RES}_{\mathrm{set}}(aa)$ is NP-hard.
In other words, given a directed graph $G$ and integer $k\in\mathbb{N}$, it is NP-hard to decide if we can remove $k$ edges of $G$ such that $G$ no longer contains any walk of length 2. We show Proposition 4.1 by reducing from the minimum vertex cover problem. Recall that a vertex cover of an undirected graph $G=(V,E)$ is a subset of vertices $V^{\prime}\subseteq V$ such that every edge of $E$ is incident to a vertex of $V^{\prime}$, i.e., for each $e\in E$, we have $e\cap V^{\prime}\neq\emptyset$. The vertex cover number of $G$ is the minimum cardinality of a vertex cover of $G$. Given an undirected graph $G$ and an integer $k\in\mathbb{N}$, the minimum vertex cover problem asks if the vertex cover number of $G$ is at most $k$. It is an NP-complete problem (karp1972reductibility, ).
Before showing the reduction, we show that the vertex cover number of an undirected graph $G$ can be computed from the vertex cover number of any odd-length subdivision of $G$. Formally, for $\ell\in\mathbb{N}$, $\ell\geq 1$, an $\ell$-subdivision of $G$ is an undirected graph $G^{\prime}$ obtained by replacing each edge of $G$ by a path of length $\ell$ (i.e., with $\ell-1$ fresh internal vertices, that are disjoint across edges). We claim:
Proposition 4.2 (Hardness of vertex cover on subdivisions).
Let $\ell\in\mathbb{N}$ be an odd integer. Then for any undirected graph $G=(V,E)$, letting $G^{\prime}$ be an $\ell$-subdivision of $G$, then the vertex cover number of $G^{\prime}$ is $k+m(\ell-1)/2$, where $k$ is the vertex cover number of $G$ and $m$ is the number of edges of $G$.
Proof.
It suffices to show that $G$ has a vertex cover of cardinality at most $k$ if and only if $G^{\prime}$ has a vertex cover of cardinality at most $k+m(\ell-1)/2$.
We first show the “only if” direction. Let $G=(V,E)$ and let $C$ be a vertex cover of $V$ of cardinality at most $k$. We pick a subset $C^{\prime}$ of the vertices of $G^{\prime}$ by taking the vertices of $C$ and picking vertices of the subdivisions in the following way. For each edge $\{u,v\}$ of $G$, letting $u-v_{1}-\cdots v_{\ell-1}-v$ be the corresponding path in $G$, we distinguish three cases:
  • If $u\in C$ and $v\notin C$ then we add to $C^{\prime}$ the vertices $v_{2},\ldots,v_{\ell-1}$
  • If $v\in C$ and $u\notin C$ then we add to $C^{\prime}$ the vertices $v_{1},\ldots,v_{\ell-2}$.
  • If $u\in C$ and $v\in C$ we do one of the above (it does not matter which).
The case where $u\notin C$ and $v\notin C$ is impossible because $C$ is a vertex cover of $G$.
Now, for each edge of $G$ we select $(\ell-1)/2$ vertices in $C^{\prime}$, so in total $C^{\prime}$ has the prescribed cardinality. To see that it is a vertex cover, note that all edges of $G^{\prime}$ occurs in a path of the form $u-v_{1}-\cdots v_{\ell-1}-v$, and the vertices of $C^{\prime}$ selected above form a vertex cover of the edges of such paths. Hence, $C^{\prime}$ is a vertex cover of $G^{\prime}$.
We next show the second direction, which is somewhat more challenging. Given a vertex cover $C^{\prime}$ of $G^{\prime}$ of cardinality at most $k+m(\ell-1)/2$, we will first modify $C^{\prime}$ without increasing its cardinality to guarantee the following property (*): for each edge $(u,v)\in E$, considering the corresponding path $u-v_{1}-\cdots v_{\ell-1}-v$ of $G^{\prime}$, then there are precisely $(\ell-1)/2$ of the intermediate vertices of the path that are in $C^{\prime}$, formally $|C^{\prime}\cap\{v_{1},\ldots,v_{\ell-1}\}|=(\ell-1)/2$.
For this, we process each path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, and consider each vertex $v_{1},\ldots,v_{\ell-1}$ in order. If for $i>1$ we have that $v_{i}\in C^{\prime}$ and the previous vertex $v_{i-1}$ was already in $C^{\prime}$, then we remove $v_{i}$ from $C^{\prime}$ and add instead $v_{i+1}$ to $C^{\prime}$. Note that the cardinality of $C^{\prime}$ does not increase, and may strictly decrease if $v_{i+1}$ was already in $C^{\prime}$. At the end of this process, given that we start with a vertex cover $C^{\prime}$ of $G^{\prime}$, we see that $C^{\prime}$ contains precisely $(\ell-1)/2$ vertices of the path: if $v_{1}\in C^{\prime}$ then we keep the $v_{i}$’s for which $i$ is odd, and otherwise the $v_{i}$’s for which $i$ is even.
We have processed $C^{\prime}$ to still be a vertex cover, and have cardinality $m(\ell-1)/2+k^{\prime}$ for some $k^{\prime}\leq k$, with $k^{\prime}=|C^{\prime}\cap V|$. Let us define $C\coloneq C^{\prime}\cap V$. It has cardinality at most $k$, and we claim that $C$ is a vertex cover. Indeed, assuming by contradiction that some edge $(u,v)$ of $G$ is not covered by $C$, so that $u\notin C$ and $v\notin C$, and considering the path $u=v_{0}-v_{1}-\cdots v_{\ell-1}-v_{\ell}=v$, we know that $C^{\prime}$ (after the processing of the previous paragraph) was a vertex cover of $G^{\prime}$, but if $u\notin C$ and $v\notin C$ then it cannot be the case that precisely $(\ell-1)/2$ of the $v_{i}$’s are in $C^{\prime}$ and cover all the edges of the path, a contradiction.
We have thus shown both directions of the claimed equivalence, which concludes the proof. ∎
Let us now show the hardness of resilience for the language $aa$:
Proof of Proposition 4.1.
vvuuwwvvuuww
(a) Undirected graph $G$ (top) and arbitrary orientation $G^{\prime}$ of $G$ (bottom)
(c) Graph of $aa$-matches of Figure 3(b)
(d) Database $\Xi$ constructed to encode $G^{\prime}$. All edges are labeled by $a$.
Figure 3. Illustration of the proof of Proposition 4.1.
(b) Completed gadget for aaaa
(a) Completed gadget for $axb|cxd$
(b) Hypergraph of matches of $axb|cxd$ on the completed gadget in Figure 4(a)
(c) Condensed hypergraph of matches of Figure 4(b)
Figure 4. Illustration of hardness gadgets and corresponding condensed hypergraphs of matches for the proof of Proposition 4.13.
We reduce from the vertex cover problem. Let $G$ be an undirected graph and $k\in\mathbb{N}$ be an integer. We perform the following construction, illustrated in Figure 3(a). We start by picking an arbitrary orientation for each edge of $G$, giving a directed graph $G^{\prime}=(V,E)$: see Figure 3(a). We then construct a database $\Xi$ in the following way. First, for each vertex $u\in V$ we create a fact $s_{u}\xrightarrow{a}t_{u}$. Second, for each edge $e=(u,v)$ of $E$ we create the following set $D_{e}$ of facts: $t_{u}\xrightarrow{a}t_{e,1}$, $t_{e,1}\xrightarrow{a}t_{e,2}$, $t_{e,2}\xrightarrow{a}t_{e,3}$, $t_{v}\xrightarrow{a}t_{e,2}$. Together with $s_{u}\xrightarrow{a}t_{u}$ and $s_{v}\xrightarrow{a}t_{v}$, the facts of $D_{e}$ form an instance that we call the completed gadget $D_{e}^{\prime}$ for $e$: we illustrate $D_{e}^{\prime}$ in Figure 3(b), abbreviating $t_{e,i}$ as $i$ for brevity. We introduce the notations $F_{e,\textup{in}}=s_{u}\xrightarrow{a}t_{u}$ and $F_{e,\textup{out}}=s_{v}\xrightarrow{a}t_{v}$ to refer to the two facts of $D_{e}^{\prime}\setminus D_{e}$, called the endpoint facts: they are shown in bold orange in Figure 3(b). The overall database $\Xi$ built from the orientation $G^{\prime}$ of $G$ is illustrated in Figure 3(d), where for brevity we abbreviate $t_{e,i}$ as $i_{e}$. Building $\Xi$ from $G$ is in $O(|G|)$ (as $D_{e}$ has constant size).
We then claim that the resilience of $aa$ on $\Xi$ is equal to the vertex cover number of a 5-subdivision of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-walk. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(c)). Thus, the graph of $aa$-matches of $\Xi$ is (isomorphic to) a 5-subdivision of $G$, and we conclude by Proposition 4.2. ∎
4.2. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(b) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, as in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd path. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
Definition 4.3 (Pre-gadget).
A pre-gadget is a tuple $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ consisting of a database $D$, two distinguished domain elements $t_{\textup{in}}\neq t_{\textup{out}}$ of $D$ respectively called the in-element and out-element, and a label $a\in\Sigma$. We require that $t_{\textup{in}}$ and $t_{\textup{out}}$ never occur as heads of facts of $D$. The completion of $\Gamma$ is the database $D^{\prime}$ obtained from $D$ by adding two fresh distinct domain elements $s_{\textup{in}}$ and $s_{\textup{out}}$ and the two facts $s_{\textup{in}}\xrightarrow{a}t_{\textup{in}}$ and $s_{\textup{out}}\xrightarrow{a}t_{\textup{out}}$ using letter $a$.
Example 4.4.
Figure 3(b) gives an example completion of a pre-gadget. The in-element and out-element of the pre-gadget are respectively $t_{u}$ and $t_{v}$, and the facts $s_{u}\xrightarrow{a}t_{u}$ and $s_{v}\xrightarrow{a}t_{v}$ are the facts added in the completion (so the pre-gadget has 4 edges and the completion has 6).
We next explain how we use gadgets to encode directed graphs (as in Figure 3(a)):
Definition 4.5 (Encoding a graph with a pre-gadget).
Given a directed graph $G$ and a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$, the encoding of $G$ with $\Gamma$ is the database $\Xi$ obtained from $G$ as follows:
  1. (1)
    For each node $u$ in $G$, create a fact $s_{u}\xrightarrow{a}t_{u}$ in $\Xi$, where $s_{u}$ and $t_{u}$ are fresh domain elements.
  2. (2)
    For each edge $e=(u,v)$ in $G$, add to $\Xi$ a copy $D_{e}$ of $D$ defined by the following isomorphism:
    • domain element $t_{\textup{in}}$ is renamed to $t_{u}$ and domain element $t_{\textup{out}}$ is renamed to $t_{v}$
    • all other elements are renamed to fresh elements called the internal elements of $D_{e}$.
Let us observe the following property about the encoding, which is the reason why we require that in-elements and out-elements never occur as the head of facts in pre-gadgets (before completion).
Claim 4.6 (Walks in encodings).
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a pre-gadget, and let $\Xi$ be the encoding of $G$ with $\Gamma$. Then no walk of $\Xi$ passes through two internal elements belonging to different copies of $D$.
Proof.
Consider a walk in $\Xi$ and consider the first element used in the walk which is an internal element of a copy $D_{e}$ of $D$. The only elements shared between $D_{e}$ and the rest of $\Xi$ are the images of the elements $t_{\textup{in}}$ and $t_{\textup{out}}$, i.e., elements of the form $t_{v}$ for $v$ a vertex of $D$. However, by our requirement on $D$, there is no fact in $D$ having $t_{\textup{in}}$ or $t_{\textup{out}}$ as head. Hence, the remainder of the walk can never reach the images of $t_{\textup{in}}$ and $t_{\textup{out}}$, so it can never reach an internal element of another copy. ∎
For now, our definition of pre-gadgets is independent of the query. Let us now formalize what we require from query matches in gadgets, in order to reduce from the vertex cover problem. We do this by introducing the hypergraph of matches, generalizing the graph of matches from the proof of Proposition 4.1, along with a condensation operation to simplify it.
4.3. Condensed Hypergraph of Matches
Let $L$ be a language and $D$ a database. A match (or witness) of $L$ on $D$ is a set of edges corresponding to an $L$-walk (as defined in Section 2): formally, each $L$-walk $e_{1},\ldots,e_{m}$ defines a match which is the set $\{e_{1},\ldots,e_{m}\}$, noting that several $L$-walks may define the same match. Then:
Definition 4.7 (Hypergraph of matches).
The hypergraph of matches of $D$ and $L$, denoted $\mathcal{H}_{L,D}$, is the hypergraph whose nodes are the facts of $D$ and whose hyperedges are the matches of $L$ on $D$.
Recall that a hitting set (also called hypergraph transversal (gottlob2004hypergraph, )) of a hypergraph $\mathcal{H}=(V,E)$ is a subset $V^{\prime}\subseteq V$ such that $V^{\prime}\cap e\neq\emptyset$ for every $e\in E$. Thus, hitting sets are a generalization of vertex covers to hypergraphs. It is clear from the definition of resilience that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of the hypergraph of matches $\mathcal{H}_{L,D}$. We now introduce two rewriting rules, called the condensation rules, that can be applied to any hypergraph $\mathcal{H}=(V,E)$ to simplify it without affecting the minimum size of hitting sets. For $v\in V$, write $E(v)=\{e\in E\mid v\in e\}$. The condensation rules are:
Edge-domination.:
If there are $e,e^{\prime}\in E$ with $e\neq e^{\prime}$ and $e\subseteq e^{\prime}$, we can remove $e^{\prime}$ to obtain $\mathcal{H}^{\prime}=(V,E\setminus\{e^{\prime}\})$. Intuitively, hitting sets must intersect $e$ so they automatically intersect $e^{\prime}$.
Node-domination.:
If there are $v,v^{\prime}\in V$ with $v\neq v^{\prime}$ such that $E(v)\subseteq E(v^{\prime})$, we can remove $v$ to obtain $\mathcal{H}^{\prime}=(V\setminus\{v\},\{e\setminus\{v\}\mid e\in E\})$. Intuitively, we can always replace $v$ by $v^{\prime}$ in hitting sets.
We say that $\mathcal{H}^{\prime}$ is a condensed hypergraph of $\mathcal{H}$ if $\mathcal{H}^{\prime}$ can be obtained from $\mathcal{H}$ by applying some sequence of condensation rules (note that we do not require that $\mathcal{H}^{\prime}$ cannot be further simplified using the rules). It can be remarked that the condensation rules are confluent in the sense that we always get isomorphic hypergraphs if we apply them to an input hypergraph until it is no longer possible, no matter the order in which we apply them: see (amarilli2025confluence, ). However, we will not need confluence for our purposes in the present paper. One can easily see that the rules preserve the minimum size of a hitting set:
Claim 4.8 (Condensation preserves minimum size of hitting sets).
Let $\mathcal{H}^{\prime}$ be a condensed hypergraph of $\mathcal{H}$. Then the minimum size of a hitting set of $\mathcal{H}$ and $\mathcal{H}^{\prime}$ are the same.
Proof.
We show the claim by induction on the length on the sequence of applications of the condensation rules: the base case, corresponding to the empty sequence, is trivial.
For the induction case, let us first assume that $\mathcal{H}^{\prime}$ is obtained from $\mathcal{H}=(V,E)$ by applying the edge-domination rule. We then immediately see that any $V^{\prime}\subseteq V$ is a hitting set of $\mathcal{H}$ iff it is a hitting set of $\mathcal{H}^{\prime}$, so the minimal size of a hitting set in $\mathcal{H}$ and $\mathcal{H}^{\prime}$ is the same.
Second, let us assume that $\mathcal{H}^{\prime}=(V^{\prime},E^{\prime})$ is obtained from $\mathcal{H}=(V,E)$ by applying the node-domination rule on two vertices $v$ and $v^{\prime}$ with $E(v)\subseteq E(v^{\prime})$. Write $V^{\prime}=V\setminus\{v\}$ and $E^{\prime}=\{e\setminus\{v\}\mid e\in E\}$. We show both directions of the claim. First, any hitting set of $\mathcal{H}^{\prime}$ is also a hitting set of $\mathcal{H}$, so the minimum size of a hitting set of $\mathcal{H}$ is at most the minimum size of a hitting set of $\mathcal{H}^{\prime}$. Second, let us consider a hitting set $S^{\prime}$ of $\mathcal{H}$. If $v\notin S^{\prime}$ then $S^{\prime}$ is a hitting set of $\mathcal{H}^{\prime}$. Otherwise, if $v\in S^{\prime}$, then $(S^{\prime}\setminus\{v\})\cup\{v^{\prime}\}$ is a hitting set of $\mathcal{H}^{\prime}$ of cardinality at most $|V^{\prime}|$. In both cases $\mathcal{H}^{\prime}$ has a hitting set of cardinality at most $|S^{\prime}|$, so that the minimum size of a hitting set of $\mathcal{H}^{\prime}$ is at most the minimum set of a hitting set of of $\mathcal{H}$. Hence, the equality is proven, which concludes the induction step and finishes the proof. ∎
By 4.8, we know that $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)$ is equal to the minimum size of a hitting set of any condensed hypergraph of $\mathcal{H}_{L,D}$, so we can always apply condensation rules in hardness proofs.
We can now define the requirement for a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ to be a gadget [^4] Note that a notion generalizing gadgets has been introduced under the name of Independent Join Paths in (makhija2024unified, ), where it is used to show the hardness of resilience of conjunctive queries.: we require that the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on the completion $D^{\prime}$ of $\Gamma$ can be condensed into an odd path. Formally, let $\mathcal{H}=(V,E)$ be a hypergraph and $u,v\in V$. We call $\mathcal{H}$ an odd path from $u$ to $v$ if all hyperedges of $\mathcal{H}$ have size 2 and the corresponding graph is an odd path from $u$ to $v$; i.e., there is a sequence $u=w_{1},\ldots,w_{2k}=v$ such that $E=\{\{w_{i},w_{i+1}\}\mid 1\leq i<2k\}$.
Definition 4.9 (Gadget).
A gadget for language $L$ is a pre-gadget $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ that satisfies the following condition: letting $D^{\prime}$ be the completion of $\Gamma$ (see Definition 4.3) where we add two $a$-facts $F_{\textup{in}}$ and $F_{\textup{out}}$ from fresh elements to $t_{\textup{in}}$ and $t_{\textup{out}}$ respectively, there exists a condensation of the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of $L$ on $D^{\prime}$ which is an odd path from $F_{\textup{in}}$ to $F_{\textup{out}}$.
Example 4.10.
The hardness pre-gadget whose completion is pictured in Figure 3(b) is a gadget for $aa$, as can be seen on the hypergraph of matches of Figure 3(c) (with no condensation rules applied).
We claim that the existence of a gadget implies that the resilience problem is hard, namely:
Proposition 4.11 (Gadgets imply hardness).
Let $L$ be an infix-free language. If there is a gadget $\Gamma$ for $L$, then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
All our NP-hardness results in the paper will be proved using Proposition 4.11, by building gadgets for the queries that we consider. We note that we build the gadgets by hand, so it is possible that some queries may have shorter gadgets than the ones that we exhibit.
We prove Proposition 4.11 in the remaining of Section 4.3, by reducing from the vertex cover problem. Let $G^{\prime}$ be an undirected graph, and let $G$ be an arbitrary orientation of $G$. We use the gadget to encode the input graph. The idea is that the definitions that we impose on gadgets are designed to ensure that the encoding of $G$ has a condensed hypergraph of matches $\mathcal{H}^{\prime}$ which is an odd-length subdivision of the original graph $G^{\prime}$, so that by Proposition 4.2 and 4.8 computing a minimum vertex cover on $G^{\prime}$ reduces to computing a minimum hitting set of $\mathcal{H}^{\prime}$. Let us prove this formally.
We call an odd subdivision of an undirected graph an $\ell$-subdivision for some odd $\ell\in\mathbb{N}$. The technical claim used in the proof of Proposition 4.11 is the following:
Claim 4.12.
Let $G$ be a directed graph, let $\Gamma=(D,t_{\textup{in}},t_{\textup{out}},a)$ be a gadget for $L$, and let $\Xi$ be the encoding of $G$ using $\Gamma$. Then there exists a condensed hypergraph of $\mathcal{H}_{L,\Xi}$ that has only hyperedges of size $2$ and which is (when seen as an undirected graph) an odd subdivision of $G^{\prime}$, the undirected graph corresponding to $G$.
Proof.
Thanks to 4.6, one can check that the hypergraph of matches $\mathcal{H}_{L,\Xi}$ is precisely the union of the hypergraphs of matches of the various copies $D_{e}$ of $D$, but where in- and out-elements have been merged as in Definition 4.5. Then, considering the sequence of condensation rules to be done to obtain the condensed hypergraph of matches of the completion $D^{\prime}$ of $D$ that is an odd path as in Definition 4.9, we can do this sequence of simplifications on every copy $D_{e}$ of the gadget $D$ (noticing that by Definition 4.9 the source and target facts are not affected by the condensation rules), thus indeed obtaining a condensed hypergraph of matches of $\mathcal{H}_{L,\Xi}$ that is an odd subdivision of $G^{\prime}$. ∎
We can now prove Proposition 4.11 from 4.12:
Proof of Proposition 4.11.
We reduce from the vertex cover problem. Let $\Gamma$ be a gadget for $L$. Let $G^{\prime}$ be an undirected graph, and pick an arbitrary orientation $G$ of $G^{\prime}$. Construct the encoding $\Xi$ of $G$ by $\Gamma$: clearly this can be done in linear time. Use the oracle to compute the resilience of $Q_{L}$ on $\Xi$, written $r$, which is equal to the minimum size of a hitting set in $\mathcal{H}_{L,\Xi}$. By 4.12, $\mathcal{H}_{L,\Xi}$ has a condensed hypergraph $\mathcal{H}^{\prime}$ that is an odd subdivision of $G^{\prime}$ and moreover by 4.8 we have that the size of a minimum hitting set of $\mathcal{H}^{\prime}$ is also equal to $r$. We can then conclude using Proposition 4.2 (noticing that the length of the odd subdivision is a constant that depends only on $L$). ∎
Hence, from now on, all hardness results in the paper will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo, ) to automatically perform some sanity checks on the gadgets that we construct. Namely, our implementation takes as input a description of a gadget along with a query. The query can be either explicit (e.g., $aa$, or $axb|cxd$), or it can be implicitly defined via requirements (e.g., words that are known to satisfy the query, or words for which it is known that no infix satisfies the query). The implementation then checks that the completion of the gadget satisfies the requirements outlined in Definition 4.3. Further, it exhaustively enumerates the directed paths in the gadget and checks which infixes of these paths are known to satisfy the query. This can be used to construct the condensed hypergraph of matches and check that it is indeed an odd path, to satisfy Definition 4.9. The implementation of (gadgetrepo, ) is also used to prepare the gadget drawings and condensed hypergraph of matches presented in the paper.
4.4. Hardness for $axb|cxd$
We illustrate the use of gadgets by building one for the language $axb|cxd$:
Proposition 4.13.
$\mathtt{RES}_{\mathrm{set}}(axb|cxd)$ is NP-hard.
Proof.
By Proposition 4.11, it is enough to exhibit a gadget for $axb|cxd$. We claim that the pre-gadget $\Gamma$ whose completion $D^{\prime}$ is depicted in Figure 4(a) is a gadget. First, $\Gamma$ is clearly a pre-gadget, so we only need to show that it is a gadget, i.e., that it satisfies Definition 4.9.
We show in Figure 4(b) the hypergraph of matches $\mathcal{H}_{L,D^{\prime}}$ of the gadget whose completion $D^{\prime}$ is depicted in Figure 4(a). We now explain why $\mathcal{H}_{L,D^{\prime}}$ can be condensed to the hypergraph of Figure 4(c), which is an odd path. In this case, it suffices to apply node-domination rules to $\mathcal{H}_{L,D^{\prime}}$ to obtain $\mathcal{H}^{\prime}$ by eliminating all vertices corresponding to facts that are not endpoint facts and occur in only one hyperedge. For instance, we can first remove $t_{out}\xrightarrow{x}15$ because it is dominated by $15\xrightarrow{b}16$, then remove $12\xrightarrow{x}15$ because it is dominated by $13\xrightarrow{x}12$, and so on, until we obtain $\mathcal{H}^{\prime}$.
Thus, $\Gamma$ is indeed a gadget, which concludes the proof. ∎
5. Hardness of Four-Legged Languages
In the previous section, we have presented hardness gadgets, and shown the hardness of resilience for the two specific languages $aa$ and $axb|cxd$. In this section we generalize the latter result to a subclass of the infix-free non-local languages, called the four-legged languages: these languages intuitively feature a specific kind of counterexample to locality. We show that every four-legged language admits a gadget, so that resilience for such languages is always NP-hard by Proposition 4.11. We give two consequences of this result: it implies that resilience is hard for all non-star-free regular languages; and it also implies a dichotomy on resilience for languages featuring a neutral letter.
5.1. Four-Legged Languages
Recall the notion of local languages (Definition 3.1) from Section 3, which hinges on recognizability by certain automata (namely, local DFAs), and which was equivalent to the notion of being letter-Cartesian (Proposition 3.5).
If a language $L$ is not letter-Cartesian, by definition there are $x\in\Sigma$ and $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ such that $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We then define the four-legged languages as those infix-free languages that feature such a violation in which none of the words $\alpha$, $\beta$, $\gamma$, and $\delta$ is empty. Formally:
Definition 5.1 (Four-legged).
A language $L$ is four-legged if it is infix-free and has a letter $x\in\Sigma$ (the body) and four non-empty words $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ (the legs) with $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$.
Example 5.2.
The languages $axb|cxd$ (from Proposition 4.13) and $axb|cxd|cxb$ are non-local and four-legged. The languages $aa$ (from Proposition 4.1) and $ab|bc$ are non-local but not four-legged.
Note that four-legged languages are never local, because they are not letter-Cartesian. The main result of this section is that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard when $L$ is four-legged; formally:
Theorem 5.3 (Hardness of four-legged languages).
If $L$ is four-legged then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
We prove Theorem 5.3 in the remaining of Section 5.1. To this end, we first show an intermediate lemma: for languages that are four-legged, we can assume without loss of generality that the legs are stable, i.e., no infix of $\alpha x\delta$ is in $L$:
Definition 5.4.
Let $L$ be a four-legged language, and let $x\in\Sigma$ be a body and let $\alpha,\beta,\gamma,\delta\in\Sigma^{+}$ be legs that witness that $L$ is four-legged, i.e., $\alpha x\beta\in L$ and $\gamma x\delta\in L$ but $\alpha x\delta\notin L$. We say that the legs $\alpha,\beta,\gamma,\delta$ are stable if, in addition, there is no infix of $\alpha x\delta$ which is in $L$.
Lemma 5.5.
Any infix-free language which is four-legged with body $x\in\Sigma$ also has stable legs with body $x$.
Proof.
Let $L$ be an infix-free language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $\eta^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $\eta^{\prime}$ is in $L$) then we are done. Otherwise, as $\eta^{\prime}\notin L$, it means that there is a strict infix $\eta$ of $\eta^{\prime}$ that is in $L$. We know that $\eta$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is infix-free. Similarly, $\eta$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $\eta=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $\eta=\eta^{\prime}$ but we assumed that $\eta$ is a strict infix of $\eta^{\prime}$.
We distinguish two cases:
  • We have $\delta_{2}\neq\epsilon$. Then we claim that $\alpha\coloneq\gamma^{\prime}$, $\beta\coloneq\delta^{\prime}$, $\gamma\coloneq\alpha_{1}$ and $\delta\coloneq\delta_{1}$ are stable legs. Indeed, observe first that none of $\alpha,\beta,\gamma,\delta$ is empty. Moreover we have $\alpha x\beta=\gamma^{\prime}x\delta^{\prime}\in L$ and $\gamma x\delta=\alpha_{1}x\delta_{1}\in L$. Last, no infix of $\alpha x\delta=\gamma^{\prime}x\delta_{1}$ can be in $L$ because $\gamma^{\prime}x\delta_{1}\delta_{2}=\gamma^{\prime}x\delta^{\prime}$ is in $L$ and $\delta_{2}\neq\epsilon$ so any infix of $\alpha x\delta$ is a strict infix of a word of $L$; but $L$ is infix-free so no strict infix of a word of $L$ is a word of $L$, hence no infix of $\alpha x\delta$ is a word of $L$.
  • We have $\alpha_{2}\neq\epsilon$. Then we do the symmetric choice of taking $\alpha\coloneq\alpha_{1},\beta\coloneq\delta_{1}$, $\gamma\coloneq\alpha^{\prime}$, and $\delta\coloneq\beta^{\prime}$ and claim that they are stable legs. Instead, none of $\alpha,\beta,\gamma,\delta$ is empty. Further, $\alpha x\beta=\alpha_{1}x\delta_{1}$ and $\gamma x\delta=\alpha^{\prime}x\beta^{\prime}$ are in $L$. Moreover we have $\alpha x\delta=\alpha_{1}x\beta^{\prime}$ and all its infixes are infixes of $\alpha^{\prime}x\beta^{\prime}$ which are strict because $\alpha_{2}\neq\epsilon$; as $\alpha^{\prime}x\beta^{\prime}\in L$ and $L$ is infix-free, no infix of $\alpha x\delta$ is a word of $L$.
Hence, in either case, we can define stable legs for $L$ with body $x$, which concludes the proof. ∎
Figure 5. Completed gadget for Case 1 of Theorem 5.3.
Figure 6. Completed gadget for Case 2 of Theorem 5.3.
We can now prove that resilience is hard for four-legged languages (Theorem 5.3):
Proof of Theorem 5.3.
Let $L$ be an infix-free and four-legged language. Using Lemma 5.5, let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be four stable legs, i.e., we have that $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but no infix of $\alpha^{\prime}x\delta^{\prime}$ is in $L$. We then distinguish two cases:
Case 1.:
The first case is when, furthermore, no infix of $\gamma^{\prime}x\beta^{\prime}$ is in $L$. Then in this case, let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$, which is well defined since $\alpha^{\prime}\neq\epsilon$) and likewise write $\beta^{\prime}=\beta b$ and $\gamma^{\prime}=c\gamma$ and $\delta^{\prime}=\delta d$ (note that some of the letters $a$, $b$, $c$, and $d$ may be equal). In this case we claim that the (completed) gadget depicted in Figure 5 is a valid gadget. [^5] We point out that this gadget is essentially the generalization of the gadget in Figure 4(a) for $L=axb|cxd$. Edges in the gadgets labeled with words $\alpha,\beta,\gamma,\delta\in\Sigma^{*}$ represent paths labeled with the letters in the words. If any of the words is empty, then the head node of the edge is merged with the tail node. Such a drawing allows us to show hardness of any language that can be obtained by replacing the $\alpha,\beta,\gamma,\delta$ with any words in $\Sigma^{*}$. This is possible since, for any language that meets the conditions of case 1, a condensed hypergraph of matches can be obtained that does not contain any vertices corresponding to a word (i.e., a Greek symbol). We can then check exhaustively that all directed walks in the completed gadget are either accounted for by the condensed hypergraph of matches or are not in the language. Thus, we know that for any such language, a condensed hypergraph of matches will be an odd path as shown in Figure 5, and hence the gadget is a valid gadget.
Case 2.:
The second case is when there are some infixes of $\gamma^{\prime}x\beta^{\prime}$ that are in $L$. Observe that such infixes must intersect non-trivially both $\gamma^{\prime}$ and $\beta^{\prime}$. Indeed, infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\gamma^{\prime}$ are strict infixes of $\alpha^{\prime}x\beta^{\prime}\in L$ so they are not in $L$ because $L$ is infix-free, and likewise infixes of $\gamma^{\prime}x\beta^{\prime}$ which do not intersect $\beta^{\prime}$ are strict infixes of $\gamma^{\prime}x\delta^{\prime}\in L$ so again they are not in $L$. Let us write $\alpha^{\prime}=a\alpha$ (i.e., $a$ is the first letter of $\alpha^{\prime}$) and $\beta^{\prime}=b\beta$ ($b$ is the first letter of $\beta^{\prime}$) and $\gamma^{\prime}=c_{1}\gamma_{1}$ ($c_{1}$ is the first letter of $\gamma^{\prime}$) and $\gamma^{\prime}=\gamma_{2}c_{2}$ ($c_{2}$ is the last letter of $\gamma^{\prime}$) and $\delta^{\prime}=\delta d$ ($d$ is the last letter of $\delta^{\prime}$). Then we claim that the (completed) gadget depicted in Figure 6 is a valid gadget. Indeed, we can again check exhaustively all directed walks in the gadget and observe that the following holds for all of them: they (1) either contain or are subsumed by $\alpha^{\prime}x\beta^{\prime}$ or $\gamma^{\prime}x\delta^{\prime}$ matches; (2) are an infix of $\alpha^{\prime}x\delta^{\prime}$ (and hence can never be a match); or (3) contain $c_{2}xb$ as an infix. We notice that any infix of $\gamma^{\prime}x\beta^{\prime}$ that is a match must contain $c_{2}xb$ as an infix, and we know that such a match exists. For any such match, the other letters of the match can be eliminated in the condensed hypergraph of matches, as they are subsumed (via the edge domination rule) by the $c_{2}xb$ - thus the node domination rule can remove the other letters of the $\gamma^{\prime}x\beta^{\prime}$-infix match. Thus, it suffices to consider $\alpha^{\prime}x\beta^{\prime}$, $\gamma^{\prime}x\delta^{\prime}$, and $c_{2}xb$ as the only matches in the condensed hypergraph of matches. When we do so, we can see that there exists a condensed hypergraph of matches that is an odd path as shown in Figure 6, and hence the gadget is valid.
This concludes the proof of Theorem 5.3. ∎
Note that Theorem 5.3 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not in NP).
5.2. Implications of Hardness for Four-Legged Languages
Next, we spell out two consequences of Theorem 5.3. The first one concerns non-star-free languages, the second concerns languages with a neutral letter.
Hardness for non-star-free languages.
We say that a regular language $L$ is star-free if there exists a $k>0$ such that, for every $\rho,\sigma,\tau\in\Sigma^{*}$, for every $m\geq k$, we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{m}\tau\in L$. Star-free languages have many equivalent characterizations (see, e.g., (mpri, )), including being counter-free (mcnaughton1971counter, ). Further, if a language $L$ is star-free then so is $\mathtt{IF}(L)$ (see LABEL:apx:starfreered). We can show that being star-free is a necessary condition to enjoy tractable resilience, because non-star-free infix-free regular languages are four-legged and hence resilience for them is NP-hard by Theorem 5.3:
Lemma 5.6 (Hardness of non-star-free languages).
Let $L$ be a regular language which is infix-free and non-star-free. Then $L$ is four-legged.
Proof.
Let $A$ be a DFA which recognizes $L$, and let $k>0$ be greater than the number of states of $A$. From the fact that $L$ is not star-free, applying the definition to this choice of $k$, we obtain $\rho,\sigma,\tau\in\Sigma^{*}$ and $m\geq k$ such that exactly one of $\rho\sigma^{k}\tau$ and $\rho\sigma^{m}\tau$ is in the language. In particular this implies that $\sigma\neq\epsilon$; let us write $\sigma=x\sigma^{\prime}$.
As $k$ is greater than the number of states of $A$, by the pigeonhole principle, there must be $0\leq i<j\leq k$ such that the partial runs of $\rho\sigma^{i}$ and $\rho\sigma^{j}$ finish at the same state when read by $A$. Let $\omega\coloneq j-i$. We know that we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+\omega}\tau\in L$, because exchanging the prefixes $\rho\sigma^{i}$ and $\rho\sigma^{j}$ does not change the state that we reach. Iterating the argument, we get the following fact, dubbed (*): for any $p\geq 0$ we have $\rho\sigma^{k}\tau\in L$ iff $\rho\sigma^{k+p\omega}\tau\in L$. For similar reasons, we get the fact dubbed (**): for any $p\geq 0$ we have $\rho\sigma^{m}\tau\in L$ iff $\rho\sigma^{m+p\omega}\tau\in L$.
Let $(\phi,\psi)=(k,m)$ if $\rho\sigma^{k}\tau\in L$ and $\rho\sigma^{m}\tau\notin L$, and $(\phi,\psi)=(m,k)$ if it is the opposite. Recall that $m,k>0$ so $\phi>0$ and $\psi>0$. Further, by (*) and (**) we know the following fact, dubbed ($\dagger$): we have $\rho\sigma^{\phi+p\omega}\tau\in L$ for all $p\geq 0$, and $\rho\sigma^{\psi+p\omega}\tau\notin L$ for all $p\geq 0$. Let $p$ be large enough so that $\phi+p\omega-1>\psi$. Now let us take:
  • $\alpha=\rho\sigma^{2\omega-1}$,
  • $\beta=\sigma^{\prime}\sigma^{\phi}\tau$,
  • $\gamma=\rho\sigma^{\phi+p\omega-1-\psi}$ and
  • $\delta=\sigma^{\prime}\sigma^{\psi}\tau$.
Note that all of $\alpha,\beta,\gamma,\delta$ are non-empty because they all contain $v\neq\epsilon$ at a power which is greater than 0. We claim that they witness that $L$ is four-legged with body $x$. Indeed:
  • $\alpha x\beta=\rho\sigma^{2\omega-1}x\sigma^{\prime}\sigma^{\phi}\tau=\rho\sigma^{2\omega+\phi}\tau$
  • $\gamma x\delta=\rho\sigma^{\phi+p\omega-1-\psi}x\sigma^{\prime}\sigma^{\psi}\tau=\rho v^{\phi+p\omega}\tau$
  • $\alpha x\delta=\rho\sigma^{2\omega-1}x\sigma^{\prime}\sigma^{\psi}\tau=\rho\sigma^{2\omega+\psi}\tau$.
So by ($\dagger$) we have that:
  • $\alpha x\beta\in L$ because $\rho\sigma^{\phi}\tau\in L$
  • $\gamma x\delta\in L$ because $\rho\sigma^{\phi}\tau\in L$
  • $\alpha x\delta\notin L$ because $\rho\sigma^{\psi}\tau\notin L$
This establishes that $L$ is four-legged. As it is infix-free, we can conclude by Proposition 5.3. ∎
Dichotomy for languages with a neutral letter.
A second consequence of Theorem 5.3 concerns languages $L$ with a neutral letter, i.e., a letter $e\in\Sigma$ such that inserting or deleting $e$ anywhere does not change the membership of words to $L$. Formally, $e$ is neutral for $L$ if for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\beta\in L$ iff $\alpha e\beta\in L$. In formal language theory, one often assumes the existence of a neutral letter as a technical assumption on languages (koucky2005bounded, ; barrington2005first, ). Under this assumption, we can show:
Proposition 5.7 (Dichotomy for languages with neutral letters).
Let $L$ be a language with a neutral letter. If $\mathtt{IF}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME, otherwise $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Note that we can read this result as implying that, in the presence of a neutral letter, there are not many tractable cases left. Indeed, intuitively, when considering a language $L$ with neutral letter whose infix-free sublanguage is local, then membership to $L$ must be guided only by the first and last non-neutral letters and the set of non-neutral letters that occur in the word.
We prove Proposition 5.7 in the remaining of Section 5.2. To prove this result, we first show that, for languages featuring a neutral letter, whenever the infix-free language is non-local then it must either be four-legged or contain a word with a repeated letter, specifically a word of the form $aa$ for $a\in\Sigma$:
Lemma 5.8.
Let $L$ be a language with a neutral letter such that $\mathtt{IF}(L)$ is not local. Then one of the following two cases holds (possibly both):
  • $\mathtt{IF}(L)$ is four-legged;
  • there is $x\in\Sigma$ such that $xx\in\mathtt{IF}(L)$.
To prove this lemma, let us prove the following intermediate result:
Claim 5.9.
Let $L$ be a language with a neutral letter $e$ and let $L^{\prime}=\mathtt{IF}(L)$. Then for any non-empty words $\theta,\eta\in\Sigma^{+}$, if $\theta\eta\in L^{\prime}$ then $\theta e\eta\in L^{\prime}$.
Proof.
We have $\theta\eta\in L^{\prime}$ so $\theta\eta\in L$, and $e$ is neutral for $L$ so we have $\theta e\eta\in L$. Let us proceed by contradiction and assume that $\theta e\eta\notin L^{\prime}$. This means that there is a strict infix $\tau$ of $\theta e\eta$ which is in $L$. We first claim that $\tau$ must include the $e$. Indeed, if it does not, then there are two cases:
  • $\tau$ is an infix of $\theta$. Then it is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the fact that $L^{\prime}$ is infix-free, so this case is impossible.
  • $\tau$ is an infix of $\eta$. This is impossible for the same reason as the previous case.
Now, we know that $\tau$ is a strict infix of $\theta e\eta$ which includes the $e$. Let $\tau^{\prime}$ be obtained from $\tau$ by removing the $e$. We have $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$. What is more, it is easily seen that $\tau^{\prime}$ is an infix of $\theta\eta$, and it is a strict infix because $\tau$ includes the $e$ so from $|\tau|<|\theta e\eta|$ (because $\tau$ is a strict infix) we deduce $|\tau^{\prime}|=|\tau|-1<|\theta e\eta|-1=|\theta\eta|$. So $\tau^{\prime}\in L$ is a strict infix of $\theta\eta\in L^{\prime}$, which contradicts the assumption that $L^{\prime}$ is infix-free. ∎
We can now prove Lemma 5.8:
Proof of Lemma 5.8.
[…]
Now, because $L^{\prime}$ is infix-free, it is easy to check that we must have the following claim, dubbed ($\star$): we have $\alpha^{\prime}\beta^{\prime}\neq\epsilon$ and $\gamma^{\prime}\delta^{\prime}\neq\epsilon$ and $\alpha^{\prime}\gamma^{\prime}\neq\epsilon$ and $\beta^{\prime}\delta^{\prime}\neq\epsilon$. Indeed:
  • If $\alpha^{\prime}\beta^{\prime}=\epsilon$ then $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ is equal to $x$ which is an infix of $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$, so by infix-freeness it must be the case that $\gamma^{\prime}\delta^{\prime}=\epsilon$, but now $x=\alpha^{\prime}x\delta^{\prime}=x$ and we assumed $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$, contradiction.
  • If $\gamma^{\prime}\delta^{\prime}=\epsilon$, the reasoning is similar: in this case $\gamma^{\prime}x\delta^{\prime}$ is equal to $x$, which is an infix of $\alpha^{\prime}x\beta^{\prime}$, so by infix-freeness we must have $\alpha^{\prime}\beta^{\prime}=\epsilon$, but then $x=\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ and $x=\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ is a contradiction.
Thus, we have established ($\star$).
We now start proving the statement. Let us first show that (call this $\dagger$) if both $\alpha^{\prime}$ and $\delta^{\prime}$ are empty then the result holds. Indeed, in that case we have $x\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}x\in L^{\prime}$. By ($\star$) (first two points) we have that $\beta^{\prime}$ and $\gamma^{\prime}$ are both non-empty. By 5.9 we deduce from $x\beta^{\prime}\in L^{\prime}$ that $xe\beta^{\prime}\in L^{\prime}$, and we deduce from $\gamma^{\prime}x\in L^{\prime}$ that $\gamma^{\prime}ex\in L^{\prime}$. Now, take $e$ as a candidate choice of body, and $\alpha=x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x$, as a candidate choice of legs: all of these are non-empty. We have $\alpha e\beta\in L^{\prime}$ and $\gamma e\delta\in L^{\prime}$. So, considering the word $\alpha e\delta=xex$, there are two possibilities: either $xex\notin L^{\prime}$ or $xex\in L^{\prime}$. In the first case, we conclude with our choice of body and legs that $L^{\prime}$ is four-legged, which allows us to conclude. In the second case, we have $xx\in L$ because $e$ is neutral for $L$. Further, we have $xx\in L^{\prime}$ because the strict infixes of $xx$ are the empty word and $x$ which are both strict infixes of the word $x\beta^{\prime}\in L^{\prime}$: we conclude because $L^{\prime}$ is infix-free. The fact that $xx\in L^{\prime}$ allows us to conclude: indeed, we can then do the exact same reduction from vertex cover as in Proposition 4.1 (noticing that we do not use other letters than ’$x$’).
The same reasoning shows that (call this $\dagger^{\prime}$) if both $\beta^{\prime}$ and $\gamma^{\prime}$ are empty then the result holds. Indeed, in this case we have $\alpha^{\prime}x\in L^{\prime}$ and $x\delta^{\prime}\in L^{\prime}$, and we have $\alpha^{\prime}$ and $\delta^{\prime}$ both non-empty by ($\star$) (first two points). Applying 5.9 to each gives us that $\alpha^{\prime}ex\in L^{\prime}$ and $xe\delta^{\prime}\in L^{\prime}$, and again we conclude that $L^{\prime}$ is four-legged or that $xex\in L^{\prime}$ from which we show $xx\in L^{\prime}$ in the same way.
We have ruled out the case of $\alpha^{\prime}$ and $\delta^{\prime}$ being both empty, and of $\beta^{\prime}$ and $\gamma^{\prime}$ being both empty. Combining with $(\star)$, one gets that either $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty, or $\beta^{\prime}$ and $\delta^{\prime}$ are both non-empty. Indeed, assume that the first term in the disjunction is false, i.e., that it is not the case that $\alpha^{\prime}$ and $\gamma^{\prime}$ are both non-empty. If we have $\alpha^{\prime}=\epsilon$, then $\delta^{\prime}$ is non-empty from ($\dagger$), and $\beta^{\prime}$ is non-empty by $(\star)$ (first point). If we have $\gamma^{\prime}=\epsilon$, then $\beta^{\prime}$ is non-empty by ($\dagger^{\prime}$), and $\delta^{\prime}$ is non-empty by $(\star)$ (second point).
We finish the proof by distinguishing these two cases:
  • Assume that $\alpha^{\prime}\neq\epsilon$ and $\gamma^{\prime}\neq\epsilon$. We have $\alpha^{\prime}x\beta^{\prime}\in L^{\prime}$ so by 5.9 we have $\alpha^{\prime}ex\beta^{\prime}\in L^{\prime}$, and we have $\gamma^{\prime}x\delta^{\prime}\in L^{\prime}$ so by 5.9 we have $\gamma^{\prime}ex\delta^{\prime}\in L^{\prime}$. We then claim that $\alpha^{\prime}ex\delta^{\prime}\notin L^{\prime}$, implying that $L^{\prime}$ is four-legged (taking $e$ as the body, $\alpha=\alpha^{\prime}$, $\beta=x\beta^{\prime}$, $\gamma=\gamma^{\prime}$ and $\delta=x\delta^{\prime}$, which are all not empty). Indeed, assume by way of contradiction that $\alpha^{\prime}ex\delta^{\prime}\in L^{\prime}$. We would then have $\alpha^{\prime}ex\delta^{\prime}\in L$, so $\alpha^{\prime}x\delta^{\prime}\in L$ (as $e$ is neutral). As $\alpha^{\prime}x\delta^{\prime}\notin L^{\prime}$ by assumption, there exists a strict infix $\tau$ of $\alpha^{\prime}x\delta^{\prime}$ that is in $L^{\prime}$. Now, adding $e$ to $\tau$ would give a strict infix $\tau^{\prime}$ of $\alpha^{\prime}ex\delta^{\prime}$, with $\tau^{\prime}\in L$ because $\tau\in L$ and $e$ is neutral for $L$, a contradiction because $L^{\prime}$ is infix-free.
  • Assume that $\beta^{\prime}\neq\epsilon$ and $\delta^{\prime}\neq\epsilon$. The reasoning is similar: using 5.9 we have $\alpha^{\prime}xe\beta^{\prime}\in L^{\prime}$ and $\gamma^{\prime}xe\delta^{\prime}\in L^{\prime}$ and we obtain that $L^{\prime}$ is four-legged by taking $e$ as the body, $\alpha=\alpha^{\prime}x$, $\beta=\beta^{\prime}$, $\gamma=\gamma^{\prime}x$ and $\delta=\delta^{\prime}$.
This concludes the proof. ∎
For completeness we give here examples of languages for which only one of the two cases in the statement of Lemma 5.8 holds. Let $L_{1}=e^{*}be^{*}ce^{*}|e^{*}de^{*}fe^{*}$ (for which $e$ is neutral), with $\mathtt{IF}(L_{1})=be^{*}c|de^{*}f$. Then $\mathtt{IF}(L_{1})$ is four-legged (taking $x=e$, $\alpha=b$, $\beta=c$, $\gamma=d$, $\delta=f$), hence $\mathtt{IF}(L_{1})$ is not local, but it does not have a word of the form $xx$ for $x\in\Sigma$. Now take $L_{2}=e^{*}(a|c)e^{*}(a|d)e^{*}$, with $\mathtt{IF}(L_{2})=(a|c)e^{*}(a|d)$. Then $\mathtt{IF}(L_{2})$ is not local (taking $x=a$, $\alpha=c$, $\beta=\epsilon$, $\gamma=\epsilon$, $\delta=d$), it contains the word $aa$, but one can check that it is not four-legged.
We can then use Lemma 5.8 to prove our complexity characterization of resilience for languages featuring a neutral letter (Proposition 5.7):
Proof of Proposition 5.7.
If $\mathtt{IF}(L)$ is local then $\mathtt{RES}_{\mathrm{bag}}(L)$ is PTIME by Theorem 3.13. Otherwise, by Lemma 5.8 either $\mathtt{IF}(L)$ is four-legged, in which case $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard by Theorem 5.3, or $aa\in\mathtt{IF}(L)$ for $a\in\Sigma$. In this case one can easily adapt the proof of Proposition 4.1: one can check that the reduction only builds instances using the letter $a$, and the restriction of $\mathtt{IF}(L)$ to words using only the letter $a$ must be exactly $aa$ given that $\mathtt{IF}(L)$ is infix-free. This concludes the proof. ∎
The dichotomy above relies on the presence of a neutral letter, so it does not apply to all languages. In particular, it does not apply at all to finite languages. This motivates further study of finite languages, and in fact all the languages studied in the remainder of the paper will be finite languages.
6. Hardness of Finite Languages Having Repeated Letters
We now conclude the presentation of our main hardness results by showing that resilience is hard for a class of languages that generalizes $aa$ (Proposition 4.1). We say that a word $\alpha$ has a repeated letter if we have $\alpha=\beta a\gamma a\delta$ for some $a\in\Sigma$ and $\beta,\gamma,\delta\in\Sigma^{*}$. We show that, for infix-free finite languages, resilience is NP-hard whenever the language contains a word with a repeated letter:
Theorem 6.1 (Hardness with repeated letters).
Let $L$ be an infix-free finite language that contains a word with a repeated letter. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-complete.
Before proving Theorem 6.1, we discuss its relation with existing work and with the other results we have obtained so far.
Connections to UCQs and self-joins.
Observe that finite languages $L$ correspond to the case of RPQs that are in fact unions of conjunctive queries (UCQs). The UCQs obtained from RPQs have a specific form: they are posed on an arity-two signature corresponding to the alphabet $\Sigma$, and their constituent conjunctive queries (CQs) are directed paths corresponding to words of the language $L$. In this sense, repeated letters in words correspond to self-joins in CQs. Thus, Theorem 6.1 implies that, for those UCQs that correspond to finite RPQs, the resilience problem is NP-hard unless every CQ is self-join-free. This contrasts with the setting of general UCQs, where resilience is sometimes tractable even in the presence of self-joins (DBLP:conf/pods/FreireGIM20, ). Specifically, in that setting, the following Boolean CQ on an arity-two signature is known to be in PTIME (DBLP:conf/pods/FreireGIM20, ) despite having two self-joins: Q:A(x),B(x),R(x,y),R(z,y),A(z),C(z)Q{\,:\!\!-\,}A(x),B(x),R(x,y),R(z,y),A(z),C(z).
Connections to other language classes.
We now explain how Theorem 6.1 relates to the languages discussed so far. We first perform a sanity check: such languages are never local.
Lemma 6.2.
Let $L$ be a finite language containing a word with a repeated letter. Then $L$ is not local.
Proof.
As $L$ is finite and contains a word with a repeated letter, we can take a word with a repeated letter whose length is maximum among words of $L$ with a repeated letter. Let $\alpha=\beta a\gamma a\delta\in L$ be such a word. Assume by contradiction that $L$ is local. Then, by Proposition 3.5, $L$ is letter-Cartesian. Applying the definition to $\alpha$ and $\alpha$, we must have that the following word is also in $L$:
$\beta a\gamma a\gamma a\delta\in L.$
This word is a word of $L$ which contains a repeated letter and is strictly longer than $\alpha$. This contradicts the maximality of $\alpha$, and concludes the proof. ∎
Note that the finiteness hypothesis is essential for Lemma 6.2, e.g., $L=ax^{*}b$ contains words with repeated letters but it is local (see Example 3.2). Further notice that there is no converse to Lemma 6.2: there are finite infix-free non-local languages that feature no words with repeated letters and for which resilience is hard. This is for instance the case of the four-legged language $axb|cxd$ of Proposition 4.13. Thus, the languages covered by Theorem 6.1 and Theorem 5.3 are incomparable: they each generalize Proposition 4.1 and Proposition 4.13, respectively, without implying the other result. (However, the proof of Theorem 6.1 relies on Theorem 5.3.) These two results also do not cover all intractable cases, as we will see in the next sections.
We prove Theorem 6.1 in the remaining of this section. Because the proof is technical, we sketch it here before giving the details. We first make a preliminary observation: the tractability of resilience is preserved under the mirror operation on regular languages. We then start from a word featuring a repeated letter, and pick it to optimize certain criteria (maximizing the gap between the repeated letters, and then maximizing the total length of the word): we call this a maximal-gap word, and the existence of such words is the only place where we use the fact that $L$ is finite.
We then consider such a maximal-gap word $\beta a\gamma a\delta$, and consider various cases. We first exclude the case where $\beta$ and $\delta$ are both non-empty, by showing that $L$ is then four-legged. Then, thanks to the mirror operation, we assume without loss of generality that $\delta$ is empty. We then consider the word $\gamma a\gamma$ and study whether it has infixes in $L$. If it does not, we easily build a gadget. If it does, we can now show that $\beta$ must be empty or otherwise the language is four-legged. We then study the choice of infix $\alpha^{\prime}$ of $\gamma a\gamma$ and distinguish two cases by considering the suffix of the left $\gamma$ and the prefix of the right $\gamma$: either they overlap or they don’t. We build suitable gadgets for each case.
The mirror operation.
We first make a straightforward observation on the mirror of languages. The mirror of a word $\alpha=a_{1}\cdots a_{n}\in\Sigma^{*}$ is $\alpha^{\mathrm{R}}=a_{n}\cdots a_{1}\in\Sigma^{*}$ (in particular $\epsilon^{\mathrm{R}}=\epsilon$). The mirror of a language $L\subseteq\Sigma^{*}$ is $L^{\mathrm{R}}\coloneq\{\alpha^{\mathrm{R}}\mid\alpha\in L\}$. Observe that $(L^{\mathrm{R}})^{\mathrm{R}}=L$. It is immediate by symmetry that a language $L$ is infix-free iff $L^{\mathrm{R}}$ is infix-free. Further, it is also immediate that a language and its mirror have the same complexity for resilience:
Proposition 6.3.
Let $L$ be any language. There are PTIME reductions between $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$ and $\mathtt{RES}_{\mathrm{set}}(L)$, and between $\mathtt{RES}_{\mathrm{bag}}(L^{\mathrm{R}})$ and $\mathtt{RES}_{\mathrm{bag}}(L)$.
The point of this claim will be to eliminate symmetric cases, up to replacing the language by its mirror. We only use the claim about set semantics in the sequel (given that we will show hardness of resilience in set semantics, from which hardness for bag semantics follows): the claim for bag semantics in Proposition 6.3 is given just for completeness.
Proof of Proposition 6.3.
We give the proof for set semantics, the proof for bag semantics is the same mutatis mutandis. Given a database instance $D$ for $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$, simply create a database instance $D^{\mathrm{R}}$ by reversing the direction of all arrows of $D$. It is then obvious that a subset $D^{\prime}\subseteq D$ satisfies $Q_{L^{\mathrm{R}}}$ iff the corresponding subset $(D^{\prime})^{\mathrm{R}}$ of $D^{\mathrm{R}}$, which has the same cardinality, satisfies $Q_{L}$. Hence, the answer to $\mathtt{RES}_{\mathrm{set}}(L^{\mathrm{R}})$ on $D$ and to $\mathtt{RES}_{\mathrm{set}}(L)$ on $D^{\mathrm{R}}$ are identical, giving the forward PTIME reduction. The converse reduction is obtained by applying the previous argument to reduce $(L^{\mathrm{R}})^{\mathrm{R}}=L$ to $L^{\mathrm{R}}$. ∎
Maximal-gap words
Let $L$ be an infix-free finite language that contains a word with a repeated letter, and let us show that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard, to establish Theorem 6.1. We will pick a word with a repeated letter for the rest of the proof, but we will choose such a word carefully. Here is the precise notion we will need:
Definition 6.4.
For $L$ a finite language, we say that a word $\alpha=\beta a\gamma a\delta$ that features a repeated letter is a maximal-gap word if:
  • The gap between the repeated letters is maximal, i.e., there is no word $\alpha^{\prime}=\beta^{\prime}a^{\prime}\gamma^{\prime}a^{\prime}\delta^{\prime}$ in $L$ such that $|\gamma^{\prime}|>|\gamma|$
  • Among such words, the length of $\alpha$ is maximal, i.e., there is no word $\alpha^{\prime}=\beta^{\prime}a^{\prime}\gamma^{\prime}a^{\prime}\delta^{\prime}$ in $L$ such that $|\gamma^{\prime}|=|\gamma|$ and $|\alpha^{\prime}|>|\alpha|$.
It is clear that any finite language that contains a word with a repeated letter must have maximal-gap words. Hence, in the rest of the proof, we fix $\alpha=\beta a\gamma a\delta$ to be an arbitrary such word. We will do case disjunctions on this word in the rest of the proof.
Reducing to the case $\beta=\epsilon$.
Let us first assume that $\beta$ and $\delta$ are both non-empty. In this case we conclude because $L$ is four-legged. Indeed:
Claim 6.5.
Let $L^{\prime}$ be an infix-free language containing a maximal-gap word $\beta a\gamma a\delta$ such that $\beta\neq\epsilon$ and $\delta\neq\epsilon$. Then $L^{\prime}$ is four-legged.
Proof.
Consider the two decompositions of the word $\alpha$ as $\alpha=(\beta a\gamma)a\delta$ and $\alpha=\beta a(\gamma a\delta)$. Note that $\beta a\gamma$, $\delta$, $\beta$, and $\gamma a\delta$ are all non-empty, so they can be used as legs. So consider the word $\alpha^{\prime}=(\beta a\gamma)a(\gamma a\delta)$. The word $\alpha^{\prime}$ contains two $a$’s separated by $\gamma a\gamma$, which is longer than $\gamma$. Hence, $\alpha^{\prime}$ is not in $L^{\prime}$, as it would contradict the fact that $\alpha$ is maximal-gap (the first criterion of the definition of maximal-gap suffices). So we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
If $\beta$ and $\delta$ are both non-empty, then applying 6.5 and Theorem 5.3 with $L^{\prime}\coloneq L$ concludes. Hence, in the rest of the argument, we assume that one of $\beta$ and $\delta$ is empty. Further, without loss of generality, up to replacing $L$ by $L^{\mathrm{R}}$ using Proposition 6.3, we assume that $\beta=\epsilon$. Hence, in the rest of the proof, we can assume that $\beta=\epsilon$: we have a maximal-gap word $\alpha=a\gamma a\delta\in L$ and this word is maximal-gap.
Reducing to the case where some infix of $\gamma a\gamma$ is in $L$.
Continuing the proof, we will now consider the word $\gamma a\gamma$ and study whether some infixes of this word are in $L$. (Intuitively, when designing a gadget for $L$, such paths will appear, and so the construction of the gadget will be different depending on whether these paths constitute query matches or not.) Let us first assume that no infix of $\gamma a\gamma$ is in $L$: this case can easily be ruled out by designing a gadget for $L$ based on the word $\alpha=a\gamma a\delta$:
Lemma 6.6.
Let $L^{\prime}$ be an infix-free language containing a word $a\gamma a\delta$ where no infix of $\gamma a\gamma$ is in $L^{\prime}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 7. Completed gadget for the proof of Lemma 6.6 in the case when $\delta=\epsilon$ (left), and condensed hypergraph of matches (right).
Figure 8. Completed gadget for Lemma 6.6 in the case where $\delta\neq\epsilon$ (left), and condensed hypergraph of matches (right).
Proof.
We first present the case where $\delta=\epsilon$, and then present the case where $\delta\neq\epsilon$.
In the case where $\delta=\epsilon$, we use the completed gadget pictured on Figure 7 (left), where the edge labeled $\gamma$ stands for a (possibly empty) path. Note that, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as their head (except the two facts of the gadget completion, pictured in orange), so indeed it is a pre-gadget. The condensed hypergraph of matches is shown on Figure 7 (right), and it is an odd path, so indeed the instance pictured on Figure 7 (left) is (the completion of) a gadget and we can conclude by Proposition 4.11. Note how we use the assumption that the matches on the completed gadget are the ones indicated: directed walks including two $a$’s must include $a\gamma a$ as an infix and so they are dominated, and directed walks including only one $a$ cannot be a match of the query.
In the case where $\delta\neq\epsilon$, we use the completed gadget pictured on Figure 8 (left), again with the edge labeled $\gamma$ standing for a possibly empty path. Again, no matter what $\gamma$ stands for, there will be no facts having $t_{u}$ and $t_{v}$ as head (except for the facts added in the completion), so again we have a pre-gadget. The condensed hypergraph of matches is shown on Figure 8 (right). Now, to show that it is actually like this, first observe that the facts of the $\delta$-walks are always dominated by the preceding $a$: there is no match of the query within $\delta$ (as it is a strict infix of $a\gamma a\delta$), so all matches using $\delta$ must use the preceding $a$, and we can remove the facts of $\delta$-walks in the condensation via the node domination rule. We are then essentially back to the preceding case. More precisely, again, no directed walk containing only one copy of the $a$-edges of the gadget can be a word of the language, because these paths are either infixes of $\gamma a\delta$ (hence strict infixes of $a\gamma a\delta\in L$) or they are infixes of $\gamma a\gamma$ and so they are not in the language by assumption. Further, for any two consecutive $a$’s in the gadget, there is a $a\gamma a\delta$ match containing them. Further, after observing that $\delta$ facts are eliminated by node domination, we note that all query matches that are infixes of a word of the form $a(\gamma a)^{*}$ must contain two consecutive $a$’s, so they are subsumed by the matches corresponding to $a\gamma a\delta$, which after node domination contain just two consecutive $a$’s and the $\gamma$ between them. (The facts corresponding to $\gamma$ are then eliminated by node domination.)
Hence, we indeed have a condensed hypergraph of matches corresponding to Figure 8 (right), and we can conclude by Proposition 4.11. ∎
Applying this lemma with $L^{\prime}\coloneq L$ concludes. So in the rest of the proof we can assume that there is an infix of $\gamma a\gamma$ which is in $L$.
Reducing to the case $\delta=\epsilon$.
The maximal-gap word $\alpha=a\gamma a\gamma\delta$, together with the infix of $\gamma a\gamma$ which is in $L$, now allows us to show that $L$ is four-legged unless $\delta=\epsilon$. To show this, we need to discuss the structure of the infix of $\gamma a\gamma$:
Claim 6.7.
Let $L^{\prime}$ be an infix-free language containing the word $\alpha=a\gamma a\delta$ and containing an infix of $\gamma a\gamma$. Then $L^{\prime}$ contains a word of the form $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ where $\gamma_{1}$ is a non-empty suffix of $\gamma$ and $\gamma_{2}$ is a non-empty prefix of $\gamma$.
Proof.
Consider the infix of $\gamma a\gamma$ which exists by hypothesis. This infix must include the middle $a$, as otherwise it is an infix of $\gamma$, which is a strict infix of $\alpha\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is infix-free.
So write the infix as $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ being a (not necessarily strict) suffix of $\gamma$ and $\gamma_{2}$ being a (not necessarily strict) prefix of $\gamma$. All that remains is to show that $\gamma_{1}$ and $\gamma_{2}$ are non-empty.
First, we must have $\gamma_{1}\neq\epsilon$, otherwise $\alpha^{\prime}$ is an infix of $a\gamma$ which is a strict infix of $a\gamma a\delta\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is infix-free.
Second, we must have $\gamma_{2}\neq\epsilon$ otherwise $\alpha^{\prime}$ is an infix of $\gamma a$ which is a strict infix of $a\gamma a\delta$ and we conclude in the same way. ∎
Hence, applying 6.7 to $L^{\prime}\coloneq L$, we know that $L$ contains a maximal-gap word $\alpha=a\gamma a\delta$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ and $\gamma_{2}$ being respectively a non-empty suffix of $\gamma$ and a non-empty prefix of $\gamma$. We now claim that, if $\delta\neq\epsilon$, then $L$ is four-legged, allowing us to conclude:
Claim 6.8.
Let $L^{\prime}$ be an infix-free language containing a word $a\gamma a\delta$ such that $\delta\neq\epsilon$ and a word $\gamma_{1}a\gamma_{2}$ such that $\gamma_{1}\neq\epsilon$, $\gamma_{2}\neq\epsilon$, and $\gamma_{1}$ is a suffix of $\gamma$. Then $L^{\prime}$ is four-legged.
Proof.
Indeed, consider $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ and $\alpha=(a\gamma)a\delta$. The words $\delta$, $a\gamma$, $\gamma_{1}$, and $\gamma_{2}$ are all non-empty, so they can be used as legs. So consider the word $\alpha^{\prime\prime}=\gamma_{1}a\delta$ and show that $\alpha^{\prime\prime}\notin L^{\prime}$, so that $L^{\prime}$ is four-legged.
Indeed, the word $\alpha^{\prime\prime}$ is an infix of $\gamma a\delta$ (because $\gamma_{1}$ is a suffix of $\gamma$), which is a strict infix of $a\gamma a\delta$ which is in $L^{\prime}$. Thus, as $L^{\prime}$ is infix-free, $\alpha^{\prime\prime}\notin L^{\prime}$, so we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
Applying 6.8 with $L^{\prime}\coloneq L$, we can conclude unless $\delta=\epsilon$, which we assume in the rest of the proof.
Distinguishing two cases depending on whether $\gamma_{1}$ and $\gamma_{2}$ overlap.
To summarize, we now know that $L$ contains a maximal-gap word $\alpha=a\gamma a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$ for $\gamma_{1}$ a non-empty suffix of $\gamma$ and $\gamma_{2}$ a non-empty prefix of $\gamma$. We distinguish two cases, intuitively based on whether $\gamma_{1}$ and $\gamma_{2}$ cover $\gamma$ in a way that overlaps or not:
Non-overlapping case.:
The non-overlapping case is when $|\gamma_{1}|+|\gamma_{2}|\leq|\gamma|$, so that we can write $\gamma=\gamma_{2}\eta\gamma_{1}$ for some $\eta\in\Sigma^{*}$.
Overlapping case.:
The overlapping case is when we have $|\gamma_{1}|+|\gamma_{2}|>|\gamma|$, so that we can write $\gamma=\eta^{\prime\prime}\eta\eta^{\prime}$ such that $\gamma_{1}=\eta\eta^{\prime}$ and $\gamma_{2}=\eta^{\prime\prime}\eta$, with the “overlap” $\eta$ being non-empty.
Showing hardness in the overlapping case.
Let us first deal with the overlapping case. We have $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, meaning $\alpha^{\prime}=\eta\eta^{\prime}a\eta^{\prime\prime}\eta\in L$, and we know that the “overlap” $\eta$ is non-empty. We then show:
Claim 6.9.
If the words $\eta^{\prime}$ and $\eta^{\prime\prime}$ are not both empty, then $L$ is four-legged.
Proof.
Let us first assume that $\eta^{\prime}$ is non-empty and show that $L$ is four-legged in this case. Let us write $\eta^{\prime}=x\sigma$. Substituting this into $\alpha^{\prime}$ and $\alpha$, we have $\alpha^{\prime}=\eta x(\sigma a\eta^{\prime\prime}\eta)\in L$ and $\alpha=(a\eta^{\prime\prime}\eta)x(\sigma a)\in L$. Now, taking $x$ as body, let us consider the cross-product word $\alpha^{\prime\prime}=\eta x\sigma a=\eta\eta^{\prime}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta\eta^{\prime}a\in L$. Hence, as $L$ is infix-free, $\alpha^{\prime\prime}\notin L$, so that we can conclude that $L$ is four-legged unless $\eta^{\prime}$ is empty.
Let us substitute $\eta^{\prime}=\epsilon$ in $\alpha$ and $\alpha^{\prime}$. We know that $\alpha^{\prime}=\eta a\eta^{\prime\prime}\eta\in L$ and $\alpha=a\eta^{\prime\prime}\eta a\in L$. Let us now assume that $\eta^{\prime\prime}$ is non-empty and show that $L$ is four-legged in that case as well. Let us write $\eta^{\prime\prime}=x\sigma$. We have $\alpha=ax(\sigma\eta a)\in L$ and $\alpha^{\prime}=(\eta a)x(\sigma\eta)\in L$. Now, taking $x$ as body again, let us consider the cross-product word $\alpha^{\prime\prime}=ax\sigma\eta=a\eta^{\prime\prime}\eta$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta a\in L$. Hence, as $L$ is infix-free, $\alpha^{\prime\prime}\notin L$, so we can conclude again that $L$ is four-legged unless $\eta^{\prime\prime}$ is empty.
We conclude the claim: $L$ is four-legged unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$. ∎
From 6.9, we can conclude unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$, so we assume this in the rest of the proof of hardness for the overlapping case. We now know that $\alpha^{\prime}=\eta a\eta\in L$ and $\alpha=a\eta a\in L$.
Let us now re-use the fact that $\alpha$ was chosen to be a maximal-gap word of $L$. Now, as $\eta$ is non-empty, observe that $\alpha^{\prime}=\eta a\eta$ contains repeated letters across the two copies of $\eta$. These letters are spaced at least by the length of $\eta$ (and possibly more, if $\eta$ contains repeated letters – even though this cannot happen by the maximal-gap condition). In any case, $\alpha^{\prime}$ achieves a gap between two repeated letters which is at least that of $\alpha$: so $\alpha$ and $\alpha^{\prime}$ are tied for the first criterion in the definition of maximal-gap. Hence, by the second criterion, since $\alpha$ is maximal-gap, we know that $\alpha^{\prime}$ cannot be strictly longer than $\alpha$. So from $|\alpha^{\prime}|\leq|\alpha|$, as $|\alpha^{\prime}|=2|\eta|+1$ and $|\alpha|=|\eta|+2$, we conclude $|\eta|\leq 1$. As $\eta$ is non-empty, we have $|\eta|=1$. This letter may be different from $a$ or not. So we can conclude the overlapping case by showing two hardness results:
Claim 6.10.
Let $L^{\prime}$ be any infix-free language containing abaaba and babbab. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is hard.
Figure 9. Completed gadget used in the proof of 6.10
Proof.
We use the gadget whose completion is pictured on Figure 9 (left). It is obviously a pre-gadget, and one can check that the hypergraph of matches can be reduced to the odd path of Figure 9 (right), so that it is indeed a gadget and we can conclude by Proposition 4.11. More precisely, observe that there are three kinds of directed walks in the gadget:
  • Paths of length 2 or less: all of these form a word which is a strict infix of abaaba or babbab, so they are not in the language by infix-freeness.
  • Paths of length 3: all of these form the word abaaba or babbab, so they are in the language.
  • Paths of length 4 or more: all of these have a path of length 3 as infix, and these always form words of the language by the previous point, so they are all subsumed by edge domination.
Hence, the hypergraph of matches of the language $L^{\prime}$ on the completion of the gadget corresponds precisely to the hyperedges given by the paths abaaba and babbab. One can then obtain the condensation of Figure 9 (right) by observing that $5\xrightarrow{b}1$ is subsumed by the next $a$-fact ($1\xrightarrow{a}2$) and that $8\xrightarrow{b}7$ is subsumed by the next $a$-fact ($7\xrightarrow{a}4$). The matches $\{1\xrightarrow{a}2,2\xrightarrow{b}3\}$ and $\{7\xrightarrow{a}4,4\xrightarrow{b}6\}$ then eliminate by edge domination the matches $\{t_{u}\xrightarrow{b}1,1\xrightarrow{a}2,2\xrightarrow{b}3\}$ and $\{1\xrightarrow{a}2,2\xrightarrow{b}3,3\xrightarrow{a}4\}$ and $\{t_{v}\xrightarrow{b}7,7\xrightarrow{a}4,4\xrightarrow{b}6\}$, and we can then see that $3\xrightarrow{a}4$ is eliminated by node domination (e.g., by $2\xrightarrow{b}3$) and that $t_{u}\xrightarrow{b}1$ is eliminated by node domination with $1\xrightarrow{a}2$ and $t_{v}\xrightarrow{b}7$ is eliminated by node domination with $7\xrightarrow{a}4$, leading to the pictured hypergraph. This allows us to conclude by Proposition 4.11. ∎
Claim 6.11.
Let $L$ be any infix-free language containing the word aaaaaa. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Figure 10. Completed gadget used in the proof of 6.11
Proof.
We use the gadget whose completion is pictured on Figure 10 (left) (which incidentally turns out to be identical to that of Figure 3(b) in Proposition 4.1). This is indeed a pre-gadget, and the hypergraph of matches can be condensed to Figure 10 (right), so it is indeed a gadget. Note that we use the fact that all directed walks in the gadget form words of the form $a^{*}$: by infix-freeness of the language, the matches precisely correspond to the words of length 3. ∎
Showing hardness in the non-overlapping case.
We conclude the proof of Theorem 6.1 by covering the non-overlapping case. Recall that, in this case, we know that the language $L$ contains a maximal-gap word $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, and recall that we already know that $\gamma_{1}$ and $\gamma_{2}$ are both non-empty.
We show:
Claim 6.12.
The language $L$ is four-legged unless $\gamma_{1}$ and $\gamma_{2}$ both have length 1.
Proof.
As $\gamma_{1}$ is non-empty, let us write $\gamma_{1}=\chi^{\prime}x$. If $\gamma_{1}$ has length 2 or more, then $\chi^{\prime}$ is non-empty, and we have $\alpha^{\prime}=\chi^{\prime}x(a\gamma_{2})\in L$ and $\alpha=(a\gamma_{2}\eta\chi^{\prime})xa\in L$. Let us take $x$ as body and consider the cross-product word $\alpha^{\prime\prime}=\chi^{\prime}xa=\gamma_{1}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$. So, as $L$ is infix-free, we have $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged. Hence, it suffices to focus on the case where $\gamma_{1}=x$ has length 1.
Now, as $\gamma_{2}$ is non-empty, let us write $\gamma_{2}=y\chi^{\prime\prime}$. If $\gamma_{2}$ has length 2 or more, then $\chi^{\prime\prime}$ is non-empty, and we have $\alpha=ay(\chi^{\prime\prime}\eta\gamma_{1}a)\in L$ and $\alpha^{\prime}=(\gamma_{1}a)y\chi^{\prime\prime}\in L$. Let us take $y$ as body and consider the cross-product word $\alpha^{\prime\prime}=ay\chi^{\prime\prime}=a\gamma_{2}$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha\in L$, so as $L$ is infix-free we know that $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged.
Hence, we conclude that $L$ is four-legged unless $\gamma_{1}$ and $\gamma_{2}$ both have length 1. ∎
Thanks to 6.12, we know that $\alpha=ax\eta ya\in L$ and $\alpha^{\prime}=yax\in L$, for some letters $x,y\in\Sigma$ and some word $\eta\in\Sigma^{*}$. So all that remains is to show:
Claim 6.13.
Let $L^{\prime}$ be an infix-free language that contains words $ax\eta ya$ and yaxyax for some $a,x,y\in\Sigma$ and $\eta\in\Sigma^{*}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 11. Completed gadget used in the proof of 6.14
To show this claim, we make an auxiliary claim:
Claim 6.14.
Let $L^{\prime}$ be an infix-free language that contains a word of the form aabaab for $a,b\in\Sigma$ with $a\neq b$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Proof.
We use the gadget whose completion is pictured on Figure 11 (left). Let us now discuss the condensed hypergraph of matches. First observe that there are three matches forming the word aabaab, and walks of length at most 2 must form strict infixes of aabaab so they do not correspond to words of $L^{\prime}$. Note that, as $ab$ is not in $L^{\prime}$ (as a strict infix of $aab\in L^{\prime}$), then we know that every match using $3\xrightarrow{a}1$ must also use the fact $t_{v}\xrightarrow{a}3$; hence $3\xrightarrow{a}1$ can be eliminated by node-domination. Likewise, $b$ is not in $L^{\prime}$, so every match using $3\xrightarrow{b}4$ must also use $t_{v}\xrightarrow{a}3$, and $3\xrightarrow{b}4$ is eliminated by node domination. This ensures that the only walk of length greater than 3 which does not contain aabaab as an infix, namely, the aaaaaa path from $s_{v}$ to $1$, is subsumed in the condensation in the case where $aaa\in L^{\prime}$: namely, after eliminating $3\xrightarrow{a}1$ by node elimination, the remaining facts $\{s_{v}\xrightarrow{a}t_{v},t_{v}\xrightarrow{a}3\}$ correspond to the match $\{s_{v}\xrightarrow{a}t_{v},t_{v}\xrightarrow{a}3,3\xrightarrow{b}4\}$ after we have used node domination to eliminate $3\xrightarrow{b}4$. Thus, we can build a condensed hypergraph of matches by considering only the walks that form the word aabaab, and then applying condensation we obtain Figure 11 (right). This establishes that we have a gadget, so we can conclude by Proposition 4.11. ∎
We can now conclude the proof of 6.13:
Figure 12. Completed gadget used in the proof of 6.13
Proof of 6.13.
We first rule out the case where $y=a$. Then $L^{\prime}$ contains $yax=aax$, and we conclude directly by 6.14 or 6.11 depending on whether $x=a$ or not.
We next rule out the case where $x=a$. Then, using Proposition 6.3, let us show hardness for $L^{\prime\mathrm{R}}$. This language contains xayxay, and as $x=a$ we can apply 6.14 or 6.11 again depending on whether $y=a$ or not, and conclude.
Hence, in the sequel, we assume that $x$ and $y$ are both distinct from $a$. (We may have $x=y$ or not.)
We then use the gadget whose completion is pictured in Figure 12 (left), with a condensed hypergraph of matches shown on figure Figure 12 (right). The same construction, and the same condensed hypergraph of matches, is used no matter whether $x$ and $y$ are the same letter or not.
To see why the condensed hypergraph of matches is as pictured on Figure 12, let us describe the directed walks in the completion of the gadget. The ones forming words of the form $ax\eta ya$, and yaxyax, are guaranteed to be matches: the ones forming strict infixes of these are guaranteed not be matches because the language is infix-free; and the ones containing one of these as an infix are guaranteed to be eliminated by node domination. In particular, walks that start within an $\eta$-edge, or end within an $\eta$-edge, or both, are not relevant. Indeed, observe that $\eta$-edges are always followed by (some prefix of) $yax\eta$. So a match starting within an $\eta$-edge (possibly using all facts of the edge) must include the subsequent yaxyax (because $\eta ya$ is a strict infix of $ax\eta ya$ and so is not in the language), and the match is then subsumed by the yaxyax-match. Likewise, $\eta$-edges are always preceded by (some suffix of $\eta yax$), so matches ending within an $\eta$-edge (possibly using all facts of the edge) must include the preceding yaxyax (because $ax\eta$ is a strict infix of $ax\eta ya$) and then they are subsumed by the yaxyax match.
For this reason, to build the condensed hypergraph of matches, it suffices to consider walks that start and end at the edges labeled $a$ and $x$ and $y$. One can check that such walks either are infixes of $ax\eta ya$ or contain yaxyax as an infix, so their membership to $L$ is determined by the fact that $L$ contains yaxyax and $ax\eta ya$ and is infix-free. Building the corresponding condensed hypergraph of matches and further applying condensation rules gives Figure 12 (right), which is an odd path, so we conclude by Proposition 4.11. ∎
Applying 6.13 with $L^{\prime}\coloneq L$, we conclude that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard in the remaining cases of the non-overlapping case. This concludes the case distinction and concludes the proof of Theorem 6.1.
7. Additional Complexity Results
We have shown the tractability of resilience for local languages (Theorem 3.13), and shown hardness for four-legged languages (Theorem 5.3) and for finite infix-free languages featuring a word with a repeated letter (Theorem 6.1). Our results imply a dichotomy over languages with a neutral letter (Proposition 5.7), but not in general. In this section, we classify some languages not covered by these results, and highlight remaining open cases. We only consider languages that are infix-free, and mostly study finite languages.
We first illustrate two cases of languages that are tractable but not local. One case are the languages we call bipartite chain languages, such as $ab|bc$, which are finite languages intuitively obtained from a bipartite graph on letter endpoints. For them, we can compute resilience via a variant of the product construction of Theorem 3.13, simply by “reversing” some matches. Another case are the languages we call one-dangling languages, such as $abc|be$, which are intuitively obtained by extending a (possibly infinite) infix-free local language with a single two-letter word with which it shares at most one letter. For them, we can again compute resilience by a more involved encoding to a flow problem. We conclude by showing examples of queries not covered by our previous hardness results but for which resilience can be shown intractable.
7.1. Tractability of Bipartite Chain Languages
Let us define the chain languages:
Definition 7.1 (Chain languages).
A language $L$ is a chain language if:
  1. (1)
    No word in $L$ contains a repeated letter.
  2. (2)
    Each word of $L$ of length at least 2 may only share its first or last letter with other words in $L$: formally, for each word $\alpha=a\gamma b$ in $L$, no letter of $\gamma$ occurs in another word of $L$.
Note that chain languages are always finite, because they contain at most $|\Sigma|$ words of length >2>2: indeed, each word of length 3 or more contains intermediate letters which cannot occur in another word. We will focus on chain languages satisfying an extra condition, called being bipartite:
Definition 7.2 (Endpoint graph, BCL).
The endpoint graph of a language $L$ is the undirected graph $(\Sigma,E)$ having as vertices the letters of $\Sigma$, and having an edge $\{a,b\}$ for $a\neq b$ in $\Sigma$ iff $a$ and $b$ are the endpoints of some word of $L$ of length at least $2$ (i.e., $L$ contains a word of the form $a\alpha b$ or $b\alpha a$).
A chain language is a bipartite chain language (BCL) if its endpoint graph is bipartite.
Example 7.3 (Figure 14).
The languages $ab|bc$ (from Example 5.2) and $axyb|bztc|cd|dea$ are BCLs; the language $ab|bc|ca$ is a chain language but not a BCL. None of these languages are local.
Thus, there are two kinds of chain languages: those that are bipartite, and those which are not. For chain languages $L$ that are not bipartite, we conjecture that $\mathtt{RES}_{\mathrm{set}}(L)$ is always NP-hard. We leave this open, but show hardness for the specific non-bipartite chain language of Example 7.3:
Proposition 7.4.
The resilience problem for the language $ab|bc|ca$ is NP-hard.
Proof.
Figure 13. Completed gadget used in the proof of Proposition 7.4
We claim that the gadget shown in Figure 13 is valid for $ab|bc|ca$. Indeed, the gadget has the following properties: We can see that it meets all the criteria of being a pre-gadget. We can also generate a condensed hypergraph of matches that forms an odd path, and hence this gadget is a valid hardness gadget for $ab|bc|ca$. ∎
Figure 14. Illustrations of the endpoint graphs for the three languages mentioned in Example 7.3.
We now focus on chain languages that are bipartite, and we will show tractability of resilience for such languages in the rest of this section. Notice that the definition of BCLs is symmetric: any word $\alpha$ of $L$ can be replaced by the mirror word of $\alpha$ and the resulting language is still a BCL. Also notice that the class of BCLs is incomparable to local languages (even finite local languages): indeed, $ax^{*}b$ and $axb|axc$ are local languages but not BCLs, and Example 7.3 shows BCLs that are not local. Further note that BCLs containing no word of length less than 2 are always infix-free. Further, we have the immediate analogue of Lemma 3.14, which we state here and prove in LABEL:apx:redbcl2:
Lemma 7.5.
lem:redbcl22 For any BCL $L$, the language $\mathtt{IF}(L)$ is also a BCL.
We can now state our tractability result: the resilience problem is tractable for all BCLs, even in combined complexity and in bag semantics:
Proposition 7.6 (Bipartite chain languages are tractable).
[…]
Lemma 7.7.
[…]
We can then prove Proposition 7.6 from LABEL:lem:nfachain:
Proof of Proposition 7.6.
Thanks to LABEL:lem:nfachain, we compute in time $O(|A|\times|\Sigma|^{2})$ the language $L\coloneq\mathcal{L}(A)$ as an explicit list of words.
We first process $L$ in linear time to eliminate the words of length 0 or $1$. If $L$ contains $\epsilon$, then resilience is trivial; so in the sequel we assume that $L$ does not contain $\epsilon$. For every letter $a\in\Sigma$ such that $L$ contains the single-letter word $a$, then we know that in the input database $D$ we must remove every $a$-fact for the query to be false; so we can preprocess $D$ to remove all such facts and remove the corresponding one-letter words from $L$. Hence, at the end of this preprocessing, in time $O(|A|+|D|)$, we can assume that $L$ consists of words of length 2 or more.
We now point out that $L$ is infix-free. This is by definition of a chain language and from the assumption that the words have length at least 2. Indeed, if $\alpha\in L$ is a strict infix of $\beta\in L$ then $\beta$ must have length at least 3, so there are letters in $\beta$ that are neither the first nor the last letter, by definition of $L$ being a chain language these letters cannot occur in a word of $L\setminus\{\beta\}$, so they do not occur in $\alpha$, so $\alpha$ being an infix of $\beta$ implies that $\alpha$ has length $1$, contradicting the assumption that the words of $L$ have length 2 or more.
We call an endpoint letter a letter of $\Sigma$ which occurs as the first or last letter of a word of $L$, i.e., has some incident edges in the endpoint graph $G$ of $L$. The set of such letters can be computed in linear time from the explicit list. We can also build explicitly the endpoint graph $G$ of $L$ from the explicit list, in linear time in the list. We know by assumption that $G$ is bipartite: we compute a witnessing bipartition in linear time by the obvious greedy algorithm. Let us assign each endpoint letter of $L$ to either the source or target partition. Now, the words of $L$ can be partitioned between the forward words, in which the first letter is assigned to the source partition and the last letter to the target partition; and the reversed words, where the opposite occurs.
Then we construct a flow network as follows:
  • We add two fresh vertices $t_{\text{source}}$ and $t_{\text{target}}$ which are the source and the target respectively of the network.
  • For every fact $a(v,v^{\prime})\in D$ we define two vertices $\mathrm{start}_{a,v,v^{\prime}}$ and $\mathrm{end}_{a,v,v^{\prime}}$.
  • For every letter $a$, there is an edge with capacity $\mathtt{mult}(a(v,v^{\prime}))$ from $\mathrm{start}_{a,v,v^{\prime}}$ to $\mathrm{end}_{a,v,v^{\prime}}$ for every $v,v^{\prime}$ such that we have $a(v,v^{\prime})\in D$.
  • For every forward word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity $\infty$ from $\mathrm{end}_{a,v,v^{\prime}}$ to $\mathrm{start}_{b,v^{\prime},v^{\prime\prime}}$
  • For every reversed word containing an infix $ab$, where $a,b$ are letters in $\Sigma$, and $a(v,v^{\prime})$ and $b(v^{\prime},v^{\prime\prime})\in D$, there is an edge with capacity $\infty$ from $\mathrm{end}_{b,v^{\prime},v^{\prime\prime}}$ to $\mathrm{start}_{a,v,v^{\prime}}$.
  • There is an edge with capacity $\infty$ from $t_{\text{source}}$ to every vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is an endpoint assigned to the source partition, and from every vertex of the form $\mathrm{end}_{a,v,v^{\prime}}$ to $t_{\text{target}}$ where $a$ is an endpoint in the target partition.
The explicit list of words has size $O(|A|\times|\Sigma|^{2})$, and the construction runs in time $O(|D|^{2}\times|A|\times|\Sigma|^{2})$.
We can see that this construction contains a single edge for every fact, whose capacity is the multiplicity of said fact. Every witness of the instance corresponds to a source-to-target walk by construction. Indeed, for each forward word, we have a path from $t_{source}$ to the edge corresponding to its first letter, then following each letter in succession until we reach the last letter and then $t_{target}$. Further, for each backward word, we have a path from $t_{source}$ to the edge corresponding to its last letter, then following each letter in succession in reverse order, until we reach the first letter and $t_{target}$.
We must also show that there is no path from source to target in the network that does not correspond to a witness. Any path from $t_{source}$ to $t_{target}$ must start with a vertex of the form $\mathrm{start}_{a,v,v^{\prime}}$ where $a$ is assigned to the source partition (in particular it is not a letter that occurs in the middle of any word). We then move to a vertex $\mathrm{start}_{b,v^{\prime},v^{\prime\prime}}$ and there are two cases: either $a$ was the first letter of a forward word $\alpha$ and $ab$ occurs as an infix in that word, necessarily at the beginning because no middle letters are repeated, or $a$ was the last letter of a backward word $\alpha$ and $ba$ occurs in that word, again necessarily at the end because no middle letters are repeated. We describe the first case (corresponding to forward words), the second (corresponding to backward words) is symmetric. From the condition that intermediate letters of words do not occur in other words, whenever we reach a vertex, then the next vertex must correspond to the next letter of $\alpha$. Eventually we reach the last letter of $\alpha$ and the corresponding node has no outgoing edge in the flow network except the edge to $t_{\text{target}}$, so we terminate. Thus, all paths in this network correspond to witnesses.
Thus, this construction maintains the one-to-one correspondence between the contingency sets of $D$ for $Q_{L}$ and the cuts of the network that have finite cost: selecting facts to remove in the contingency set corresponds to cutting edges in the network, and ensuring that the contingency set does not satisfy the query corresponds to ensuring that no source to target paths remain. Furthermore, this is the sum of multiplicities of the facts in a contingency set is equal to the cost of the corresponding cut, so that minimum cuts correspond to minimum contingency sets. We can then, as in the case of local languages, simply solve the MinCut problem for this network in $\tilde{O}(|A|\times|D|^{2}\times|\Sigma|^{2})$, concluding the proof. ∎
7.2. One-Dangling Languages
We now present another example of tractable languages, dubbed one-dangling languages, for which tractability is shown by a more involved reduction to a flow problem.
Definition 7.8 (One-dangling language).
A one-dangling language is a language that can be written as $L\cup\{xy\}$ where $L$ is a local language over some alphabet $\Sigma$ and where $x$ and $y$ are distinct letters such that at least one of them is not in $\Sigma$.
Note that one-dangling languages cover in particular the more restricted class of languages for which tractability had been shown in the conference version of this paper (amarilli2025resilience, , Proposition 7.7). They also cover the two languages $abcd|be$ and $ax^{*}b|xd$ for which the complexity of resilience had been left open in (amarilli2025resilience, ). So we can subsume that result by showing the following:
Proposition 7.9.
[…]
In addition to covering more languages, this result is a combined tractability result instead of a data complexity result, and it gives a better polynomial exponent than (amarilli2025resilience, , Proposition 7.7). This is because our proof will proceed via a reduction to a flow problem, instead of leveraging the more complicated technique of submodular function minimization used in (MCCORMICK2005321, ).
We prove Proposition 7.9 in the rest of this subsection, before finally turning to other NP-hard cases in the rest of the section. We note that an argument used the proof was first posted online (cstheory_martin, ). We do a slight abuse of notation throughout the proof for convenience: for $D^{\prime\prime}$ a bag database and $L^{\prime\prime}$ a language, we will talk of a contingency set of $D^{\prime\prime}$ for $L^{\prime\prime}$ simply to mean a contingency set of $D^{\prime\prime}$ for $Q_{L^{\prime\prime}}$.
Due to Proposition 6.3, if there is a PTIME solution for $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\})$, then there is a PTIME solution for $\mathtt{RES}_{\mathrm{bag}}(L^{R}\cup\{yx\})$, and the mirror of a local language is local (as can easily be seen from Proposition 3.5). Hence, we can assume without loss of generality that $y\not\in\Sigma$.
[…]
Let us now define how the input bag database $D$ over $\Sigma\cup\{y\}$ is rewritten to a bag database $D^{\prime}$ over $\Sigma\cup\{z\}$. To do this, we will extend the semantics of bag graph databases to what we call the extended bag semantics, where we allow the function $\mathtt{mult}$ to give negative multiplicities, or multiplicity zero, to edges of the database. The rest of the semantics remains unchanged, and the definition of resilience is unchanged as well – noting that, when computing resilience over database in extended bag semantics, one can always assume that facts with a non-positive multiplicity are removed. For this reason, it is easy to see that $\mathtt{RES}_{\mathrm{bag}}(L)$ under the extended semantics reduces in linear time to $\mathtt{RES}_{\mathrm{bag}}(L)$ under the normal bag semantics. Let us use the notation $\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L)$ for a fixed language $L$ to denote the resilience problem when posed over databases in extended bag semantics.
We can now describe the transformation. Given a bag database $D$, let $V$ be the domain of $D$. For every node $v\in V$, let $D^{-}_{x}(v)$ be the set of edges of the form $u\xrightarrow{x}v$ for some $u\in V$, and let $D^{+}_{y}(v)$ be the set of edges of the form $v\xrightarrow{y}u$ for some $u\in V$. We rewrite $D$ by doing the following for each $v\in V$:
  1. (1)
    add a fresh node $(v,{\sf in})$ into $V$;
  2. (2)
    replace all edges of the form $u\xrightarrow{x}v\in D^{-}_{x}(v)$ by the edge $u\xrightarrow{x}(v,{\sf in})$, keeping the same multiplicity;
  3. (3)
    add an edge $(v,{\sf in})\xrightarrow{z}v$ into $D$ and define $\mathtt{mult}((v,{\sf in})\xrightarrow{z}v)\coloneq\sum_{e\in D^{-}_{x}(v)}\mathtt{mult}(e)-\sum_{e\in D^{+}_{y}(v)}\mathtt{mult}(e)$ (which may be negative); and
  4. (4)
    erase any edges in $D^{+}_{y}(v)$ from $D$.
Let us call $D^{\prime}$ the resulting bag database: it is a bag database in extended semantics, and it is over the alphabet $\Sigma\cup\{z\}$ because applying the last point of the list above over all $v\in V$ will remove all $y$-facts. Let us define the value $\kappa\coloneq\sum_{v\in V}\sum_{e\in D^{+}_{y}(v)}\mathtt{mult}(e)$ (evaluating this expression in $D$). Note that $\kappa$ can be computed in linear time from $D$ as it is simply the sum of multiplicities of all edges labeled $y$ in $D$.
We now claim the following:
[…]
which suffices to conclude, because we have computed $\kappa$ and we can compute the resilience expression in the right-hand side using Theorem 3.13: the language $L^{\prime}$ is local, and as we explained the complexity of computing resilience is unchanged when working over extended bag semantics databases. To show Equation ($\dagger$7.2), we will show the following claim, and explain after its proof how the lemma implies Equation ($\dagger$7.2):
Claim 7.10.
(i) For any minimal contingency set $C$ of $D$ for $L\cup\{xy\}$, there exists a contingency set $C^{\prime}$ of $D^{\prime}$ for $L^{\prime}$ such that $\sum_{e\in C}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$; and (ii) for any minimal contingency set $C^{\prime}$ of $D^{\prime}$ for $L^{\prime}$, there exists a contingency set $C$ of $D$ for $L\cup\{xy\}$ such that $\sum_{e\in C}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$.
We end the proof of Proposition 7.9 by explaining why Equation ($\dagger$7.2) follows from 7.10. First, suppose that $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\},D)<\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L^{\prime},D^{\prime})+\kappa$. Let $C^{*}$ be a minimal contingency set of $D$ for $L\cup\{xy\}$, and let $C^{\prime*}$ be a minimal contingency set of $D^{\prime}$ for $L^{\prime}$. We have that $\sum_{e\in C^{*}}\mathtt{mult}(e)<\kappa+\sum_{e\in C^{\prime*}}\mathtt{mult}(e)$. However, due the to point (i) in 7.10, we know that there exists a contingency set $C^{\prime}$ of $D$ for $L^{\prime}$ such that $\sum_{e\in C^{*}}\mathtt{mult}(e)=\kappa+\sum_{e\in C^{\prime}}\mathtt{mult}(e)$, and so $\sum_{e\in C^{\prime}}\mathtt{mult}(e)<\sum_{e\in C^{\prime*}}\mathtt{mult}(e)$, which contradicts the fact that $C^{\prime*}$ is minimal. Now suppose that $\mathtt{RES}_{\mathrm{bag}}(L\cup\{xy\},D)>\mathtt{RES}_{\mathrm{bag}}^{\sf ex}(L^{\prime},D^{\prime})+\kappa$. We reach a contradiction using an analogous argument and using point (ii) in 7.10. Thus, we have concluded the proof of Proposition 7.9.
7.3. Additional NP-Hard Cases
Already in the case of finite infix-free languages, there are gaps remaining between the tractable cases (local languages, and languages covered by Proposition 7.6) and the hard cases (languages containing a word with a repeated letter, and four-legged languages) that we have identified so far. We show below that an unclassified languages is NP-hard by building a specific gadget:
Proposition 7.11.
$\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$ and $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$ are NP-hard.
Proof.
Figure 15. Completed gadget used in the proof of Proposition 7.11 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|be|ef)$
Figure 16. Completed gadget used in the proof of Proposition 7.11 regarding $\mathtt{RES}_{\mathrm{set}}(abcd|bef)$
We claim that the gadget shown in Figure 15 is valid for $abcd|be|ef$ and the gadget shown in Figure 16 is valid for $abcd|bef$ (also, they happen to be identical). We can see that the gadgets meet all the criteria of a pre-gadget, and show that there exists a condensed hypergraph of matches of the pre-gadgets for these queries that is an odd path. These properties can be checked by exhaustively considering all walks in the gadget, and we conclude that these gadgets are valid hardness gadgets for $abcd|be|ef$ and $abcd|bef$. ∎
8. Conclusion and Future Work
We have studied the complexity of resilience for RPQs, towards the goal of characterizing the languages for which the problem is tractable. We study bag and set semantics, but our results so far show no difference between the two settings, unlike for CQs (makhija2024unified, ). Refer back to Figure 1 on page 1 for a summary of our results and of some remaining open cases. We leave open the general question of whether there are languages $L$, regular or not, for which $\mathtt{RES}_{\mathrm{set}}(L)$ is PTIME but $\mathtt{RES}_{\mathrm{bag}}(L)$ is NP-hard. We also leave open the question of classifying the complexity of resilience in intermediate settings, e.g., bag semantics where only weights $1$ and $\infty$ are allowed; or the setting where some relations are indicated as exogenous in the signature, which amounts to giving a weight of $\infty$ to all their facts. Other settings left open by our work are those where we delete vertices (or just vertex labels) instead of edges, or where vertices and/or edges can be labeled by words or by non-empty words instead of single letters.
Our work does not fully characterize the tractability boundary and leaves some cases open. In particular, at a high level, we still do not fully understand the non-local languages that are not four-legged, i.e., non-letter-Cartesian languages featuring counterexamples where some of the “legs” are empty.
Another open question concerns non-Boolean RPQs, i.e., where endpoints are fixed (or given as input). For instance, the RPQ $aa$ then seems tractable as possible contingency sets have a simple structure; but longer RPQs may stay hard. This problem may relate to length-bounded cuts (itai1982complexity, ; 10.1145/1868237.1868241, ), or multicuts (garg1994multiway, ) and skew multicuts (kratsch2015fixed, ). Note that this non-Boolean setting corresponds to the problem of deletion propagation with source side effects (buneman2002propagation, ), which was studied before resilience: this problem asks about finding a minimum contingency set to remove a specific query answer. The case of removing multiple tuples has also been investigated (kimelfeld2013multituple, ). We do not know whether a dichotomy in one of these settings implies a dichotomy for our problem, or vice-versa. We also do not understand how our results relate to VCSPs (bodirsky2024complexity, ), e.g., whether classes of VCSPs with better algorithms (in combined complexity, or in terms of the polynomial degree) can imply tractable algorithms for some RPQ classes; or whether VCSP dichotomies hold in set semantics (or in bag semantics with weights restricted to $1$ and $\infty$).
Last, resilience can be studied for recursive queries beyond RPQs, in particular two-way RPQs or Datalog, though this would require new techniques (unlike RPQs, these queries are not “directional”). We note that for two-way RPQs, and more generally for queries expressible as monadic Datalog, the existence of finite duals implies a dichotomy on the resilience problem in bag semantics via the VCSP techniques of (bodirsky2024complexity, ).
9. Acknowledgments
The authors are grateful to Charles Paperman for help about local languages, to Rémi de Pretto for pointing out a minor problem, to Blaz Korecic and Martín Andrighetti for insightful discussions towards the proof of Proposition 7.9, and to the reviewers of the conference version for their valuable feedback. Part of this work was conducted while four of the authors were participating in the semester-long program Logic and Algorithms in Database Theory and AI at the Simons Institute for the Theory of Computing. Amarilli was partially supported by the ANR project EQUUS ANR-19-CE48-0019, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 431183758, and by the ANR project ANR-18-CE23-0003-02 (“CQFD”). Gatterbauer and Makhija were partially supported by the National Science Foundation (NSF) under award number IIS-1762268.